### v2

# Bourbaki degree of pairs of projective surfaces

Abstract
The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials - defined as the kernel of the associated Jacobian matrix - naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation. As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
The present work focuses on sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly-free cases. In particular, one of the nearly-free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
1 Introduction
Let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$ variables with coefficients in an algebraically closed field $\kappa$. For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{1}+1,\ldots,d_{k}+1$, respectively, with $d_{1}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
$\nabla\sigma=\begin{pmatrix}\nabla f_{1}\\ \nabla f_{2}\\ \vdots\\ \nabla f_{k}\end{pmatrix}:\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i}).$
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$. It is considered (see [11]) in analogy with the case of the tangent logarithmic sheaf associated to hypersurfaces in $\mathbb{P}^{n}$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
In the case $k=1$ of divisors on $\mathbb{P}^{n}$, one has a short exact sequence of the form:
$0\rightarrow{\mathcal{T}}_{f}\rightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\mathcal{I}_{J_{f}}(d)\rightarrow 0,$
where $J_{f}=\langle\partial_{0}f,\ldots,\partial_{n}f\rangle\trianglelefteq R$ is the Jacobian ideal of the homogeneous polynomial $f$. The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated to the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$. For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated to the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
One should consider sequences rather than the associated subvarieties $X=V(\sigma)$ if $k>1$, because of the following observation: different choices of sequences $\sigma^{\prime},\sigma$ generating the same ideal may have different Jacobian syzygy modules (see [11, Example 2.7]). When $k=1$, however, any different choice $f^{\prime}$ for generator of the ideal $\langle f\rangle$ gives a linear multiple of $\nabla f$, so that ${\mathcal{T}}_{f}\simeq{\mathcal{T}}_{f^{\prime}}$. Similarly, when $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and in particular ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [11, Lemma 2.14]).
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$. Following the terminology of [11], we say that a sequence $\sigma=(f_{1},\ldots,f_{k})$ is compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$. This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$. Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [11, Lemma 2.8]).
One may also consider the relationship with Bourbaki ideals, as in [17] and [7]. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{n}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B_{\nu}}(e-d)\rightarrow 0$
where $d=-c_{1}({\mathcal{T}}_{\sigma})$ and $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection, since ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank two. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [17]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The focus of this work is to understand the above concept for $n=3$ and $k=2$.
For $k=1$, the singular scheme of $V(f)$, defined by the Jacobian ideal $\Sigma_{f}\doteq V(J_{f})\subset\mathbb{P}^{n}$, plays an important role in this study. If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma)}{s!}t^{s}+O(t^{s-1}),$
and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [17]):
$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}),$
where $e=\operatorname{indeg}({\mathcal{T}}_{f})$.
Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for ${\mathcal{T}}_{\sigma}$ to get an ideal sheaf as the cokernel. This problem is addressed in the recent work [7] where the authors consider tame hypersurfaces and other notions of Bourbaki degrees. In this paper, we focus on the case of sequences $\sigma=(f,g)$ on $\mathbb{P}^{3}$, so our logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has also rank two, and we consider the same approach as in [17], since we only have to choose one syzygy to do the construction.
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$. In [11], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$. In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$, and $c_{1}(\mathcal{Q}_{\sigma})$ coincides with the degree of the greatest common divisor among all $(k\times k)$-minors of $\nabla\sigma$. The generic case is when $c_{1}(\mathcal{Q}_{\sigma})=0$, and this $g.c.d.$ is one, in which case we call $\sigma$ a normal sequence (see 2.1).
If we assume $c_{1}(\mathcal{Q}_{\sigma})=0$, we obtain a Hilbert polynomial of the form
$H(\mathcal{Q}_{\sigma},t)=\frac{\deg(\mathcal{Q}_{\sigma})}{s!}t^{s}+O(t^{s-1}),$
where $s\leq n-2$. We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ has codimension at least three.
In the case $n=3$ and $k=2$, we have sequences $\sigma=(f,g)$ of homogeneous polynomials with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, setting $d\doteq d_{f}+d_{g}$, so our objects are pairs of projective surfaces. We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and thus $c_{1}({\mathcal{T}}_{\sigma})=-d$. Since ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we may reproduce analogous considerations as in [17]. We show a formula for the Bourbaki degree of $\sigma$ given by
$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma),$
where $m(\sigma)=\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$ is the degree of the sheaf $\mathcal{Q}_{\sigma}$ as defined above. The discrete invariant $m(\sigma)$ plays the role of the Tjurina number in the previous context. We compare the different scheme structures between $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ in Example 2, considering a free sequence $\sigma=(f,g)$ of polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with $d_{f}=1,d_{g}=2$ and so that $m(\sigma)=5$, but $\deg(\Xi_{\sigma})=6$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))=4$ as non-reduced schemes, showing the relationship between $\deg(\mathcal{Q}_{\sigma})$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))$ is more complicated than in the case $k=1$, where these two coincide for the sheaf $\mathcal{O}_{J_{f}}(d)$.
About the initial degree and the invariant $m=m(\sigma)$, we show some bounds that are analogous to the case of plane projective curves:
Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
  • (a)
    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
  • (b)
    $m(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$;
  • (c)
    The following are equivalent:
    • (1)
      $m(\sigma)=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$;
    • (2)
      $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$;
    • (3)
      ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$;
    • (4)
      $\sigma$ is compressible;
We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 3.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 9): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable. We also show the bound $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, which is attained for example for regular pencils with singular members having only isolated singularities (see Remark 10).
In Section 3.2, we relate our notion to the notion of the Bourbaki degree for a projective plane curve, by considering a reduced polynomial $g\in\kappa[x_{0},x_{1},x_{2}]$ together with its associated surface $V(g)\subset\mathbb{P}^{3}$ and projective curve $X=V(g)\subset\mathbb{P}^{2}$, to obtain $\operatorname{Bour}(X)=\operatorname{Bour}(g,x_{3})$.
Afterwards, in Section 3.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is a Bourbaki scheme obtained by the choice of a minimal syzygy. Related to the free resolutions, we introduce the notions of nearly-free sequences and $3$-syzygy sequences. We have the following chain of implications:
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3-\text{syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
and the converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$.
In Section 4, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [11, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [17, Corollary 2.11]). The main theorem of the section is:
Theorem B. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
  • (a)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, then $\operatorname{Bour}(\sigma)\in\{0,1,2\}$;
  • (b)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
In Section 5, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. In these two classes, we classify all free and nearly free cases in terms of their discrete invariants and establish some stability results, which are summarized below:
Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 41), which is strictly $\mu$-semistable;
  • (d)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 16), but the bound above for stability may not be sharp.
Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 54 and Example 15);
  • (d)
    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Example 15 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable. This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [2]:
Conjecture. If the tangent sheaf of a codimension one foliation on $\mathbb{P}^{3}$ is not split, then it is $\mu$-semistable.
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([12]) for the two families of sequences.
Acknowledgments
FM is financed by the São Paulo Research Foundation (FAPESP), Brasil. Process Number #2021/10550-4. FM is currently a PhD. student under the advice of Marcos Jardim and Daniele Faenzi, with partial funding by the Bridges ”Brazil-France interplays in Gauge Theory, extremal structures and stability” projects ANR-21-CE40-0017 and ANR-17-EURE-0002. We thank Marcos Jardim, Daniele Faenzi and Alan Muniz for fruitful discussions and suggestions regarding this work.
2 Basic setting and $m$-syzygy sequences on $\mathbb{P}^{n}$
We work over an algebraically closed field $\kappa$ of characteristic zero, although most of the definitions can be considered under relaxed assumptions. We denote by $R=\kappa[x_{0},\ldots,x_{n}]$ the graded ring of polynomials over $\kappa$ and $\mathbb{P}^{n}\doteq\mathbb{P}^{n}_{\kappa}=\operatorname{Proj}R$ the projective $n$ space over $\kappa$.
In this section, we recall from [11] the definition of the logarithmic tangent sheaf associated to a sequence of $k$ polynomials, and introduce some notation inspired by the literature about the case $k=1$.
For a sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials with degrees $d_{1}+1,\ldots,d_{k}+1$, we consider the Jacobian matrix $\nabla\sigma$ as a morphism between locally free sheaves:
$\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\xrightarrow{\nabla\sigma}\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i})$
so the kernel sheaf ${\mathcal{T}}_{\sigma}\doteq\ker(\nabla\sigma)$, is called the logarithmic tangent sheaf associated to $\sigma$. We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel sheaf and by $\mathcal{M}_{\sigma}\doteq\operatorname{im}(\nabla\sigma)$ the image sheaf. We denote by $d\doteq\sum_{i=1}^{n}d_{i}$ the total degree of a sequence $\sigma$. Assuming the generic rank of $\operatorname{im}(\nabla\sigma)$ is $k$, it follows that ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank n+1kn+1-k.
2.1 Normal sequences
By construction, the reduced support of the sheaf $\mathcal{Q}\sigma$ coincides with the vanishing locus $\Xi\sigma\doteq V(\bigwedge^{k}\nabla\sigma)$ of the $k\times k$-minors of the matrix $\nabla\sigma$. Assuming $\sigma$ is algebraically independent, the generic rank of $\nabla\sigma$ is $k$, so that $\operatorname{{codim}}\Xi_{\sigma}\geq 1$.
Moreover, since $\mathcal{Q}\sigma$ is a torsion sheaf, the condition $c_{1}(\mathcal{Q}\sigma)=0$ is equivalent to $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$. In terms of the matrix $\nabla\sigma$, $\operatorname{{codim}}\Xi_{\sigma}\geq 2$ if and only if the greatest common divisor among all $(k\times k)$-minors is one. When $k=1$, we have
$\operatorname{{codim}}(Z_{f})=\operatorname{{codim}}(\mathcal{Q}_{f})\geq 2\iff V(f)\text{ is normal}.$
Inspired by the behavior above, we call a sequence $\sigma$ normal when $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$.
We note that the two schemes $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ may have different scheme structures and, in particular, different degrees. On one hand, the scheme structure of the support $\operatorname{supp}(\mathcal{Q}_{\sigma})=V(\operatorname{Ann}\mathcal{Q}_{\sigma})$ is induced by the annihilator ideal sheaf of $\mathcal{Q}_{\sigma}$, defined locally as the annihilator ideal of the associated module. On the other hand, the scheme structure of $\Xi_{\sigma}$ is induced locally by the 0-th Fitting ideal of the $\mathcal{Q}_{\sigma}$. These schemes have the same reduced locus, since the support of these two ideals coincide, but in general $\operatorname{Fitt}_{0}\subseteq\operatorname{Ann}$ (see, for example, [9, Proposition 20.7]).
For $k=1$, when $\sigma=f$, we have an isomorphism $\mathcal{Q}f\simeq\mathcal{O}{Z_{f}}$. This holds because $\mathcal{M}f\simeq\mathcal{I}{Z_{f}}(d)$ is the ideal sheaf of the scheme $Z_{f}$, and whenever $f$ is reduced, the schemes $\Xi_{f}$ and $\operatorname{supp}(\mathcal{Q}{\sigma})$ coincide.
If $\sigma$ is a normal sequence, then the Hilbert polynomial of the cokernel sheaf $\mathcal{Q}_{\sigma}$ is of the form
$\operatorname{Hilb}(\mathcal{Q}_{\sigma},t)=\frac{m(\sigma)}{(n-2)!}t^{n-2}+\ldots$
where $m(\sigma)\geq 0$ is a discrete invariant so that $m(\sigma)=0$ if and only if $\operatorname{{codim}}(\mathcal{Q}_{\sigma})\geq 3$. However, $m(\sigma)$ generally coincides with neither the schematic degree of $\Xi_{\sigma}$ nor that of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, as we explore in Example 2 at the end of this section.
2.2 $m$-syzygy sequences and exponents
We may consider a minimal free resolution of the sheaf ${\mathcal{T}}_{\sigma}$, which will be given in the following form:
$\ldots\rightarrow\bigoplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{n}}(-e_{i})\xrightarrow{\rho}{\mathcal{T}}_{\sigma}\rightarrow 0.$
In which case, we say $\sigma$ is an $m$-syzygy sequence.
We note that $m\geq n$ and $\sigma$ is said to be free if and only if ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, or, equivalently, when $m=n$.
The integers $e_{1}\leq\ldots\leq e_{m}$ will be called exponents, and the first $\operatorname{indeg}({\mathcal{T}}_{\sigma})\doteq e_{1}=e$ will be called the initial degree of the sheaf ${\mathcal{T}}_{\sigma}$. This is the minimum integer $e\geq 0$ such that there is a non-trivial syzygy $\nu$ of degree $d$ for the Jacobian matrix $\nabla\sigma$. Since there is an injection ${\mathcal{T}}_{\sigma}\hookrightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}$ and the latter is a $\mu$-semistable sheaf, all exponents satisfy $e_{i}\geq 0$.
A sequence $\sigma=(f_{1},\ldots,f_{k})$ is called compressible if, after a linear change of variables, there is a variable that does not occur in any of the $f_{1},\ldots,f_{k}$. The number of variables omitted in this way will give rise to syzygies of degree zero for $\nabla\sigma$, so the following hold (for proofs, see [11, Lemma 2.7 and Lemma 2.8]):
Lemma 1.
The following quantities coincide:
  • The number of variables omitted in $f_{1},\ldots,f_{k}$;
  • The discrete invariant $h^{0}({\mathcal{T}}_{\sigma})$;
  • The number of 0’s in a vector of exponents for ${\mathcal{T}}_{\sigma}$.
In particular, $e=0$ if and only if $\sigma$ is compressible.
For $n=3$ and when $\sigma=(f,g)$ is a normal sequence, we show the conditions above are equivalent to $m(\sigma)$ attaining its maximum value, see Section 3.1.
Example 2.
Let $\sigma=(2x_{1}x_{3}-x_{1}^{2},3x_{2}x_{3}^{2}-3x_{0}x_{1}x_{3}+x_{1}^{3})$. The Jacobian matrix is of the form
$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$
and the matrix below
$\begin{pmatrix}x_{3}&x_{0}x_{1}-x_{1}^{2}-2x_{2}x_{3}\\ 0&-x_{1}x_{3}\\ x_{1}&-2x_{2}x_{3}\\ 0&-x_{1}x_{3}+x_{3}^{2}\end{pmatrix}$
gives two linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
In this case, the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, given by:
$\begin{cases}\operatorname{Ann}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively. However, from the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ we conclude $m(\sigma)=5$.
3 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (3.1), followed by a reduction to the case of a projective plane curve (3.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (3.3), in particular introducing the class of nearly-free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
3.1 Framework and first results
By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$. We also work only with normal sequences, as defined previously in 2.1. By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
Each sequence $\sigma=(f,g)$ induces a morphism of sheaves on $\mathbb{P}^{3}$ by the Jacobian matrix:
$\nabla\sigma:\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(d_{g}),$
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [11].
The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
Lemma 3.
Let $\mathcal{Q}$ be a coherent sheaf on $\mathbb{P}^{3}$ with $\operatorname{{rk}}(\mathcal{Q})=0$ and $c_{1}(\mathcal{Q})=0$. Then the Hilbert polynomial of $\mathcal{Q}$ is given by
$\mathcal{X}(\mathcal{Q}(t))=\operatorname{ch}_{2}(\mathcal{Q})t+\operatorname{ch}_{3}(\mathcal{Q})+2\operatorname{ch}_{2}(\mathcal{Q}).$
Proof.
Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$. From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
Any saturated syzygy of the Jacobian matrix $\nu$, of degree $e\in\mathbb{Z}$, induces a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described by the Serre correspondence (see [14, Theorem $4.1$]). Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$. Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
Proposition 4.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
For any saturated syzygy $\nu\in H^{0}({\mathcal{T}}(e))$ of degree $e\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated closed subscheme of pure dimension one. Then, we have the following equation:
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+m_{0}-m(\sigma),$
where $m_{0}\doteq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$.
Proof.
From the hypothesis $c_{1}({\mathcal{Q}}_{\sigma})=0$, we conclude $c_{1}({\mathcal{T}}_{\sigma})=-(d_{f}+d_{g})$ and $\operatorname{ch}_{2}({\mathcal{Q}}_{\sigma})=m(\sigma)$. Moreover, from additivity of $\operatorname{ch}_{2}$ on short exact sequences:
$\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ $=-\operatorname{ch}_{2}({\mathcal{O}}_{\mathbb{P}^{3}}(d_{f})\oplus{\mathcal{O}}_{\mathbb{P}^{3}}(d_{g}))+\operatorname{ch}_{2}({\mathcal{Q}}_{\sigma})$
$=-\frac{d_{f}^{2}+d_{g}^{2}}{2}+m(\sigma).$
We can relate the Chern character and the Chern classes by the formula
$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
$=\frac{(d_{f}+d_{g})^{2}}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}-c_{2}({\mathcal{T}}_{\sigma})$
$=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-c_{2}({\mathcal{T}}_{\sigma}),$
so that
$c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma).$ (1)
On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(e))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(e))=c_{2}({\mathcal{T}}_{\sigma})+e\cdot c_{1}({\mathcal{T}}_{\sigma})+e^{2}.$
From this, together with 1, we obtain
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
We define the Bourbaki degree of a sequence $\sigma$, inspired by [17, Definition 2.4].
Definition 5.
Let $\sigma$ be a normal sequence and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. The Bourbaki degree of a normal sequence $\sigma$ is defined by:
$\operatorname{Bour}(\sigma)\doteq\deg(B_{\nu})=e(e-d)+m_{0}-m(\sigma),$
for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. We note that every non-trivial syzygy of minimal degree is saturated.
Remark 6.
It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$.
Remark 7.
From [13, Proposition 4.1], we have the formula
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
relating the third Chern class of ${\mathcal{T}}_{\sigma}$ and the discrete invariants of $B$, when $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $B$ is the zero locus of a non-zero section in $H^{0}({\mathcal{T}}_{\sigma}(e))$. Moreover, dualizing the sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B}(e-d)\rightarrow 0$
we conclude that the singular set of the sheaf ${\mathcal{T}}_{\sigma}$ is contained in $B$.
Now, we obtain some bounds for the quantities $\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ in terms of the degrees $d_{f},d_{g}$.
Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
  • (a)
    We have $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
  • (b)
    $m(\sigma)\leq m_{0}$;
  • (c)
    The following are equivalent:
    • (1)
      $m(\sigma)=m_{0}$;
    • (2)
      $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$;
    • (3)
      $\sigma$ is compressible;
    • (4)
      ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$;
Proof.
To show $(a)$, we build explicit syzygies of the Jacobian matrix $\nabla\sigma$ of degrees $d_{f}+d_{g}$, and at least one of them is nonzero. Writing the Jacobian matrix by
$\nabla\sigma=\begin{pmatrix}\partial_{0}f&\partial_{1}f&\partial_{2}f&\partial_{3}f\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&\partial_{3}g\\ \end{pmatrix}$,
the following vectors
$\nu_{0}$ $=\begin{pmatrix}0\\ \partial_{2}f\partial_{3}g-\partial_{3}f\partial_{2}g\\ -\partial_{1}f\partial_{3}g+\partial_{3}f\partial_{1}g\\ \partial_{1}f\partial_{2}g-\partial_{2}f\partial_{1}g\end{pmatrix};\nu_{1}=\begin{pmatrix}\partial_{2}f\partial_{3}g-\partial_{3}f\partial_{2}g\\ 0\\ -\partial_{0}f\partial_{3}g+\partial_{3}f\partial_{0}g\\ \partial_{0}f\partial_{2}g-\partial_{2}f\partial_{0}g\end{pmatrix}$;
$\nu_{2}$ $=\begin{pmatrix}\partial_{1}f\partial_{3}g-\partial_{3}f\partial_{1}g\\ -\partial_{0}f\partial_{3}g+\partial_{3}f\partial_{0}g\\ 0\\ \partial_{0}f\partial_{1}g-\partial_{1}f\partial_{0}g\end{pmatrix};\nu_{3}=\begin{pmatrix}\partial_{1}f\partial_{2}g-\partial_{2}f\partial_{1}g\\ -\partial_{0}f\partial_{2}g+\partial_{2}f\partial_{0}g\\ \partial_{0}f\partial_{1}g-\partial_{1}f\partial_{0}g\\ 0\end{pmatrix}$.
are a syzygies of degree $d_{f}+d_{g}$. Since there is at least one nonzero $2\times 2$ minor, from the hypothesis of algebraically independent, at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
For $(b)$, using that $e\leq d$, we can use that
$0\leq\operatorname{Bour}(\sigma)=e(e-d)+m_{0}-m(\sigma),$
so that $m(\sigma)\leq m_{0}+e(e-d)$, but $e(e-d)\leq 0$ and the claim follows.
For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [11, Lemma 2.7]. Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [11, Lemma 2.8]. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
To show $(1)\Rightarrow(2)$, let $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. From the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=e(e-d)$. Since $\operatorname{Bour}(\sigma)\geq 0$ and from part (a) we have $e\leq d$, it follows that $e=0$ or $e=d$. Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free. Since $c_{1}({\mathcal{T}}_{\sigma})=-d$ is additive, it follows that $e=0$, otherwise we would have $c_{1}({\mathcal{T}}_{\sigma})<-d$. ∎
Remark 8.
The bound obtained in 3.1, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [11, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [2, Theorem 3.1],
$c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C).$
Using the formula $c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$ and the equations
$c_{2}({\mathcal{T}}_{\sigma})-d+1=c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C)$
we obtain $m(\sigma)=\deg(C)-d-d_{f}d_{g}-1$, so from the bound above we get
$\deg(C)-d-d_{f}d_{g}-1=m(\sigma)$ $\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g},$
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [20, Corollary 4.8].
From simple observations about the formula $\operatorname{Bour}(\sigma)$, we are able to obtain the following inequalities related to $\mu$-stability and freeness of the logarithmic sheaves.
Proposition 9.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
  • (a)
    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if and only if $m(\sigma)=0$ and $e=d$.
  • (b)
    If $\operatorname{Bour}(\sigma)>(d_{f}-1)(d_{f}+d_{g})+d_{g}^{2}+1$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
  • (c)
    If
    $m(\sigma)<\frac{1}{2}\left(\frac{3d_{f}^{2}}{2}+\frac{3d_{g}^{2}}{2}+d_{f}d_{g}\right),$
    then $\sigma$ is not free.
Proof.
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$. This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility. The converse also follows simply from the formula.
By construction, $m(\sigma)\geq 0$, and therefore for a given value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, the Bourbaki degree of $\sigma$ can be at most
$\operatorname{Bour}(\sigma)\leq e(e-d)+m_{0},$
and this is a function $H=H(e)$ which attains its minimum at $e=d/2$. The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, and if $e>d/2$, ${\mathcal{T}}_{\sigma}$ is $\mu$-stable, since $\mu({\mathcal{T}}_{\sigma})=-d/2$. Thus, the maximum value $\operatorname{Bour}(\sigma)$ in the range $e\in\{1,\ldots,d/2\}$ for any possible value of $m(\sigma)$ is $H(1)$, which is the expression above on the right-hand side. If $\operatorname{Bour}(\sigma)$ is higher than this, then $e>d/2$, and thus we obtain $(b)$.
With the same strategy as $(b)$, since the function $H(e)$ attains its minimum at $e=d/2$, for $m(\sigma)$ satisfying the inequality of the claim we obtain that
$\operatorname{Bour}(\sigma)=H(e)\geq H(d/2)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. ∎
Remark 10.
If $\sigma=(f,g)$ is a pencil of surfaces of the same degree, say $d_{f}=d_{g}=p$, then $m(\sigma)=0$ if $\sigma$ is a regular pencil and, moreover, every singular member $V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}$ is normal. This follows from the description
$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g))$
obtained in [11, Lemma 2.17], since normal projective surfaces have isolated singularities and regular sequences have only a finite number of singular members, therefore $\dim(\operatorname{supp}(\mathcal{Q}_{\sigma}))=0$.
Example 11.
We consider the sequence
$\sigma=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$.
Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a locally free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
Remark 12.
We have not found any examples of sequences with $m(\sigma)=0$ so that $e<d$, which motivates the conjecture: does $m(\sigma)=0$ implies $e=d$?
3.2 The Bourbaki degree of a plane curve from the Bourbaki degree of a pair of surfaces
In this section, we obtain the formula of the Bourbaki degree of a projective plane curve $X=V(g)\subset\mathbb{P}^{2}$, where $g\in\kappa[x_{0},x_{1},x_{2}]$ is a reduced homogeneous polynomial, introduced in [17], as a special case of our formula, by considering the pair $\sigma=(x_{3},g)$ in $\mathbb{P}^{3}$.
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$. We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$.
The matrix $\nabla\sigma$ will be given by
$\nabla\sigma=\begin{pmatrix}0&0&0&1\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&0\end{pmatrix}$.
Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ the vector, so we denote by
${\mathcal{T}}_{\overline{g}}\doteq\ker(\nabla\overline{g})\hookrightarrow\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}\xrightarrow{\nabla\overline{g}}\mathcal{O}_{\mathbb{P}^{3}}(d_{g})$
the kernel of the multiplication by this vector. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
${\mathcal{T}}_{\overline{g}}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}$$\mathcal{O}_{\mathbb{P}^{3}}(d)$$\mathcal{O}_{Z_{g}}(d)$${\mathcal{T}}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}$$\mathcal{O}_{\mathbb{P}^{3}}(d)\oplus\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{Q}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{O}_{\mathbb{P}^{3}}$$\scriptstyle{\nabla\overline{g}}$$\scriptstyle{\nabla\sigma}$$\scriptstyle{\cdot 1}$
and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$.
From the discrete invariants considered $(d_{f}=0)$, we have the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}(\ker\nabla g(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
where $\tau(X)$ is the Tjurina number of the curve $X$, since $d_{f}=0$.
From the construction, we have that the singular scheme of ${\mathcal{T}}_{\sigma}$ (described by $Z_{g}$) intersects $V(x_{3})$ transversely, so we obtain the formula between Chern classes $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:\mathbb{P}^{2}\simeq V(x_{3})\hookrightarrow\mathbb{P}^{3}$ is the inclusion. Therefore,
$1-d[H]+(m_{0}-m(\sigma))[H]^{2}$ $=i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$
$=1-d[H]+(d^{2}-\tau(X))[H]^{2},$
and since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$.
Moreover, doing the analogous comparison for the total classes of the twist ${\mathcal{T}}_{\sigma}(e)$, we obtain
$i^{*}c({\mathcal{T}}_{\sigma}(e))=1+(e-d)[H]+\operatorname{Bour}(\sigma)[H]^{2}=1+(e-d)[H]+\operatorname{Bour}(X)[H]^{2}=c(i^{*}{\mathcal{T}}_{\sigma}(e)),$
and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$.
Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number, and thus it should be the correct generalization for the case of pairs of surfaces.
3.3 Locally free resolutions and the Bourbaki scheme
In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [17, Theorem 2.1, (c)]. This follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$.
Lemma 13.
Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a non-zero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated to $\nu$ in a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0.$
Then:
  • (a)
    Every free resolution for $\mathcal{I}_{B}$:
    $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\xrightarrow{\omega}\mathcal{I}_{B}\rightarrow 0$
    lifts for a free resolution of the form
    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0$
    for ${\mathcal{T}}_{\sigma}$.
  • (b)
    For a minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
    $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\lambda,\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0,$
    it induces a free resolution for $\mathcal{I}_{B}$ of the form:
    $0\rightarrow F_{2}(d-e)\rightarrow F_{1}(d-e)\rightarrow{}F_{0}(d-e)\xrightarrow{\lambda(d-e)}\mathcal{I}_{B}\rightarrow 0.$
Proof.
To show $(a)$, we apply the functor $\operatorname{Hom}(F_{0}(e-d),-)$ to the short exact sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0,$
to get the exact piece:
$\operatorname{Hom}(F_{0}(e-d),{\mathcal{T}}_{\sigma})\xrightarrow{\pi^{*}}\operatorname{Hom}(F_{0}(e-d),\mathcal{I}_{B}(e-d))\rightarrow\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))=0,$
since $\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(F_{0}^{\vee}(-2e-d))=0$, as $F_{0}^{\vee}$ is a direct sum of line bundles and these have vanishing first cohomology in $\mathbb{P}^{3}$. Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\omega(e-d)=\tilde{\omega}$. We now consider the map $\omega(e-d)\oplus\nu$ in the following commutative diagram with short exact sequences as the central two columns:
[…]
From the snake lemma, we obtain that $\operatorname{coker}(\tilde{\omega}\oplus\nu)=0$ and that $\ker(\tilde{\omega}\oplus\nu)\simeq\ker(\omega(e-d))$. Thus, we can continue the resolution for $\mathcal{I}_{B}$, twisting by $\mathcal{O}_{\mathbb{P}^{3}}(e-d)$, to obtain the following free resolution:
[…]
for ${\mathcal{T}}_{\sigma}$, as claimed.
To show $(b)$, we consider the diagram with exact rows induced by the fact above to obtain the short exact sequence in cokernels as the third row below:
[…]
Completing to the resolution and twisting accordingly, we obtain the resolution from the claim below.
[…]
Definition 14.
Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say $\sigma$ is:
  • nearly free if $\operatorname{Bour}(\sigma)=1$.
  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of Lemma 13, $(b)$.
Example 15 (Nearly free sequence with $d_{f}=1,d_{g}=2$).
We consider the following sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(x_{0}^{2}+x_{3}^{2},x_{0}^{3}+x_{0}x_{1}x_{2}+x_{3}^{3})$
with Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\end{pmatrix}$.
The 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the $1$-st Fitting ideal has dimension zero.
The minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=1$, $m(\sigma)=4$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=3$.
Example 16 ($3$-syzygy pencil of cubics which is not nearly-free).
Considering the following pencil of cubics:
$\sigma=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
where $f$ is a hyperplane arrangement and $g$ is the union of a plane and a smooth quadric, with the Jacobian matrix:
$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{0}x_{3}-x_{1}x_{3}&x_{0}x_{2}-x_{1}x_{2}\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
In this case, the 0-th Fitting ideal coincides with the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$, and the saturation of the first Fitting ideal is zero. The primary decomposition of $\Xi_{\sigma}$ is described in the following table:
dimension degree radical ideal
1 1 $(x_{2},x_{3})$
1 1 $(x_{3},x_{0}-x_{1})$
1 1 $(x_{2},x_{0}-x_{1})$
1 2 $(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for $\mathcal{I}_{B}$ will be of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}\rightarrow\mathcal{I}_{B}\rightarrow 0,$
presenting $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
For the rest of the section, we study some aspects of these special classes of sequences.
Proposition 17.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then:
  • (a)
    $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
    $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-1)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
    where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
  • (b)
    if $\sigma$ is nearly-free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
Proof.
This follows from Lemma 13, using the minimal free resolution for a line in $l\subset\mathbb{P}^{3}$ as the intersection of two planes:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\rightarrow\mathcal{I}_{l}\rightarrow 0$
since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for the syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. ∎
Remark 18.
The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [5], related to rational cuspidal curves. In [17, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [5] (see [17, Definition 2.17]). Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
The notion of 33-syzygy divisors is also present in a number of works in the field, for example [1], [6] and [8].
Example 19.
Consider $f=x_{0}^{3}+x_{0}x_{1}x_{2}+x_{3}^{3}$ and $g=x_{0}^{k+1}+x_{3}^{k+1}$ for $k\geq 3$. Then $d=d_{f}+d_{g}=k+2$ and:
  • For $k=2$, $(f,g)$ is a free pencil of cubics with $e=1$, $m(\sigma)=9$;
  • For $k>2$, $(f,g)$ is nearly-free with $e=1$.
For $k=2$, we note that the matrix below:
$\begin{pmatrix}0&-x_{0}x_{3}^{2}\\ x_{1}&0\\ -x_{2}&x_{2}x_{3}^{2}\\ 0&x_{0}^{3}\end{pmatrix}$
gives trivializing syzygies such that ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
For $k>2$, we set
$M=\begin{pmatrix}0&-x_{0}x_{2}x_{3}^{k}&-x_{0}x_{1}x_{3}^{k}\\ x_{1}&-3x_{0}^{k}x_{3}^{2}+3x_{0}^{2}x_{3}^{k}&0\\ -x_{2}&x_{2}^{2}x_{3}^{k}&-3x_{0}^{k}x_{3}^{2}+3x_{0}^{2}x_{3}^{k}+x_{1}x_{2}x_{3}^{k}\\ 0&x_{0}^{k+1}x_{2}&x_{0}^{k+1}x_{1}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}^{k}x_{3}^{2}-x_{0}^{2}x_{3}^{k}\\ \frac{1}{3}x_{1}\\ -\frac{1}{3}x_{2}\end{pmatrix}$,
so we obtain a free resolution of ${\mathcal{T}}_{\sigma}$ given by
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-1)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)^{\oplus 2}\xrightarrow{M}\mathcal{T}_{\sigma}\rightarrow 0,$
and in particular $\sigma$ is a nearly-free sequence.
Proposition 20.
If a normal sequence $\sigma=(f,g)$ is a $3$-syzygy, then $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$. Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
Proof.
First, if we assume $\sigma$ is $3$-syzygy, then there is a free resolution of the form:
$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we split the resolution into two short exact sequences:
$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma},$
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [14, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
$S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma}$
is a free resolution for ${\mathcal{T}}_{\sigma}$, concluding $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$.
For the equivalence stated above, if we start with a sequence $\sigma$ which is $3$-syzygy and then apply Lemma 13, $(b)$, we obtain a resolution for $\mathcal{I}_{B}$ which is of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
since $\operatorname{{rk}}(F_{0})=2$, thus concluding $B$ must be a complete intersection scheme. On the other hand, if $B$ is a complete intersection, then there is a resolution for $\mathcal{I}_{B}$ of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\xrightarrow{r,s}\mathcal{I}_{B}\rightarrow 0,$
induced by the two equations $r,s$ defining $B=V(r,s)\subset\mathbb{P}^{3}$. Thus, applying Lemma 13, $(a)$, we can lift the resolution above for a free resolution for ${\mathcal{T}}_{\sigma}$ of the form
$0\rightarrow F_{1}\rightarrow F_{0}^{\prime}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $\operatorname{{rk}}(F_{0}^{\prime})=3$, thus $\operatorname{{rk}}(F_{0})=2$ and $\sigma$ is $3$-syzygy. ∎
The following proposition follows from the formula in Remark 7 for $c_{3}({\mathcal{T}}_{\sigma})$ in terms of $e,\deg(B)$ and $p_{a}(B)$ for a Bourbaki scheme $B=B_{\nu}$.
Proposition 21.
If $\sigma=(f,g)$ is a nearly-free sequence such that ${\mathcal{T}}_{\sigma}$ is locally free, then $d=d_{f}+d_{g}$ must be even and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly-free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
$e=\frac{d+2}{2},$
which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$. ∎
Remark 22.
We note that there is a chain of implications:
$\sigma\text{ is nearly free }\Rightarrow\sigma\text{ is $3$-syzygy }\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
where the first follows from 20 since $\sigma$ is nearly free iff $B_{\nu}$ is a line for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and every line is a complete intersection of two planes. The converses do not hold, as we explore in the next examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 16), pencils of cubics which satisfy $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but are not $3$-syzygy (Example 23) and pencils of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (Example 24).
We also construct two pencils of cubics with the same discrete invariants $(m(\sigma),\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma))$ and the same Chern classes, which are distinguished by their homological behavior: one is $3$-syzygy and the other satisfies $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (see Example 23 and Example 25).
Example 23 (pencil of cubics which is not $3$-syzygy and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$).
Considering the following pencil of cubics:
$\sigma=(x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{2}^{3}+x_{1}x_{2}x_{3}+x_{3}^{3})$
with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\\ 0&x_{2}x_{3}&3x_{2}^{2}+x_{1}x_{3}&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$.
We may check computationally that the saturation of the ideal of $(2\times 2)$-minors coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$ and with the saturation of the 0-th Fitting ideal, therefore over this subset $\mathcal{Q}_{\sigma}$ has rank one. The $1$-dimensional part of this scheme has $3$ irreducible components, two lines $L_{1}=V(x_{0},x_{3}),L_{2}=V(x_{2},x_{3})$ and a quadric plane curve $Q=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})$. Moreover, $L_{1}$ and $Q$ have multiplicity two structure, adding to $m(\sigma)=5$.
This can also be obtained from the minimal free resolution of ${\mathcal{T}}_{\sigma}$:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{M}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. This implies that $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$.
Example 24 (pencil of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$).
We consider the sequence of cubics $d_{f}=d_{g}=2$ given by:
$\sigma=(x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3},x_{2}x_{3}(x_{2}-x_{1})),$
considered in [11, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{L}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
So we compute $e=3$ and $m(\sigma)=7$. This may also be seen from the primary decomposition of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, which is composed of three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$, where the last line has a multiplicity $5$ structure, and at every line the rank of $\mathcal{Q}_{\sigma}$ is one.
Example 25 (pencils of cubics with the same discrete invariants and different homological behavior).
Considering the following pencil of cubics:
$\sigma=(x_{0}^{3}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{3}^{3}+x_{1}x_{3}^{2}+x_{0}x_{1}x_{3}+x_{0}^{2}x_{2}),$
with the associated Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}+x_{3}^{2}&x_{0}^{2}&x_{0}x_{1}+2x_{1}x_{3}+3x_{3}^{2}\end{pmatrix}$.
Here, although the 0-th Fitting ideal is different from the annihilator ideal of $\mathcal{Q}_{\sigma}$, their one-dimensional component coincide with a multiplicity five structure along the line $V(x_{3},x_{0})$, and thus $m(\sigma)=5$.
Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 6}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$ and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$. A resolution for $\mathcal{I}_{B}$ will be of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 6}\rightarrow\mathcal{I}_{B}\rightarrow 0.$
This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as 16, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
In Section 5 we characterize all nearly free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly free sequences with $d_{f}=1,d_{g}=2$.
4 Extreme cases of low initial degree
In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$. We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 28. We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example [4]).
As explored in [11, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
$0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the foliation. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
[…]
The middle column of the previous diagram
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(1-e)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow G\rightarrow 0$
defines a foliation by curves of $\mathbb{P}^{3}$ of degree $e$, where $G$ is a rank two torsion-free sheaf and $G^{\vee}$ is called the conormal sheaf of the foliation. Dualizing this short exact sequence, we obtain
$0\rightarrow G^{\vee}\rightarrow\Omega^{1}_{\mathbb{P}^{3}}\rightarrow\mathcal{I}_{W}(e-1)\rightarrow 0,$
defining a subscheme $W\subset\mathbb{P}^{3}$, called the singular scheme of the associated foliation by curves. It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}$.
The classification of such foliations by curves of low degree provides the following bounds on their singular schemes:
  • (a)
    ([4, $2.3$, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1. In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
  • (b)
    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$ or is zero-dimensional.
Theorem B. Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
  • (a)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, then $\operatorname{Bour}(\sigma)\in\{0,1,2\}$;
  • (b)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
Proof.
To show $(a)$, we dualize the following short exact sequence, obtained above for $e=1$:
$0\rightarrow\mathcal{I}_{B}(2-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
to get a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-2)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+2)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}$ is defined by the short exact sequence $0\to\mathcal{U}\to\mathcal{O}_{\Gamma_{\sigma}}\to\mathcal{O}_{Y}\to 0$ and $Y$ is the one-dimensional component of $\Gamma\sigma$. Moreover, since $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$, we may consider the final piece of the long exact sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+2)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$, and from the classification $\deg(W^{\prime})\leq\deg(W)\leq 2$, hence the result follows.
For $(b)$, we proceed analogously, dualizing the sequence
$0\rightarrow\mathcal{I}_{B}(3-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0$
to obtain a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-3)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+1)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}\hookrightarrow\mathcal{O}_{\Gamma_{\sigma}}\rightarrow\mathcal{O}_{Y}$ and $Y$ is the one-dimensional component of $\Gamma_{\sigma}$. Moreover, $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$ and from the final piece of the sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+1)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{O}_{\Gamma_{\sigma}},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$. Since $W^{\prime}$ is a subscheme of $W$, we have $\deg(W^{\prime})\leq\deg(W)$. The classification of foliations by curves of degree $2$ implies $\deg(W)\leq 5$, which completes the proof. ∎
To be able to present the next applications, we will need the following result on the structure of degree $2$ space curves in $\mathbb{P}^{3}$:
Theorem 26.
[18, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree $2$ and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
  • (a)
    $a\geq-1$, and a=1a=-1 if and only if $B$ is planar;
  • (b)
    For $a\geq 1$, $B$ must be a multiplicity two structure at a line $L\subset\mathbb{P}^{3}$, and these satisfy a short exact sequence of the form
    $0\rightarrow\mathcal{O}_{L}(a)\rightarrow\mathcal{O}_{B}\rightarrow\mathcal{O}_{L}\rightarrow 0;$
  • (c)
    For $a\geq 1$, if $B$ is a multiplicity two structure on a line, then $\omega_{B}\simeq\mathcal{O}_{B}(-a-2)$. If $a=0$ and $B$ is a union of two skew lines, then $\omega_{B}\simeq\mathcal{O}_{B}(-2)$.
Proof.
Items $(a)$ and $(b)$ are in the original paper, and to obtain $(c)$, for $a\geq 1$ we can apply the functor $\operatorname{\mathcal{H}\textit{om}}(-,\omega_{\mathbb{P}^{3}})$ to the short exact sequence in $(ii)$, use that $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ and twist by $a+2$ to obtain a short exact sequence
$0\rightarrow\mathcal{O}_{L}(a)\rightarrow\omega_{B}(a+2)\rightarrow\mathcal{O}_{L}\rightarrow 0$
which coincides with the short sequence in $(b)$, hence we obtain the isomorphism in the claim. For the case $a=0$, $B$ must be either a multiplicity two structure as in $(b)$ with $a=0$ or a union of skew lines. In the first case, the proof follows as before. If $B=L_{1}\cup L_{2}$, then $\omega_{B}\simeq\omega_{L_{1}}\oplus\omega_{L_{2}}$ and the claim follows from the fact $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ in each component. ∎
Remark 27.
The results above can be used to restrict the possible values of numerical invariants. For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results above to show that the only possible discrete invariant is $c_{3}({\mathcal{T}}_{\sigma})=8$, although the existence of such an example remains an open question.
Indeed, from [19, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 7 we obtain $c_{3}({\mathcal{T}}_{\sigma})=8-2a$, where $p_{a}(B)=-1-a$. From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(4-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we obtain the Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(4-a))=10-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=10-2a,$
and additivity of Euler characteristic gives a contradiction whenever $a\neq 0$, so $c_{3}({\mathcal{T}}_{\sigma})=8$ if this occurs.
Using the classification of $\mu$-semistable reflexive sheaves on $\mathbb{P}^{3}$ with small $c_{2}$, we are able to obtain a full classification of the cases of pencils of cubics with $m(\sigma)=7$ and $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$ (see 42).
As another application, we are able to show that sequences $\sigma$ with mixed degrees $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$ cannot be unstable.
Proposition 28.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, and in particular ${\mathcal{T}}_{\sigma}$ must be $\mu$-stable.
Proof.
Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$. From Section 3.1, $(c)$, since $m(\sigma)=3\neq 7$, $\sigma$ is incompressible, and from [16, Theorem 1.1], we obtain $c_{3}\leq 4$ (for a more general formula, see 45). On the other hand, using Remark 7 we have $c_{3}({\mathcal{T}}_{\sigma})=6-2a$, so $a\in\{1,2,3\}$.
From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(3-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we get Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(3-a))=8-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=8-2a,$
which cannot coincide for the values $a\in\{1,2,3\}$, hence we obtain a contradiction. ∎
5 Pencils of cubics and degree 6 curves inside quadric surfaces
In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces.
The results are derived from the previous sections, Section 3.1 and Section 4, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [14], [21], [15], [3] and [16].
5.1 Pencils of cubics
In this section, we focus on the case of pencils of cubics, i.e., $d_{f}=d_{g}=2$. The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
$\operatorname{Bour}(\sigma)=e(e-4)+12-m(\sigma).$
When $e=1$, the formula above implies $m(\sigma)\leq 9$, from the inequality $\operatorname{Bour}(\sigma)\geq 0$.
The assumption $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ enables us to apply the result [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}(1)$ and obtain the following upper bounds for $m(\sigma)$ and the third Chern class $c_{3}({\mathcal{T}}_{\sigma})$:
Proposition 29.
If $\sigma=(f,g)$ is a normal pencil of cubics with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 8$.
Proposition 30.
If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then
  • (a)
    if $7\leq m(\sigma)\leq 8$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
  • (b)
    if $0\leq m(\sigma)<7$, then
    $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72.$
The following result is obtained using [15, Theorem 0.1] and the bound obtained in Section 3.1, $(a)$:
Proposition 31.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$, and we have the following table of sharper bounds for each possibility of $m=m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
6 $e\leq 3$
7 $e\leq 3$
8 $e\leq 2$
The following result gives a picture of the generic case of a pencil of cubics.
Proposition 32.
Let $\sigma=(f,g)$ be a general pencil of cubics in $\mathbb{P}^{3}$. Then $m(\sigma)=0$, the number of singular members is $32$, and all singular members have one singular point, in particular $m(\sigma)=0$. Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$. (see Example 11).
Proof.
This follows from intersection theory for the bundle of principal parts (see [10], Proposition $7.1$ and Proposition $7.4$). ∎
Rewriting Section 3.1, $(d)$ for pencils of cubics, we obtain:
Proposition 33.
Let $\sigma$ be a normal pencil of cubics. Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 34).
Example 34 (Free, compressible pencil of cubics).
Consider the sequence $\sigma=(x_{0}^{3}+x_{1}^{3}+x_{0}x_{1}x_{3},x_{0}x_{1}x_{3})$. This sequence is independent of the variable $x_{2}$, with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&3x_{1}^{2}+x_{0}x_{3}&0&x_{0}x_{1}\\ x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}\end{pmatrix}$
so that there are two linearly independent syzygies, one of degree zero and one of degree four, in the following matrix:
$\nu=\begin{pmatrix}0&-x_{0}x_{1}^{3}\\ 0&x_{0}^{3}x_{1}\\ 1&0\\ 0&-x_{0}^{3}x_{3}+x_{1}^{3}x_{3}\end{pmatrix}$
Here, $m(\sigma)=12$, the annihilator ideal and the 0-th Fitting ideal coincide, with the support being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
Using Section 4, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
Proposition 35.
Let $\sigma$ be a normal pencil of cubics. Then:
  • (a)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable;
  • (b)
    If $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 33, 36 and 38. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Section 4, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
$\operatorname{Bour}(\sigma)$ $=1-d+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$
$=1-4+4+4+4-m(\sigma)$
$=9-m(\sigma)>2,$
since $m(\sigma)\leq 6$. This contradicts the bound $\operatorname{Bour}(\sigma)\leq 2$ established earlier.
Moreover, for $(b)$, if we assume $m(\sigma)\leq 2$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then
$\operatorname{Bour}(\sigma)$ $=2-2d+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$
$=4-8+12-m(\sigma)$
$=8-m(\sigma)>5$
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Section 4, $(b)$. ∎
Proposition 36.
Let $\sigma=(f,g)$ be an incompressible normal pencil of cubics. Then $m(\sigma)\leq 9$, and $m(\sigma)=9$ if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$ (see Example 37).
Proof.
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 29 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
Example 37 (Free, unstable incompressible pencil of cubics ($m(\sigma)=9$)).
Consider the sequence $\sigma=(x_{1}(x_{2}^{2}-x_{1}^{2}),x_{3}x_{2}(x_{0}-x_{1}))$. Then the matrix $\nabla\sigma$ is given by:
$\nabla\sigma=\begin{pmatrix}0&-3x_{1}^{2}+x_{2}^{2}&2x_{1}x_{2}&0\\ x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\end{pmatrix}$
and it admits two linearly independent syzygies, one of degree one and one of degree $3$:
$\nu\doteq\begin{pmatrix}x_{0}-x_{1}&2x_{1}x_{2}^{2}\\ 0&2x_{1}x_{2}^{2}\\ 0&3x_{1}^{2}x_{2}-x_{2}^{3}\\ -x_{3}&-3x_{1}^{2}x_{3}+x_{2}^{2}x_{3}\end{pmatrix}$
Thus, we conclude ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, and in particular $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. In this case, $m(\sigma)=9$.
Proposition 38.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ if and only if $m(\sigma)=8$ (see Example 39).
Proof.
If $m(\sigma)=8$, by 31, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 2$. If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [3, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$. If we assume $e=1$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 7 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain $c_{3}({\mathcal{T}}_{\sigma})=4$, on the other hand, 30 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction and this case cannot happen.
On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$. ∎
Example 39 (Free, incompressible and $\mu$-semistable pencil of cubics ($m(\sigma)=8$)).
Consider the sequence $\sigma=(x_{0}^{2}x_{1}+x_{3}^{3},x_{0}^{3}+x_{0}x_{2}x_{3}+x_{3}^{3})$. The Jacobian matrix $\nabla\sigma$ is given by:
$\nabla\sigma=\begin{pmatrix}2x_{0}x_{1}&x_{0}^{2}&0&3x_{3}^{2}\\ 3x_{0}^{2}+x_{2}x_{3}&0&x_{0}x_{3}&x_{0}x_{2}+3x_{2}^{2}\end{pmatrix}$
and it admits two linearly independent syzygies of degree $2$:
$\nu\doteq\begin{pmatrix}-x_{0}x_{3}&-x_{0}x_{2}\\ 2x_{1}x_{3}&2x_{1}x_{2}-9x_{3}^{2}\\ 3x_{0}^{2}+x_{2}x_{3}&x_{2}^{2}-9x_{0}x_{3}\\ 0&3x_{0}^{2}\end{pmatrix}$
Thus, we conclude ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$.
Proposition 40.
Let $\sigma=(f,g)$ be a nearly free pencil of cubics. Then, the only possible discrete invariants are $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $m(\sigma)=7$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 41).
Proof.
Using Remark 7 for $p_{a}(B)=0$ and $\deg(B)=1$, we obtain
$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2e)$
$=-2+8-2e=2(3-e).$
On the other hand, if $m(\sigma)=7$, then $e=2$ and the formula for $c_{3}$ in Remark 7 yields $c_{3}({\mathcal{T}}_{\sigma})=2$. As we have observed in Example 39 and before, we must have $m(\sigma)\leq 7$ to be able to obtain $\operatorname{Bour}(\sigma)=1$. On the other hand, assuming $m(\sigma)\leq 6$, we obtain
$\operatorname{Bour}(\sigma)=12-m(\sigma)+e(e-4)\geq 6+e(e-4)\geq 2$
hence $\operatorname{Bour}(\sigma)\neq 1$. ∎
Example 41 (Nearly free pencil of Cubics).
We consider the following sequence of cubics:
$\sigma=(x_{0}^{2}(x_{1}-x_{2})+x_{2}^{2}(x_{1}-x_{0}+x_{3}),-x_{1}x_{2}x_{3}+x_{2}^{2}x_{3})$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$
Using Macaulay2 software, we compute $e=2$ and $m(\sigma)=7$, so that $\operatorname{Bour}(\sigma)=1$.
Proposition 42.
Let $\sigma=(f,g)$ be a normal pencil of cubics such that $m(\sigma)=7$. Then, the only possible $\mu$-semistable cases are:
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 40);
  • $e=3$ and $\sigma$ is locally free, $\operatorname{Bour}(\sigma)=2$ and $B$ is a pair of skew lines (or their degeneration) (see Example 24).
Proof.
For $e\geq 2$, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [3, Lemma 2.1], we conclude that the only possible cases are $c_{3}=0$ (stable case) or $c_{3}=2$ (strictly semistable case). Thus, for $e=2$, we must have $c_{3}({\mathcal{T}}_{\sigma})=2$, and thus we obtain the second case. For the third case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [22] fitting in a sequence of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}(2)\rightarrow\mathcal{I}_{B}(1)\rightarrow 0$
where $B$ is a pair of skew lines or their degeneration, thus $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=3$. ∎
As we mentioned in Remark 27, the only possible case with $m(\sigma)=7$ for which we do not have examples is when $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=8$. We summarize the results of this subsection in the following theorem:
Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 41), which is strictly $\mu$-semistable;
  • (d)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Item $(a)$ is 33 and item $(b)$ with 36 and Example 39. Item $(c)$ follows from 42 and item $(d)$ is 35. ∎
To finish this study, we consider a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
Example 43 ($m(\sigma)=4$, $e=2$, $\operatorname{Bour}(\sigma)=4$, $3$-syzygy).
Considering the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
$\sigma=(x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3},x_{0}^{3}+x_{1}^{3}+x_{2}x_{3}^{2}),$
with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}&3x_{1}^{2}&3x_{2}^{2}&3x_{3}^{2}\\ 3x_{0}^{2}&3x_{1}^{2}&x_{3}^{2}&2x_{2}x_{3}\end{pmatrix}$.
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$ with multiplicity $4$. The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
where the first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2,\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=4$. This is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
5.2 Degree $6$ curves inside quadric surfaces
Let us focus on the case of normal sequences $\sigma=(f,g)$ with $d_{f}=1,d_{g}=2$. Here, the Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ is given by:
$\operatorname{Bour}(\sigma)=e(e-3)+7-m(\sigma).$
Assuming $\sigma$ is incompressible, as $c_{1}({\mathcal{T}}_{\sigma})=-3$, we can apply [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}$ without further restrictions, and obtain the following upper bounds for $m(\sigma)$ and $c_{3}({\mathcal{T}}_{\sigma})$:
Proposition 44.
If $\sigma=(f,g)$ is an incompressible normal sequence with $d_{f}=1,d_{g}=2$, then $m(\sigma)\leq 5$.
Proposition 45.
Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then, the following hold:
  • (a)
    If $4\leq m(\sigma)\leq 5$, then $c_{3}\leq 5-m(\sigma))$. In particular, when $m(\sigma)=5$, $\sigma$ is locally free;
  • (b)
    If $0\leq m(\sigma)<4$, then $c_{3}\leq m(\sigma)^{2}-10m(\sigma)+25.$
From Section 3.1, $(d)$, we obtain:
Proposition 46.
Let $\sigma$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $\sigma$ is compressible if and only if $m(\sigma)=7$ (see Example 47).
Example 47 (Free and compressible sequence with $d_{f}=1,d_{g}=2$).
Considering the sequence
$\sigma=(x_{0}(x_{1}-x_{2}),x_{0}^{3}+x_{1}^{3}+x_{2}^{3}),$
which is independent of the variable $x_{3}$. The matrix
$\begin{pmatrix}0&-x_{0}(x_{1}^{2}+x_{2}^{2})\\ 0&x_{0}^{3}+x_{1}x_{2}^{2}-x_{2}^{3}\\ 0&x_{0}^{3}-x_{1}^{3}+x_{1}^{2}x_{2}\\ 1&0\end{pmatrix}$
gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
The following result is obtained using ([15], $0.1$):
Proposition 48.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$, and we have the following table of bounds for the following possibilities of $m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
4 $e\leq 2$
5 $e\leq 1$
As an easy consequence of Remark 7, we obtain:
Proposition 49.
Let $\sigma=(f,g)$ be a normal sequence which is not a pencil, (i.e., with $d_{f}\neq d_{g}$) such that $d$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
Proof.
Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore
$2p_{a}(B)$ $=2-\operatorname{Bour}(\sigma)(4+d-2e)$
$p_{a}(B)$ $=1-\operatorname{Bour}(\sigma)(2+\frac{d}{2}-e)=1-(2-e)\operatorname{Bour}(\sigma)+\frac{d\operatorname{Bour}(\sigma)}{2},$
which is not an integer if $\operatorname{Bour}(\sigma)$ is odd, so we get a contradiction. ∎
Proposition 50.
Let $\sigma$ be an incompressible sequence with $d_{f}=1,d_{g}=2$ such that $m(\sigma)$ is even, that is, $m(\sigma)\in\{0,2,4\}$. Then $\sigma$ is not locally free.
Proof.
To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 49.
From 4, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
Assuming $e=2k$ is even, we obtain
$e^{2}-3e=4k^{2}-6k=2(2k^{2}-3k),$
an even number. On the other hand, when $e=2k+1$ is odd, then
$e^{2}-3e=4k^{2}+4k+1-6k-3=4k^{2}-2k-2=2(2k^{2}-k-1),$
which is also even. ∎
Proposition 51.
Let $\sigma=(f,g)$ be an incompressible normal sequence with degrees $d_{f}=1$, $d_{g}=2$. Then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ if and only if $m(\sigma)=5$ (see Example 52).
Proof.
Assuming $\sigma$ is non-compressible, by 48, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
Example 52 (Free and incompressible sequence, $m(\sigma)=5$).
Considering the sequence
$\sigma=(x_{0}x_{1},x_{3}x_{2}(x_{0}-x_{1})),$
of arrangements of hyperplanes, with Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&0&0\\ x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\end{pmatrix}$
The matrix
$\begin{pmatrix}0&x_{0}(x_{0}-x_{1})\\ 0&-x_{1}(x_{0}-x_{1})\\ x_{2}&0\\ -x_{3}&-x_{3}(x_{0}+x_{1})\end{pmatrix}$
gives linearly-independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
Proposition 53.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=4$, then $\sigma$ is nearly free, and we have two possible cases:
  • (a)
    ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=1$ (see Example 54);
  • (b)
    ${\mathcal{T}}_{\sigma}$ is unstable with $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 15);
Furthermore, these are the only two possibilities of numerical invariants for nearly free sequences with $d_{f}=1,d_{g}=2$.
Proof.
Using 48, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$. Using Remark 7 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples. To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
$\operatorname{Bour}(\sigma)=7-m(\sigma)+e(e-3)\geq 4+e(e-3)\geq 3$
for $e\in\{1,2,3\}$. ∎
Example 54 (Nearly free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
We consider the following normal sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(x_{0}x_{1}-x_{2}x_{3},x_{1}x_{3}(x_{0}-x_{2})),$
with Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&-x_{3}&-x_{2}\\ x_{1}x_{3}&x_{3}(x_{0}-x_{2})&-x_{1}x_{3}&x_{1}(x_{0}-x_{2})\end{pmatrix}$.
In this case, we have equality of schemes $\Xi_{\sigma}=\operatorname{supp}(\mathcal{Q}_{\sigma})$. The Jacobian scheme has a structure of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$, which is outside the three lines, and therefore $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$. The free resolution of ${\mathcal{T}}_{\sigma}$ is (obtained computationally):
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\xrightarrow{M}{\mathcal{T}}_{\sigma}\rightarrow 0$
given by matrices
$M=\begin{pmatrix}x_{0}x_{1}+x_{3}(x_{2}-x_{0})&x_{0}^{2}&x_{0}x_{3}\\ -x_{1}^{2}+x_{1}x_{3}&-x_{0}x_{1}&0\\ x_{1}x_{2}&x_{2}^{2}&x_{0}x_{1}+x_{3}(x_{2}-x_{1})\\ 0&-x_{2}x_{3}&x_{1}x_{3}-x_{3}^{2}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}\\ x_{3}-x_{1}\\ -x_{2}\end{pmatrix}$.
So we obtain $m(\sigma)=4,e=2$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=1$ corresponding to the point $p$, which is an irreducible component of $\Xi_{\sigma}$ of codimension three.
Proposition 55.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$. We have examples for the cases $c_{3}=2$ (see Example 56) and $c_{3}=4$ (see Example 57).
Proof.
First, the bound $c_{3}\leq 4$ is obtained from 45 for $m=m(\sigma)=3$. As we shown in 28, $e\geq 2$. We claim that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ in this case.
We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$. Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$. This follows from [21, Proposition 1.1] for $c_{3}=0$, [3, Lemma 2.4] for $c_{3}=2$ and [14, Lemma 9.6] for $c_{3}=4$. ∎
Example 56 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=2$).
We consider the following sequence:
$\sigma=(x_{3}(x_{0}-x_{1}),x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3})$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$.
Here, the saturation of the annihilator ideal coincides with the saturation of the 0-th Fitting ideal, and the first Fitting ideal has codimension three. The support of $\mathcal{Q}_{\sigma}$ in codimension two consists of two lines $V(x_{3},x_{0})\cup V(x_{3},x_{0}-x_{1})$, and the first one has multiplicity two. Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
so we get $\operatorname{Bour}(\sigma)=2,m(\sigma)=3$ and $c_{3}({\mathcal{T}}_{\sigma})=2$.
Example 57 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=4$).
We consider the following sequence:
$\sigma=(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2},x_{3}(x_{2}-x_{3})(x_{0}-x_{1}))$
with corresponding Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{2}x_{3}-x_{3}^{2}&x_{3}^{2}-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})+2x_{3}(x_{1}-x_{0})\end{pmatrix}$.
Here, the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $R$. The codimension two locus of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of three simple lines. Thus, $m(\sigma)=3$. Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $\operatorname{Bour}(\sigma)=2$ and $c_{3}({\mathcal{T}}_{\sigma})=4$.
Corollary 58.
Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is stable if and only if $e>1$. From the result above, the only possibilities of $\operatorname{Bour}(\sigma)\leq 2$ are when $m(\sigma)\geq 3$, and moreover if $m(\sigma)=3$ we cannot have $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Section 4, $(a)$, $e\neq 1$. ∎
Proposition 59.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)=2$, then $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, $c_{3}\in\{1,3,5,7\}$ and $c_{3}=7$ if and only if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, with $\operatorname{Bour}(\sigma)=3$ and $p_{a}(B)=0$ (see Example 60).
Proof.
If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [14, Section 9]. By the proof of [14, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
From Remark 7, we obtain $p_{a}(B)=1$, and thus $B$ must be a plane cubic curve, with a resolution of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
which by Lemma 13 yields a resolution
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [14, Lemma $9.3$]. If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7). From [3, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$. ∎
Example 60 ($m(\sigma)=2,e=2$, $\operatorname{Bour}(\sigma)=3$ , $c_{3}=7$).
Consider the following sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(-x_{0}x_{1}+x_{1}x_{2}-x_{2}x_{3},x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3}),$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}-x_{1}&x_{2}-x_{0}&x_{1}-x_{3}&-x_{2}\\ x_{1}^{2}&2x_{0}x_{1}&3x_{2}^{2}+2x_{2}x_{3}&x_{2}^{2}\end{pmatrix}$.
Here the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the codimension two part of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of a line $V(x_{1},x_{2})$ with multiplicity two. The saturation of the first Fitting ideal is $(0)$ in this case. The sheaf ${\mathcal{T}}_{\sigma}$ admits the following free resolution:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $e=2$, $\operatorname{Bour}(\sigma)=3$ and $m(\sigma)=2$, with $c_{3}({\mathcal{T}}_{\sigma})=7$.
Example 61 ($m(\sigma)=2$, $e=3$, $\operatorname{Bour}(\sigma)=5$, $c_{3}=3$).
We consider the following sequence:
$\sigma=(x_{2}x_{3}-x_{0}x_{1},x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{2}x_{3}^{2}+x_{3}^{3})$
with corresponding Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}-x_{1}&-x_{0}&x_{3}&x_{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}+x_{3}^{2}&x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2}\end{pmatrix}$.
Here, the saturation of the 0-th Fitting ideal coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $(1)$. Moreover, $m(\sigma)=2$ with $\operatorname{supp}(\mathcal{Q}_{\sigma})$ being a double line structure at $V(x_{0},x_{3})$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 3}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we obtain $\operatorname{Bour}(\sigma)=5$ and $c_{3}({\mathcal{T}}_{\sigma})=3$.
We summarize the results of this subsection in the following theorem:
Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 54 and Example 15);
  • (d)
    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Item $(a)$ is 46 and item $(b)$ follows with 51. Item $(c)$ is shown in 53, and the stability result in $(d)$ is in 58. ∎

### v3

@@
 Abstract
-The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials - defined as the kernel of the associated Jacobian matrix - naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation.
+The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials — defined as the kernel of the associated Jacobian matrix — naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation.
 As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
@@
 We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis.
-The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation.
+The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation of the degrees.
 We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants.
 Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
-We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly-free cases.
-In particular, one of the nearly-free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
+We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly free cases.
+In particular, one of the nearly free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
 1 Introduction
@@
 The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$.
-It is considered (see [11]) in analogy with the case of the tangent logarithmic sheaf associated to hypersurfaces in $\mathbb{P}^{n}$.
+This definition is in analogy (see [undefj]) with the case of the tangent logarithmic sheaf associated to divisors in $\mathbb{P}^{n}$, namely for $k=1$.
 A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
-In the case $k=1$ of divisors on $\mathbb{P}^{n}$, one has a short exact sequence of the form:
+In the case of divisors on $\mathbb{P}^{n}$, where $k=1$, one has a short exact sequence of the form:
 $0\rightarrow{\mathcal{T}}_{f}\rightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\mathcal{I}_{J_{f}}(d)\rightarrow 0,$
@@
 For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated to the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
-One should consider sequences rather than the associated subvarieties $X=V(\sigma)$ if $k>1$, because of the following observation: different choices of sequences $\sigma^{\prime},\sigma$ generating the same ideal may have different Jacobian syzygy modules (see [11, Example 2.7]).
-When $k=1$, however, any different choice $f^{\prime}$ for generator of the ideal $\langle f\rangle$ gives a linear multiple of $\nabla f$, so that ${\mathcal{T}}_{f}\simeq{\mathcal{T}}_{f^{\prime}}$.
-Similarly, when $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and in particular ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [11, Lemma 2.14]).
-We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$.
-Following the terminology of [11], we say that a sequence $\sigma=(f_{1},\ldots,f_{k})$ is compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$.
+For the generalization of $k>1$, one should consider sequences rather than the associated ideals $I=(f_{1},\ldots,f_{k})$, as pointed out by [undefj, Example 2.7].
+When $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and furthermore ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [undefj, Lemma 2.14]).
+We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$, called the initial degree.
+Following the terminology of [undefj], we call a sequence $\sigma=(f_{1},\ldots,f_{k})$ compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$.
 This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$.
-Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [11, Lemma 2.8]).
-One may also consider the relationship with Bourbaki ideals, as in [17] and [7].
+Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [undefj, Lemma 2.8]).
+One may also consider the relationship with Bourbaki ideals, as in [undefr] and [undeff].
 For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{n}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B_{\nu}}(e-d)\rightarrow 0$
-where $d=-c_{1}({\mathcal{T}}_{\sigma})$ and $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection, since ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank two.
+where $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection.
 The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice.
-In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [17].
+In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [undefr].
 By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence.
-The focus of this work is to understand the above concept for $n=3$ and $k=2$.
+The focus of this work is to understand the concept above for $n=3$ and $k=2$.
 For $k=1$, the singular scheme of $V(f)$, defined by the Jacobian ideal $\Sigma_{f}\doteq V(J_{f})\subset\mathbb{P}^{n}$, plays an important role in this study.
@@
 Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$.
-For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [17]):
+For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [undefr]):
 $\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}),$
 where $e=\operatorname{indeg}({\mathcal{T}}_{f})$.
-Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for ${\mathcal{T}}_{\sigma}$ to get an ideal sheaf as the cokernel.
-This problem is addressed in the recent work [7] where the authors consider tame hypersurfaces and other notions of Bourbaki degrees.
-In this paper, we focus on the case of sequences $\sigma=(f,g)$ on $\mathbb{P}^{3}$, so our logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has also rank two, and we consider the same approach as in [17], since we only have to choose one syzygy to do the construction.
-We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$.
-In [11], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$.
-In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$, and $c_{1}(\mathcal{Q}_{\sigma})$ coincides with the degree of the greatest common divisor among all $(k\times k)$-minors of $\nabla\sigma$.
-The generic case is when $c_{1}(\mathcal{Q}_{\sigma})=0$, and this $g.c.d.$ is one, in which case we call $\sigma$ a normal sequence (see 2.1).
-If we assume $c_{1}(\mathcal{Q}_{\sigma})=0$, we obtain a Hilbert polynomial of the form
-$H(\mathcal{Q}_{\sigma},t)=\frac{\deg(\mathcal{Q}_{\sigma})}{s!}t^{s}+O(t^{s-1}),$
-where $s\leq n-2$. We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ has codimension at least three.
-In the case $n=3$ and $k=2$, we have sequences $\sigma=(f,g)$ of homogeneous polynomials with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, setting $d\doteq d_{f}+d_{g}$, so our objects are pairs of projective surfaces.
-We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and thus $c_{1}({\mathcal{T}}_{\sigma})=-d$.
-Since ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
-Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we may reproduce analogous considerations as in [17]. We show a formula for the Bourbaki degree of $\sigma$ given by
+Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for the logarithmic sheaf ${\mathcal{T}}_{f}$ to get an ideal sheaf as the cokernel.
+This problem is addressed in a recent work [undeff] where the authors consider another kind of Bourbaki sequences, recently developed in [undefq], by choosing generators of maximal degree.
+We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$.
+In [undefj], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$.
+In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$.
+The generic case is when $\dim(\mathcal{Q}_{\sigma})\leq n-2$, and in particular $c_{1}(\mathcal{Q}_{\sigma})=0$.
+In this case, we call $\sigma$ a normal sequence.
+For now on, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$.
+We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$.
+Since the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
+If $\sigma$ is a normal sequence, $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
+$H(\mathcal{Q}_{\sigma},t)=\deg(\mathcal{Q}_{\sigma})t+O(1),$
+We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional.
+From the associativity formula applied to the graded homogeneous module $\mathcal{Q}$ associated to $\mathcal{Q}_{\sigma}$, we obtain the expression (see, for example, [undefs, Theorem $14.7$])
+$m(\sigma)=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(R/P)=2}\operatorname{{length}}_{R_{P}}\mathcal{Q}_{P}\cdot\deg(R/P).$
+Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [undefr]. We show a formula for the Bourbaki degree of $\sigma$ given by
 $\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma),$
-where $m(\sigma)=\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$ is the degree of the sheaf $\mathcal{Q}_{\sigma}$ as defined above.
-The discrete invariant $m(\sigma)$ plays the role of the Tjurina number in the previous context.
-We compare the different scheme structures between $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ in Example 2, considering a free sequence $\sigma=(f,g)$ of polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with $d_{f}=1,d_{g}=2$ and so that $m(\sigma)=5$, but $\deg(\Xi_{\sigma})=6$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))=4$ as non-reduced schemes, showing the relationship between $\deg(\mathcal{Q}_{\sigma})$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))$ is more complicated than in the case $k=1$, where these two coincide for the sheaf $\mathcal{O}_{J_{f}}(d)$.
-About the initial degree and the invariant $m=m(\sigma)$, we show some bounds that are analogous to the case of plane projective curves:
-Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
+so that $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. A first result of the paper is some bounds for the discrete quantities defined above.
+Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
   • (a)
-    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
+    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d$;
   • (b)
@@
       $\sigma$ is compressible;
-We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 3.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 9): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable.
-We also show the bound $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, which is attained for example for regular pencils with singular members having only isolated singularities (see Remark 10).
+We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 8): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable.
+In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely by the matrices $\nabla\sigma$ for which the Buchsbaum–Rim complex is a minimal resolution, or equivalently $m(\sigma)=0$.
+Geometrically, all regular pencils ($d_{f}=d_{g}$) with isolated singularities satisfy this property (see Remark 9).
 In Section 2.2, we relate our notion to the notion of the Bourbaki degree for a projective plane curve, by considering a reduced polynomial $g\in\kappa[x_{0},x_{1},x_{2}]$ together with its associated surface $V(g)\subset\mathbb{P}^{3}$ and projective curve $X=V(g)\subset\mathbb{P}^{2}$, to obtain $\operatorname{Bour}(X)=\operatorname{Bour}(g,x_{3})$.
-Afterwards, in Section 3.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is a Bourbaki scheme obtained by the choice of a minimal syzygy.
-Related to the free resolutions, we introduce the notions of nearly-free sequences and $3$-syzygy sequences.
-We have the following chain of implications:
-$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3-\text{syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
-and the converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$.
-In Section 4, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [11, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections.
-For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [17, Corollary 2.11]).
+Afterwards, in Section 2.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal syzygy.
+Related to the free resolutions, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
+$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1.$
+The converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$. We point out nearly free sequences are never locally free, as we show in 19.
+In Section 3, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [undefj, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections.
+For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [undefr, Corollary 2.11]).
 The main theorem of the section is:
-Theorem B. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
+Theorem B.(Theorem B) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
   • (a)
@@
 In Section 4, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$.
-In these two classes, we classify all free and nearly free cases in terms of their discrete invariants and establish some stability results, which are summarized below:
-Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
+To do so, we use classical results on bounds for Chern classes of reflexive sheaves in $\mathbb{P}^{3}$.
+We classify all free and nearly free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized below:
+Theorem C.(Theorem C)Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
   • (a)
@@
     If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
-In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 16), but the bound above for stability may not be sharp.
-Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
+In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 15), but the bound given above for stability may not be sharp.
+Theorem D.(Theorem D) Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
   • (a)
@@
   • (d)
-    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
+    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$.
+Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
+  • (e)
+    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
 Example 14 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable.
-This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [2]:
+This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [undefa]:
 Conjecture. If the tangent sheaf of a codimension one foliation on $\mathbb{P}^{3}$ is not split, then it is $\mu$-semistable.
-Along the text, we describe examples developed computationally with aid of Macaulay2 software ([12]) for the two families of sequences.
+Along the text, we describe examples developed computationally with aid of Macaulay2 software ([undefl]) for the two families of sequences.
 Acknowledgments
@@
 student under the advice of Marcos Jardim and Daniele Faenzi, with partial funding by the Bridges ”Brazil-France interplays in Gauge Theory, extremal structures and stability” projects ANR-21-CE40-0017 and ANR-17-EURE-0002.
-We thank Marcos Jardim, Daniele Faenzi and Alan Muniz for fruitful discussions and suggestions regarding this work.
-2 Basic setting and $m$-syzygy sequences on $\mathbb{P}^{n}$
-We work over an algebraically closed field $\kappa$ of characteristic zero, although most of the definitions can be considered under relaxed assumptions.
-We denote by $R=\kappa[x_{0},\ldots,x_{n}]$ the graded ring of polynomials over $\kappa$ and $\mathbb{P}^{n}\doteq\mathbb{P}^{n}_{\kappa}=\operatorname{Proj}R$ the projective $n$ space over $\kappa$.
-In this section, we recall from [11] the definition of the logarithmic tangent sheaf associated to a sequence of $k$ polynomials, and introduce some notation inspired by the literature about the case $k=1$.
-For a sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials with degrees $d_{1}+1,\ldots,d_{k}+1$, we consider the Jacobian matrix $\nabla\sigma$ as a morphism between locally free sheaves:
-$\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\xrightarrow{\nabla\sigma}\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i})$
-so the kernel sheaf ${\mathcal{T}}_{\sigma}\doteq\ker(\nabla\sigma)$, is called the logarithmic tangent sheaf associated to $\sigma$.
-We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel sheaf and by $\mathcal{M}_{\sigma}\doteq\operatorname{im}(\nabla\sigma)$ the image sheaf.
-We denote by $d\doteq\sum_{i=1}^{n}d_{i}$ the total degree of a sequence $\sigma$.
-Assuming the generic rank of $\operatorname{im}(\nabla\sigma)$ is $k$, it follows that ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank n+1kn+1-k.
-2.1 Normal sequences
-By construction, the reduced support of the sheaf $\mathcal{Q}\sigma$ coincides with the vanishing locus $\Xi\sigma\doteq V(\bigwedge^{k}\nabla\sigma)$ of the $k\times k$-minors of the matrix $\nabla\sigma$.
-Assuming $\sigma$ is algebraically independent, the generic rank of $\nabla\sigma$ is $k$, so that $\operatorname{{codim}}\Xi_{\sigma}\geq 1$.
-Moreover, since $\mathcal{Q}\sigma$ is a torsion sheaf, the condition $c_{1}(\mathcal{Q}\sigma)=0$ is equivalent to $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$.
-In terms of the matrix $\nabla\sigma$, $\operatorname{{codim}}\Xi_{\sigma}\geq 2$ if and only if the greatest common divisor among all $(k\times k)$-minors is one.
-When $k=1$, we have
-$\operatorname{{codim}}(Z_{f})=\operatorname{{codim}}(\mathcal{Q}_{f})\geq 2\iff V(f)\text{ is normal}.$
-Inspired by the behavior above, we call a sequence $\sigma$ normal when $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$.
-We note that the two schemes $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ may have different scheme structures and, in particular, different degrees.
-On one hand, the scheme structure of the support $\operatorname{supp}(\mathcal{Q}_{\sigma})=V(\operatorname{Ann}\mathcal{Q}_{\sigma})$ is induced by the annihilator ideal sheaf of $\mathcal{Q}_{\sigma}$, defined locally as the annihilator ideal of the associated module.
-On the other hand, the scheme structure of $\Xi_{\sigma}$ is induced locally by the 0-th Fitting ideal of the $\mathcal{Q}_{\sigma}$.
-These schemes have the same reduced locus, since the support of these two ideals coincide, but in general $\operatorname{Fitt}_{0}\subseteq\operatorname{Ann}$ (see, for example, [9, Proposition 20.7]).
-For $k=1$, when $\sigma=f$, we have an isomorphism $\mathcal{Q}f\simeq\mathcal{O}{Z_{f}}$.
-This holds because $\mathcal{M}f\simeq\mathcal{I}{Z_{f}}(d)$ is the ideal sheaf of the scheme $Z_{f}$, and whenever $f$ is reduced, the schemes $\Xi_{f}$ and $\operatorname{supp}(\mathcal{Q}{\sigma})$ coincide.
-If $\sigma$ is a normal sequence, then the Hilbert polynomial of the cokernel sheaf $\mathcal{Q}_{\sigma}$ is of the form
-$\operatorname{Hilb}(\mathcal{Q}_{\sigma},t)=\frac{m(\sigma)}{(n-2)!}t^{n-2}+\ldots$
-where $m(\sigma)\geq 0$ is a discrete invariant so that $m(\sigma)=0$ if and only if $\operatorname{{codim}}(\mathcal{Q}_{\sigma})\geq 3$.
-However, $m(\sigma)$ generally coincides with neither the schematic degree of $\Xi_{\sigma}$ nor that of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, as we explore in Example 2 at the end of this section.
-2.2 $m$-syzygy sequences and exponents
-We may consider a minimal free resolution of the sheaf ${\mathcal{T}}_{\sigma}$, which will be given in the following form:
-$\ldots\rightarrow\bigoplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{n}}(-e_{i})\xrightarrow{\rho}{\mathcal{T}}_{\sigma}\rightarrow 0.$
-In which case, we say $\sigma$ is an $m$-syzygy sequence.
-We note that $m\geq n$ and $\sigma$ is said to be free if and only if ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, or, equivalently, when $m=n$.
-The integers $e_{1}\leq\ldots\leq e_{m}$ will be called exponents, and the first $\operatorname{indeg}({\mathcal{T}}_{\sigma})\doteq e_{1}=e$ will be called the initial degree of the sheaf ${\mathcal{T}}_{\sigma}$.
-This is the minimum integer $e\geq 0$ such that there is a non-trivial syzygy $\nu$ of degree $d$ for the Jacobian matrix $\nabla\sigma$.
-Since there is an injection ${\mathcal{T}}_{\sigma}\hookrightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}$ and the latter is a $\mu$-semistable sheaf, all exponents satisfy $e_{i}\geq 0$.
-A sequence $\sigma=(f_{1},\ldots,f_{k})$ is called compressible if, after a linear change of variables, there is a variable that does not occur in any of the $f_{1},\ldots,f_{k}$.
-The number of variables omitted in this way will give rise to syzygies of degree zero for $\nabla\sigma$, so the following hold (for proofs, see [11, Lemma 2.7 and Lemma 2.8]):
+We thank Marcos Jardim, Daniele Faenzi, Alan Muniz, Victor Cordeiro and Abbas Nejad for fruitful discussions and suggestions regarding this work.
+2 The Bourbaki degree of pairs of projective surfaces
+In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials.
+We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3), in particular introducing the class of nearly free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
+2.1 Framework and first results
+By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
+By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
+Each sequence $\sigma=(f,g)$ induces a morphism of sheaves on $\mathbb{P}^{3}$ by the Jacobian matrix:
+$\nabla\sigma:\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(d_{g}),$
+and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$.
+We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [undefj].
+The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
 Lemma 1.
-The following quantities coincide:
-  • The number of variables omitted in $f_{1},\ldots,f_{k}$;
-  • The discrete invariant $h^{0}({\mathcal{T}}_{\sigma})$;
-  • The number of 0’s in a vector of exponents for ${\mathcal{T}}_{\sigma}$.
-In particular, $e=0$ if and only if $\sigma$ is compressible.
-For $n=3$ and when $\sigma=(f,g)$ is a normal sequence, we show the conditions above are equivalent to $m(\sigma)$ attaining its maximum value, see Section 3.1.
+Let $\mathcal{Q}$ be a coherent sheaf on $\mathbb{P}^{3}$ with $\operatorname{{rk}}(\mathcal{Q})=0$ and $c_{1}(\mathcal{Q})=0$. Then the Hilbert polynomial of $\mathcal{Q}$ is given by
+$\mathcal{X}(\mathcal{Q}(t))=\operatorname{ch}_{2}(\mathcal{Q})t+\operatorname{ch}_{3}(\mathcal{Q})+2\operatorname{ch}_{2}(\mathcal{Q}).$
+Proof.
+Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
+We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$.
+From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
+As in the introduction, the Jacobian scheme of $\nabla\sigma$ is defined as
+$\Xi_{\sigma}\doteq V\left(\bigwedge^{2}\nabla\sigma\right)=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
+the zero-locus of the $2\times 2$ minors of $\nabla\sigma$ or, in other words, the $0$-th Fitting scheme of $\nabla\sigma$.
+From the general theory of Fitting ideals (see, for example, [undefh, Chapter 20]), the annihilator ideal of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal, and they have the same support, namely
+$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
+as closed subsets, and therefore the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$.
+In general, however, their schematic structure may be different, as we explore in the following example.
+Denote by $(\Xi_{\sigma})_{1}$ the one-dimensional part of the scheme $\Xi_{\sigma}$.
 Example 2.
@@
 so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively.
-However, from the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ we conclude $m(\sigma)=5$.
-3 The Bourbaki degree of pairs of projective surfaces
-In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials.
-We start with definitions and first results (3.1), followed by a reduction to the case of a projective plane curve (3.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (3.3), in particular introducing the class of nearly-free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
-3.1 Framework and first results
-By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
-We also work only with normal sequences, as defined previously in 2.1.
-By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
-Each sequence $\sigma=(f,g)$ induces a morphism of sheaves on $\mathbb{P}^{3}$ by the Jacobian matrix:
-$\nabla\sigma:\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(d_{g}),$
-and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$.
-We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [11].
-The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
-Lemma 3.
-Let $\mathcal{Q}$ be a coherent sheaf on $\mathbb{P}^{3}$ with $\operatorname{{rk}}(\mathcal{Q})=0$ and $c_{1}(\mathcal{Q})=0$. Then the Hilbert polynomial of $\mathcal{Q}$ is given by
-$\mathcal{X}(\mathcal{Q}(t))=\operatorname{ch}_{2}(\mathcal{Q})t+\operatorname{ch}_{3}(\mathcal{Q})+2\operatorname{ch}_{2}(\mathcal{Q}).$
-Proof.
-Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
-We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$.
-From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
+From the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$, since ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ we conclude $m(\sigma)=5$.
+The right approach to computing $m(\sigma)$ in general is using the associativity formula, as we will in the next examples.
 Any saturated syzygy of the Jacobian matrix $\nu$, of degree $e\in\mathbb{Z}$, induces a short exact sequence:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
-where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described by the Serre correspondence (see [14, Theorem $4.1$]).
+where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described as the Serre correspondence (see [undefn, Theorem $4.1$]) between the sheaf ${\mathcal{T}}_{\sigma}$ and the scheme $B$.
 Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$.
@@
 $\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
-We define the Bourbaki degree of a sequence $\sigma$, inspired by [17, Definition 2.4].
+We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [undefr, Definition 2.4].
 Definition 4.
@@
 Remark 5.
-It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$.
+It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$, since every short exact sequence of the form
+$0\to\mathcal{O}_{\mathbb{P}^{3}}(-e)\to{\mathcal{T}}_{\sigma}\to\mathcal{O}_{\mathbb{P}^{3}}(e-d)\to 0$
+splits.
 Remark 6.
-From [13, Proposition 4.1], we have the formula
+From [undefm, Proposition 4.1], we have the formula
 $c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
@@
 Now, we obtain some bounds for the quantities $\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ in terms of the degrees $d_{f},d_{g}$.
-Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
+Theorem A.
+Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
   • (a)
@@
 so that $m(\sigma)\leq m_{0}+e(e-d)$, but $e(e-d)\leq 0$ and the claim follows.
-For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [11, Lemma 2.7].
-Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [11, Lemma 2.8].
+For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [undefj, Lemma 2.7].
+Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [undefj, Lemma 2.8].
 The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
@@
 Remark 7.
-The bound obtained in 3.1, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory.
-We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [11, Section 9]).
+The bound obtained in Theorem A, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory.
+We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [undefj, Section 9]).
 Let $C$ be the one-dimensional part of the singular scheme of this foliation.
-From the formulas of discrete invariants in [2, Theorem 3.1],
+From the formulas of discrete invariants in [undefa, Theorem 3.1] given for generic foliations,
 $c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C).$
@@
 $\deg(C)-d-d_{f}d_{g}-1=m(\sigma)$ $\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g},$
-and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [20, Corollary 4.8].
+and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [undefv, Corollary 4.8].
 From simple observations about the formula $\operatorname{Bour}(\sigma)$, we are able to obtain the following inequalities related to $\mu$-stability and freeness of the logarithmic sheaves.
@@
   • (a)
-    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if and only if $m(\sigma)=0$ and $e=d$.
+    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if any of the equivalent facts hold:
+    • (1)
+      $m(\sigma)=0$;
+    • (2)
+      A minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
+      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{1})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{2})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
+      a Buchsbaum–Rim resolution, where
+      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\phi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
+      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$. In particular, $e=d$.
   • (b)
@@
 This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility.
-The converse also follows simply from the formula.
+The converse also follows simply from the formula, assuming that $m(\sigma)=0$ and $e=d$.
+To show the equivalences above, we start by noting that $m(\sigma)=0$ implies that the grade of the ideal of $(2\times 2)$-minors is three, and therefore the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$.
+From this fact, we obtain that $e=d$.
+For the converse, since $e=d$ and the Buchsbaum–Rim complex is a minimal resolution, we obtain that the grade of the ideal of minors of $\nabla\sigma$ is at least three, hence $\dim(\mathcal{Q}_{\sigma})=0$ and thus $m(\sigma)=0$.
 By construction, $m(\sigma)\geq 0$, and therefore for a given value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, the Bourbaki degree of $\sigma$ can be at most
@@
 Remark 9.
-If $\sigma=(f,g)$ is a pencil of surfaces of the same degree, say $d_{f}=d_{g}=p$, then $m(\sigma)=0$ if $\sigma$ is a regular pencil and, moreover, every singular member $V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}$ is normal.
-This follows from the description
-$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g))$
-obtained in [11, Lemma 2.17], since normal projective surfaces have isolated singularities and regular sequences have only a finite number of singular members, therefore $\dim(\operatorname{supp}(\mathcal{Q}_{\sigma}))=0$.
+Let $\sigma=(f,g)$ be a sequence of surfaces of the same degree, say $d_{f}=d_{g}=p$, called a pencil of surfaces. It defines a 11-parameter family of projective surfaces as:
+$V_{\sigma}\doteq\{V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}:z=[z_{0}:z_{1}]\in\mathbb{P}^{1}\},$
+and each hypersurface $V(z_{0}f+z_{1}g)$ is called a member of the pencil $V_{\sigma}$.
+The set where each surface is singular is the vanishing locus of a polynomial $p=p(z)$ (the discriminant) inside $\mathbb{P}^{1}$, so it must be either a finite subset or the whole family.
+When this set is finite, the pencil is called a regular pencil, otherwise is called irregular.
+In [undefj, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}$ for pencils of surfaces as the union of singular loci:
+$(\Xi_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g)).$
+Hence, whenever a normal sequence $\sigma=(f,g)$ is a regular pencil where every singular member has isolated singularities, we conclude $\dim(\Xi_{\sigma})=0$ from the description above, and thus $m(\sigma)=0$.
 Example 10.
-We consider the sequence
+Consider the sequence
 $\sigma=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
 where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$.
-Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a locally free resolution of the form:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-Remark 12.
-We have not found any examples of sequences with $m(\sigma)=0$ so that $e<d$, which motivates the conjecture: does $m(\sigma)=0$ implies $e=d$?
-3.2 The Bourbaki degree of a plane curve from the Bourbaki degree of a pair of surfaces
-In this section, we obtain the formula of the Bourbaki degree of a projective plane curve $X=V(g)\subset\mathbb{P}^{2}$, where $g\in\kappa[x_{0},x_{1},x_{2}]$ is a reduced homogeneous polynomial, introduced in [17], as a special case of our formula, by considering the pair $\sigma=(x_{3},g)$ in $\mathbb{P}^{3}$.
+Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
+given by the Buchsbaum–Rim complex.
+2.2 Plane curves as pairs of surfaces
 Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$.
 We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$.
+Denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme, given as the zero locus of the Jacobian ideal of $g$.
 The matrix $\nabla\sigma$ will be given by
 $\nabla\sigma=\begin{pmatrix}0&0&0&1\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&0\end{pmatrix}$.
-Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ the vector, so we denote by
+Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ and
 ${\mathcal{T}}_{\overline{g}}\doteq\ker(\nabla\overline{g})\hookrightarrow\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}\xrightarrow{\nabla\overline{g}}\mathcal{O}_{\mathbb{P}^{3}}(d_{g})$
-the kernel of the multiplication by this vector. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
+the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
 ${\mathcal{T}}_{\overline{g}}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}$$\mathcal{O}_{\mathbb{P}^{3}}(d)$$\mathcal{O}_{Z_{g}}(d)$${\mathcal{T}}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}$$\mathcal{O}_{\mathbb{P}^{3}}(d)\oplus\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{Q}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{O}_{\mathbb{P}^{3}}$$\scriptstyle{\nabla\overline{g}}$$\scriptstyle{\nabla\sigma}$$\scriptstyle{\cdot 1}$
 and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$.
+Denoting $i:H\hookrightarrow\mathbb{P}^{2}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely.
 From the discrete invariants considered $(d_{f}=0)$, we have the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
-$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}(\ker\nabla g(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
+$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}({\mathcal{T}}_{g}(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
 where $\tau(X)$ is the Tjurina number of the curve $X$, since $d_{f}=0$.
-From the construction, we have that the singular scheme of ${\mathcal{T}}_{\sigma}$ (described by $Z_{g}$) intersects $V(x_{3})$ transversely, so we obtain the formula between Chern classes $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:\mathbb{P}^{2}\simeq V(x_{3})\hookrightarrow\mathbb{P}^{3}$ is the inclusion.
-Therefore,
+From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:H\hookrightarrow\mathbb{P}^{3}$ and thus
 $1-d[H]+(m_{0}-m(\sigma))[H]^{2}$ $=i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$
@@
 and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$.
-Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number, and thus it should be the correct generalization for the case of pairs of surfaces.
-3.3 Locally free resolutions and the Bourbaki scheme
-In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [17, Theorem 2.1, (c)].
-This follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$.
+Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does.
+A more algebraic way to see this is through the associativity formula.
+Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$, with the effect of adding a variable to the graded ring of the plane $\mathbb{P}^{2}\hookrightarrow\mathbb{P}^{3}$, and the associated map of modules is given by the quotient map
+$0\to(x_{3})\to\kappa[x_{0},x_{1},x_{2},x_{3}]\to\kappa[x_{0},x_{1},x_{2}]\to 0.$
+Then, for any associated prime $P$ of the Jacobian ideal of $X$ at $\kappa[x_{0},x_{1},x_{2}]$, the extension will satisfy
+$\dim R/(R\cdot P)$ $=\dim(\kappa[x_{0},x_{1},x_{2}]/P)+1$
+$\deg R/(R\cdot P)$ $=\deg(\kappa[x_{0},x_{1},x_{2}]/P).$
+since the generators do not include $x_{3}$.
+Moreover, the length of the modules over $P$ will be preserved, as the element $x_{3}\not\in P$ will be invertible.
+Together with the associativity formula, we obtain:
+$\tau(X)$ $=\sum_{P\in\operatorname{Ann}(\mathcal{O}_{\operatorname{Sing}X}(d)),\dim(\kappa[x_{0},x_{1},x_{2}]/P)=1}\operatorname{{length}}(\mathcal{O}_{\operatorname{Sing}X}(d))_{P}\cdot\deg(\kappa[x_{0},x_{1},x_{2}]/P)$
+$=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(\kappa[x_{0},x_{1},x_{2},x_{3}]/P)=2}\operatorname{{length}}(\mathcal{Q}_{\sigma})_{P}\cdot\deg(R/P)=m(\sigma).$
+The following example of the nature above is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ but $m(\sigma)\neq 0$, showing that a converse for 8, $(a)$ is not possible without assuming the Buchsbaum–Rim resolution.
+Example 11.
+Let $f=x_{3}$ and $g=x_{0}x_{1}x_{2}^{2}+x_{0}^{4}+x_{1}^{4}$.
+The quartic plane curve $X=V(g)\subset\mathbb{P}^{3}_{[x_{0}:x_{1}:x_{2}]}$ is nodal and the minimum degree for a syzygy of $\nabla g$ is $e=d$.
+Hence, $m(\sigma)=\tau(X)=1\neq 0$, accounting for the nodal singularity, but $e=d=3$.
+2.3 Free resolutions and the Bourbaki scheme
+In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [undefr, Theorem 2.1, (c)].
+The next result follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$, as it is the case for $\mathbb{P}^{2}$.
 Lemma 12.
@@
 $\nabla\sigma=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\end{pmatrix}$.
-The 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the $1$-st Fitting ideal has dimension zero.
-The minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-so that $e=1$, $m(\sigma)=4$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=3$.
+The scheme $(\Xi_{\sigma})_{1}$ has three primary components in dimension one, given by the prime ideals $\mathfrak{p}_{1}=(x_{0},x_{3})$, $\mathfrak{p}_{2}=(x_{0},x_{2})$ and $\mathfrak{p}_{3}=(x_{0},x_{1})$.
+We also note that there is an easy syzygy of degree one, namely $\nu=(-x_{3},0,0,x_{0})^{T}$, hence $e=1$.
+To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated to the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
+Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{P}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}_{P}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$.
+From this, we may rewrite the matrix as
+$(\nabla\sigma)_{\mathfrak{p}_{1}}=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 1&ux_{0}x_{2}&ux_{0}x_{1}&3ux_{3}^{2}\\ \end{pmatrix}\sim\begin{pmatrix}0&-2x_{0}^{2}x_{2}u&-2ux_{0}^{2}x_{1}&2x_{3}-6ux_{0}x_{3}^{2}\\ 1&x_{0}x_{2}u&x_{0}x_{1}u&3x_{3}^{2}u\\ \end{pmatrix}$,
+first by multiplying the second row by $u$, and then adding the second row scaled by $2x_{0}$ to the first row.
+From the final form of the matrix above, it is easy to see that it sends the fourth basis vector to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and thus we may compute the cokernel as the remaining entries of the first row:
+$\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(-2x_{0}^{2}x_{2}u,-2ux_{0}^{2}x_{1},2x_{3}-6ux_{0}x_{3}^{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})},$
+a $R_{\mathfrak{p}_{1}}$-module of length two.
+Over the prime $\mathfrak{p}_{2}=(x_{0},x_{2})$, the element $x_{3}\in R_{\mathfrak{p}_{2}}$ is an unit.
+We denote by $u=x_{3}^{-1}\in R_{\mathfrak{p}_{2}}$ its inverse, and we rewrite the matrix using elementary operations
+$(\nabla\sigma)_{\mathfrak{p}_{2}}=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}-3x_{0}x_{3}&x_{0}x_{2}&x_{0}x_{1}&0\\ \end{pmatrix}\sim\begin{pmatrix}ux_{0}&0&0&1\\ 3x_{0}^{2}+x_{1}x_{2}-3x_{0}x_{3}&x_{0}x_{2}&x_{0}x_{1}&0\\ \end{pmatrix}$,
+first by adding the first row scaled by $(-\frac{3}{2}x_{3})$ to the second row, and then multiplying the first row by $u/2$.
+Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{P}^{2}$, and thus
+$\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0}x_{1},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
+$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
+$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{0}x_{2},x_{1}x_{2})}$
+$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{2})}\simeq\kappa$
+since $x_{1}\in R_{\mathfrak{p}_{2}}^{\times}$.
+Thus, it follows $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+For the ideal $\mathfrak{p}_{3}$, we note that the approach is analogous as the previous one, using the same elementary operations, and obtaining the same isomorphism at the end $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{3}}\simeq\kappa$, so that from the associativity formula:
+$m(\sigma)=\sum_{i=1}^{3}\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4.$
+Using the Bourbaki formula with $m_{0}=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}=7$, $d=3$ and the obtained discrete invariants $m(\sigma)=4$ and $e=1$, we conclude that $\operatorname{Bour}(\sigma)=1$.
+From the minimal free resolution of the ideal of a line in $\mathbb{P}^{3}$, using Lemma 12, the minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
+From the resolution, one obtains $c_{3}({\mathcal{T}}_{\sigma})=3$.
 Example 15 ($3$-syzygy pencil of cubics which is not nearly-free).
@@
 where $f$ is a hyperplane arrangement and $g$ is the union of a plane and a smooth quadric, with the Jacobian matrix:
-$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{0}x_{3}-x_{1}x_{3}&x_{0}x_{2}-x_{1}x_{2}\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
-In this case, the 0-th Fitting ideal coincides with the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$, and the saturation of the first Fitting ideal is zero.
-The primary decomposition of $\Xi_{\sigma}$ is described in the following table:
-dimension degree radical ideal
-1 1 $(x_{2},x_{3})$
-1 1 $(x_{3},x_{0}-x_{1})$
-1 1 $(x_{2},x_{0}-x_{1})$
-1 2 $(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
-0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
-0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
-Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
+$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
+The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals of $R$:
+$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
+$\mathfrak{p}_{2}$ $=(x_{3},x_{0}-x_{1})$
+$\mathfrak{p}_{3}$ $=(x_{2},x_{0}-x_{1})$
+$\mathfrak{p}_{4}$ $=(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
+We claim that $m(\sigma)=5$, a quantity which will be obtained using the associativity formula. For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we start by noting that the quantities:
+$h$ $=3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$
+$v$ $=x_{0}x_{1}$
+$t$ $=(x_{0}-x_{1})$
+are all invertible inside the ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$.
+Then, we do two elementary operations:
+$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}ux_{2}x_{3}+\frac{1}{2}hx_{2}x_{3}&0&utx_{3}+x_{0}x_{2}^{2}x_{3}&utx_{2}+x_{0}x_{2}x_{3}^{2}\\ h&2v&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$
+$\sim\begin{pmatrix}ux_{2}x_{3}+\frac{1}{2}hx_{2}x_{3}&0&utx_{3}+x_{0}x_{2}^{2}x_{3}&utx_{2}+x_{0}x_{2}x_{3}^{2}\\ \frac{1}{2}uh&1&ux_{0}x_{2}&ux_{0}x_{3}\end{pmatrix}$.
+The first is to add $\frac{1}{2}x_{2}x_{3}$ times the second row to the first row, scaled by $u=v^{-1}$.
+Then, we scale the second row by $\frac{1}{2}u$.
+Since the matrix sends the second basis vector into $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, it suffices to look at the first row of the remaining matrix, so that
+$\operatorname{coker}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{((u+\frac{1}{2}h)x_{2}x_{3},utx_{3}+x_{0}x_{2}(x_{2}x_{3}),utx_{2}+x_{0}x_{3}(x_{2}x_{3}))}$
+$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3}+x_{0}x_{2}(x_{2}x_{3}),utx_{2}+x_{0}x_{3}(x_{2}x_{3}))}$
+$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3},utx_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2},x_{3})}\simeq\kappa,$
+and hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
+For $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, note that $v=x_{0}x_{2}\in R_{\mathfrak{p}_{2}}$ is a unit and denote $u=v^{-1}$. We consider some elementary operations:
+$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})h&-x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{1}&0&x_{2}(x_{0}-x_{1})-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{3}\\ h&2x_{0}x_{1}&2v&2x_{0}x_{3}\end{pmatrix}$
+$\sim\begin{pmatrix}x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})h&-x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{1}&0&x_{2}(x_{0}-x_{1})-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{3}\\ \frac{1}{2}uh&ux_{0}x_{1}&1&ux_{0}x_{3}\end{pmatrix}$.
+The first transformation is to add the scaled second row (by $\frac{1}{2}x_{3}(x_{0}-x_{1})u$) to the first one, and the second transformation is rescaling the second row by $\frac{1}{2}u$.
+Then, we see that the matrix sends the third basis vector to $(0,1)\in R_{\mathfrak{p}_{2}}^{2}$, so we may compute the cokernel as the first row of the remaining matrix:
+$\operatorname{coker}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3}(x_{2}-\frac{1}{2}u(x_{0}-x_{1})h),-x_{3}(x_{2}+\frac{1}{2}u(x_{0}-x_{1})x_{0}x_{1}),(x_{0}-x_{1})(x_{2}-\frac{1}{2}ux_{0}x_{3}^{2})))}$
+$\simeq\frac{R_{\mathfrak{p}_{2}}}{x_{3},-x_{3},(x_{0}-x_{1})}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}\simeq\kappa,$
+since for example the quantity
+$x_{2}-\frac{1}{2}u(x_{0}-x_{1})h$
+is of the form a unit $x_{2}$ plus something inside the maximal ideal $(x_{3},x_{0}-x_{1})_{\mathfrak{p}_{2}}$, it is also a unit.
+Thus, we obtain $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
+An analogous consideration also holds for the case of the prime $\mathfrak{p}_{3}=(x_{2},x_{0}-x_{1})$, with similar elementary operations, now to trivialize the fourth column as $(0,1)$, and then to show that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa$.
+For the prime $\mathfrak{p}_{4}=(x_{0},h\doteq x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$, we note that $v=x_{2}x_{3}$ is invertible in $R_{\mathfrak{p}_{4}}$, denote $u=v^{-1}$ and do the elementary operations
+$(\nabla\sigma)_{\mathfrak{p}_{4}}$ $\sim\begin{pmatrix}v&-v&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ h+2x_{0}x_{1}&0&2x_{0}x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})&2x_{0}x_{2}x_{3}(x_{0}-x_{1})\\ \end{pmatrix}$
+$\sim\begin{pmatrix}-1&1&-ux_{3}(x_{0}-x_{1})&-ux_{2}(x_{0}-x_{1})\\ h+2x_{0}x_{1}&0&2x_{0}x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})&2x_{0}x_{2}x_{3}(x_{0}-x_{1})\\ \end{pmatrix}$,
+first adding the scaled first row $(by2x_{0}x_{1}u)$ to the second row, and then scaling the first row by u-u.
+The matrix sends the second basis vector to $(1,0)\in R_{\mathfrak{p}_{4}}^{2}$, and thus to compute the cokernel we may consider only the second row of the remaining matrix, so that
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}$ $\simeq\frac{R_{\mathfrak{p}_{4}}}{(h+2x_{0}x_{1},x_{0}(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})),2x_{0}x_{2}x_{3}(x_{0}-x_{1}))}$
+$\simeq\frac{R_{\mathfrak{p}_{4}}}{(h+2x_{0}x_{1},x_{0},x_{0})}$
+$\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},h)}\simeq\kappa,$
+noting that the factors $(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1}))$ and $2x_{2}x_{3}(x_{0}-x_{1})$ are units in the local ring.
+Thus, we obtain that the length at each prime is one, but one of them has multiplicity two, therefore $m(\sigma)=5$ from the associativity formula.
+The main result of the next section is:
+$e=1$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 2$
+$e=2$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 5,$
+so for $m_{0}=12$ we conclude that these two correspond to $m(\sigma)\geq 7$ and $m(\sigma)\geq 6$, respectively. Since $m(\sigma)=5$, we conclude automatically that $e\geq 3$.
+Using Macaulay2, we obtain a minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the 33-syzygy kind, given by:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for $\mathcal{I}_{B}$ will be of the form:
+so $e=3$ and, with $m(\sigma)=5$, we conclude $\operatorname{Bour}(\sigma)=4$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for the Bourbaki ideal $\mathcal{I}_{B}$ will be of the form:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}\rightarrow\mathcal{I}_{B}\rightarrow 0,$
@@
   • (b)
-    if $\sigma$ is nearly-free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
+    if $\sigma$ is nearly free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
 Proof.
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\rightarrow\mathcal{I}_{l}\rightarrow 0$
-since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for the syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. ∎
+since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for a choice of syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$.
+The isolated zeros of the minors will be in the support of the sheaf $\operatorname{supp}({\mathcal{T}}_{\sigma})$, and these will be inside the line $l$ from the resolution obtained.
+∎
 Remark 17.
-The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [5], related to rational cuspidal curves.
-In [17, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [5] (see [17, Definition 2.17]).
+The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [undefd], related to rational cuspidal curves.
+In [undefr, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [undefd] (see [undefr, Definition 2.17]).
 Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
-The notion of 33-syzygy divisors is also present in a number of works in the field, for example [1], [6] and [8].
+The notion of 33-syzygy divisors is also present in a number of previous works, for example [undef], [undefe] and [undefg].
 Example 18.
@@
   • For $k=2$, $(f,g)$ is a free pencil of cubics with $e=1$, $m(\sigma)=9$;
-  • For $k>2$, $(f,g)$ is nearly-free with $e=1$.
+  • For $k>2$, $(f,g)$ is nearly free with $e=1$.
+The Jacobian matrix for $\sigma=(f,g)$ is
+$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\\ (k+1)x_{0}^{k}&0&0&(k+1)x_{3}^{k}\end{pmatrix}$.
 For $k=2$, we note that the matrix below:
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-1)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)^{\oplus 2}\xrightarrow{M}\mathcal{T}_{\sigma}\rightarrow 0,$
-and in particular $\sigma$ is a nearly-free sequence.
+and in particular $\sigma$ is a nearly free sequence.
+Proposition 19.
+If $\sigma=(f,g)$ is a nearly free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
+Proof.
+Since
+$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
+assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
+$e=\frac{d+2}{2},$
+which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
+If we denote by $E\doteq{\mathcal{T}}_{\sigma}(d/2)$ the normalized vector bundle, we must have $c_{1}(E)=0$ and
+$c_{2}(E)$ $=c_{2}({\mathcal{T}}_{\sigma})-\frac{d^{2}}{2}+\frac{d^{2}}{4}$
+$=m_{0}-m(\sigma)-\frac{d^{2}}{4}.$
+On the other hand, using the Bourbaki degree formula with $\operatorname{Bour}(\sigma)=1$ and $e=\frac{d+2}{2}$, we obtain
+$1=\operatorname{Bour}(\sigma)$ $=\frac{(d+2)^{2}}{4}-\frac{d(d+2)}{2}+m_{0}-m(\sigma)$
+$=\frac{d^{2}}{4}-\frac{d^{2}}{2}-d+\frac{4d}{4}+1+m_{0}-m(\sigma),$
+which yields
+$m_{0}-m(\sigma)-\frac{d^{2}}{4}=0,$
+so that $c_{2}(E)=0$.
+But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ ([undefn, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis.
+∎
 Proposition 20.
@@
 and focus on the second one.
-The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [14, Proposition 1.1].
+The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [undefn, Proposition 1.1].
 Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
@@
 so that $\operatorname{{rk}}(F_{0}^{\prime})=3$, thus $\operatorname{{rk}}(F_{0})=2$ and $\sigma$ is $3$-syzygy. ∎
-The following proposition follows from the formula in Remark 7 for $c_{3}({\mathcal{T}}_{\sigma})$ in terms of $e,\deg(B)$ and $p_{a}(B)$ for a Bourbaki scheme $B=B_{\nu}$.
-Proposition 21.
-If $\sigma=(f,g)$ is a nearly-free sequence such that ${\mathcal{T}}_{\sigma}$ is locally free, then $d=d_{f}+d_{g}$ must be even and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
-Proof.
-Since
-$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
-assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly-free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
-$e=\frac{d+2}{2},$
-which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$. ∎
 Remark 21.
@@
 $\nabla\sigma=\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\\ 0&x_{2}x_{3}&3x_{2}^{2}+x_{1}x_{3}&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$.
-We may check computationally that the saturation of the ideal of $(2\times 2)$-minors coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$ and with the saturation of the 0-th Fitting ideal, therefore over this subset $\mathcal{Q}_{\sigma}$ has rank one.
-The $1$-dimensional part of this scheme has $3$ irreducible components, two lines $L_{1}=V(x_{0},x_{3}),L_{2}=V(x_{2},x_{3})$ and a quadric plane curve $Q=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})$.
-Moreover, $L_{1}$ and $Q$ have multiplicity two structure, adding to $m(\sigma)=5$.
-This can also be obtained from the minimal free resolution of ${\mathcal{T}}_{\sigma}$:
+The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines and a plane quadric curve:
+$L_{1}$ $=V(x_{0},x_{3})\doteq V(\mathfrak{p}_{1})$
+$L_{2}$ $=V(x_{2},x_{3})=V(\mathfrak{p}_{2})$
+$Q$ $=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})=V(\mathfrak{p}_{3}),$
+Now, we proceed to computing the length of the cokernel module $Q$ over each prime ideal $\mathfrak{p}_{i}$, in order to show that $m(\sigma)=5$ using the associativity formula.
+Starting with the prime ideal $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that the element $3x_{2}^{2}\in R_{\mathfrak{p}_{1}}$ is invertible, and hence the sum $v=3x_{2}^{2}+x_{1}x_{3}$ is also invertible, as $x_{1}x_{3}\in\mathfrak{p}_{1}$.
+Let $u=v^{-1}$, so we perform two elementary operations on the matrix $\nabla\sigma$ over $\mathfrak{p}_{1}$:
+$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&x_{2}x_{3}&v&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$
+$\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&ux_{2}x_{3}&1&u(x_{1}x_{2}+3x_{3}^{2})\end{pmatrix}$.
+The first transform is to add the scaled second row (by $ux_{0}^{2}$) to the first, and then we scale the second row by $u$.
+Then, the third basis vector goes via $\nabla\sigma$ to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and therefore we may compute the cokernel from the quotient
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(2x_{0}x_{2}+x_{1}x_{3},x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3},x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2}))}\simeq\frac{R_{\mathfrak{p}_{1}}}{I},$
+where $I=(f_{1},f_{2},f_{3})$ are the generators above.
+We may factor $f_{2}=x_{0}x_{3}(1-ux_{0}x_{2})$, so that since $(1-ux_{0}x_{2})$ is a unit in the local ring, we conclude $x_{0}x_{3}\in I$.
+From this, multiplying $x_{3}f_{1}$ we conclude that $x_{1}x_{3}^{2}\in I$, and since $x_{1}$ is also a unit, we obtain $x_{3}^{2}\in I$, and analogously taking $x_{0}f_{1}$ we obtain $x_{0}^{2}x_{2}\in I$, and since $x_{2}$ is a unit we get $x_{0}^{2}\in I$.
+Thus
+$(x_{0}^{2},x_{0}x_{3},x_{3}^{2})\subset I.$
+Now, from this, looking at $f_{3}$:
+$f_{3}=x_{0}x_{1}+3x_{3}^{2}-x_{0}^{2}u(x_{1}x_{2}+3x_{3}^{2})\in I\Rightarrow x_{0}x_{1}\in I\Rightarrow x_{0}\in I,$
+and looking at $f_{1}$, we obtain
+$f_{1}=2x_{0}x_{2}+x_{1}x_{3}\in I\Rightarrow x_{1}x_{3}\in I\Rightarrow x_{3}\in I,$
+so that $I=(x_{0},x_{3})$ and thus $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\kappa$, with $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
+For the case $\mathfrak{p}_{2}=(x_{2},x_{3})$, let $u=(x_{0}^{2})^{-1}$, $q=(3x_{2}^{2}+x_{1}x_{3})$ and consider the elementary operations given by $R_{2}^{\prime}=R_{2}-uqR_{1}$ and then $R_{1}^{\prime}=uR_{1}$:
+$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&u(x_{0}x_{1}+3x_{3}^{2})\\ -uq(2x_{0}x_{2}+x_{1}x_{3})&x_{2}x_{3}-uqx_{0}x_{3}&0&x_{1}x_{2}+3x_{3}^{2}-uq(3x_{3}^{2}+x_{0}x_{1})\end{pmatrix}$.
+The Jacobian matrix sends the third basis vector to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, and thus we may compute:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I}$
+where
+$f_{1}$ $=(3x_{2}^{2}+x_{1}x_{3})(2x_{0}x_{2}+x_{1}x_{3})$
+$f_{2}$ $=x_{3}(x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3})$
+$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3ux_{3}^{2}(3x_{2}^{2}+x_{1}x_{3})-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})$
+First, we consider
+$f_{2}-\frac{x_{3}}{x_{1}}f_{3}=\frac{3x_{3}^{3}}{x_{1}}(1-u(3x_{2}^{2}+x_{1}x_{3}))\in I,$
+and thus $x_{3}^{3}\in I$. Then
+$x_{3}f_{2}=x_{3}^{2}x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3}^{3}\in I\Rightarrow x_{3}^{2}x_{2}\in I,$
+and moreover
+$x_{3}f_{1}$ $=x_{2}x_{3}(6x_{0}x_{2}^{2}+3x_{1}x_{2}x_{3}+2x_{0}x_{1}x_{3})+x_{1}^{2}x_{3}^{3}\in I$
+$\Rightarrow 6x_{0}x_{2}^{3}x_{3}+x_{3}^{2}x_{2}(3x_{1}x_{2}+2x_{0}x_{1})\in I$
+$\Rightarrow 6x_{0}x_{2}^{3}x_{3}\in I$
+$\Rightarrow x_{2}^{3}x_{3}\in I.$
+On the other hand,
+$x_{2}x_{3}f_{3}$ $=-3ux_{3}^{3}q+3x_{2}x_{3}^{3}+x_{2}x_{3}(x_{1}x_{2}-ux_{0}x_{1}q)\in I$
+$\Rightarrow x_{1}x_{2}^{2}x_{3}-3ux_{0}x_{1}x_{2}^{3}x_{3}-ux_{0}x_{1}^{2}x_{2}x_{3}^{2}\in I$
+$\Rightarrow x_{1}x_{2}^{2}x_{3}\in I\Rightarrow x_{2}^{2}x_{3}\in I.$
+Now, consider
+$x_{3}f_{3}$ $=x_{1}x_{2}x_{3}+3x_{3}^{3}-uqx_{3}^{3}-uqx_{0}x_{1}x_{3}\in I$
+$\Rightarrow x_{3}x_{2}(x_{2}-3ux_{0}x_{2}^{2}-ux_{0}x_{1}x_{3})\in I$
+$\Rightarrow x_{1}x_{3}(x_{2}-ux_{0}x_{1}x_{3})\doteq A\in I,$
+and
+$x_{3}A\in I\Rightarrow x_{1}x_{3}^{2}x_{2}-ux_{0}x_{1}^{2}x_{3}^{2}\in I\Rightarrow x_{3}^{2}\in I.$
+This simplifies
+$f_{2}=x_{3}(x_{2}-uqx_{0})\in I\Rightarrow x_{2}x_{3}(1-3ux_{2})\in I\Rightarrow x_{2}x_{3}\in I.$
+Considering
+$f_{1}$ $=6x_{0}x_{2}^{3}+x_{2}x_{3}(3x_{1}x_{2}+2x_{0}x_{1})+x_{1}^{2}x_{3}^{2}\in I$
+$\Rightarrow x_{2}^{3}\in I,$
+and then
+$x_{2}f_{3}$ $=x_{1}x_{2}^{2}+x_{2}(3x_{3}^{2}-3ux_{3}^{2}q)-uqx_{0}x_{1}x_{2}\in I$
+$\Rightarrow x_{1}x_{2}^{2}-3ux_{0}x_{1}x_{2}^{3}-ux_{0}x_{1}^{2}x_{2}x_{3}\in I$
+$\Rightarrow x_{1}x_{2}^{2}\in I\Rightarrow x_{2}^{2}\in I,$
+simplifying
+$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3uqx_{3}^{2}-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})\in I$
+$\Rightarrow x_{1}(x_{2}-ux_{0}x_{1}^{2}x_{3})\in I\Rightarrow x_{2}-ux_{0}x_{1}x_{3}\in I,$
+so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa\langle 1,x_{2}\rangle$, meaning $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=2$, since the equations $f_{1},f_{2}$ are trivialized after $x_{3}^{2},x_{2}^{2},x_{2}x_{3}\in I$.
+Let us consider the case $\mathfrak{p}_{3}=(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})\doteq(g_{1},g_{2})$.
+One may be tempted to choose $u=(2x_{0}x_{2}+x_{1}x_{3})^{-1}$ and do the elementary operations $R_{1}^{\prime}=uR_{1}$, but
+$2x_{0}x_{2}+x_{1}x_{3}=2x_{0}g_{1}+g_{2}.$
+Take $u\doteq(x_{0}^{2})^{-1}\in R_{\mathfrak{p}_{3}}$. If $x_{0}^{2}\in\mathfrak{p}_{3}$, then:
+$x_{0}\in\mathfrak{p}_{3}\Rightarrow x_{2}\in\mathfrak{p}_{3}\Rightarrow x_{1}x_{3}\in\mathfrak{p}_{3}$
+and $\text{height}(\mathfrak{p}_{3})\geq 3$ would give a contradiction. Thus, we may perform the same elementary operations as we did for $\mathfrak{p}_{2}$, and thus
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{I}$
+with generators $I=(f_{1},f_{2},f_{3})$ given as before. By the same reasoning, $(3x_{2}^{2}+x_{1}x_{3})\not\in\mathfrak{p}_{3}$, and thus
+$f_{1}\in I\Rightarrow 2x_{0}x_{2}+x_{1}x_{3}\in I.$
+From the previous description of the element above, it suffices to show $g_{i}\in I$ for some $i\in\{1,2\}$ to conclude $I=\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$.
+Let $A\doteq(g_{2}+x_{2}^{2})$, so that $u=A^{-1}$.
+We can rewrite $f_{3}$ as:
+$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-u(g_{2}+x_{2}^{2})(3x_{3}^{2}+x_{0}x_{1})$
+$=3x_{3}^{2}(1-u(g_{2}+x_{2}^{2}))+x_{1}(x_{2}-ux_{0}(g_{2}+x_{2}^{2}))$
+$=3x_{3}^{2}(1-1)+x_{1}(x_{2}-ux_{0})$
+$=-x_{1}g_{1}\in I,$
+and since $x_{1}\notin\mathfrak{p}_{4}$, $g_{1}\in I$ and $I=\mathfrak{p}_{4}R_{\mathfrak{p}_{4}}$, concluding $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=1$.
+From the associativity formula, we obtain
+$m(\sigma)=1\cdot 1+2\cdot 1+1\cdot 2=5.$
+Since $m(\sigma)\leq 6$, we conclude using the bounds from the next section that $e\geq 3$ The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{M}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$.
-This implies that $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$.
+so that $e=3$ and $\operatorname{Bour}(\sigma)=4$.
+This means $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$, but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$ or, equivalently, the degree four Bourbaki scheme is not a complete intersection.
 Example 23 (pencil of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$).
@@
 $\sigma=(x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3},x_{2}x_{3}(x_{2}-x_{1})),$
-considered in [11, Theorem 8.1].
+considered in [undefj, Theorem 8.1].
 From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle.
 Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{L}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-So we compute $e=3$ and $m(\sigma)=7$.
-This may also be seen from the primary decomposition of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, which is composed of three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$, where the last line has a multiplicity $5$ structure, and at every line the rank of $\mathcal{Q}_{\sigma}$ is one.
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
+So $e=3$ and $m(\sigma)=7$.
+The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$.
+Localizing the cokernel at the first two lines gives length one, so over last line should have length $5$.
 Example 24 (pencils of cubics with the same discrete invariants and different homological behavior).
@@
 $\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}+x_{3}^{2}&x_{0}^{2}&x_{0}x_{1}+2x_{1}x_{3}+3x_{3}^{2}\end{pmatrix}$.
-Here, although the 0-th Fitting ideal is different from the annihilator ideal of $\mathcal{Q}_{\sigma}$, their one-dimensional component coincide with a multiplicity five structure along the line $V(x_{3},x_{0})$, and thus $m(\sigma)=5$.
+Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$.
 Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 6}\rightarrow\mathcal{I}_{B}\rightarrow 0.$
-This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as 16, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
+This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as Example 15, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
 In Section 4 we characterize all nearly free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly free sequences with $d_{f}=1,d_{g}=2$.
@@
 We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two.
-From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 28.
-We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example [4]).
-As explored in [11, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
+From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 27.
+We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example, [undefc]).
+As explored in [undefj, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
 $0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
-where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the foliation.
+where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the associated foliation.
 Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
@@
   • (a)
-    ([4, $2.3$, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1.
+    ([undefk, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1.
 In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
@@
     (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$ or is zero-dimensional.
-Theorem B. Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
+For completeness, we include an argument for $(b)$, from V. Cordeiro: using the short exact sequence
+$0\to G^{\vee}\to\Omega^{1}_{\mathbb{P}^{3}}\to\mathcal{I}_{W}(1)\to 0,$
+one obtains $4\leq c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [undefc, 4.1, 4.2]).
+Now, if $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$.
+Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a section to form a sequence
+$0\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\to G^{\vee}\to\mathcal{I}_{Y}(-3)\to 0$
+where $Y$ is the vanishing locus of this section.
+Hence, from additivity one gets $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$.
+From the first inequality, we conclude $c_{2}(G^{\vee})\geq 6$, and thus $\deg(W)\leq 5$.
+Theorem B.
+Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
   • (a)
@@
 Theorem 25.
-[18, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree $2$ and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
+[undeft, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree 22 and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
   • (a)
@@
 The results above can be used to restrict the possible values of numerical invariants.
-For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results above to show that the only possible discrete invariant is $c_{3}({\mathcal{T}}_{\sigma})=8$, although the existence of such an example remains an open question.
-Indeed, from [19, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 7 we obtain $c_{3}({\mathcal{T}}_{\sigma})=8-2a$, where $p_{a}(B)=-1-a$.
+For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results to show the only possible third Chern class is $c_{3}({\mathcal{T}}_{\sigma})=8$.
+Indeed, from [undefu, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 6 we obtain $c_{3}({\mathcal{T}}_{\sigma})=2(p_{a}(B)+5)=8-2a$, where $p_{a}(B)=-1-a$.
 From the short exact sequence
@@
 and additivity of Euler characteristic gives a contradiction whenever $a\neq 0$, so $c_{3}({\mathcal{T}}_{\sigma})=8$ if this occurs.
-Using the classification of $\mu$-semistable reflexive sheaves on $\mathbb{P}^{3}$ with small $c_{2}$, we are able to obtain a full classification of the cases of pencils of cubics with $m(\sigma)=7$ and $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$ (see 42).
+In the next section, we show the above cannot happen either (see 43).
 As another application, we are able to show that sequences $\sigma$ with mixed degrees $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$ cannot be unstable.
 Proposition 27.
-Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$.
-If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, and in particular ${\mathcal{T}}_{\sigma}$ must be $\mu$-stable.
+Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$ and $m(\sigma)=3$.
+If ${\mathcal{T}}_{\sigma}$ is unstable, then the possible discrete invariants are $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=6$.
 Proof.
 Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$.
-From Section 3.1, $(c)$, since $m(\sigma)=3\neq 7$, $\sigma$ is incompressible, and from [16, Theorem 1.1], we obtain $c_{3}\leq 4$ (for a more general formula, see 45).
-On the other hand, using Remark 7 we have $c_{3}({\mathcal{T}}_{\sigma})=6-2a$, so $a\in\{1,2,3\}$.
+From Theorem A Theorem A, $\sigma$ is incompressible.
+Using Remark 6 we have for $\deg(B)=2$, $p_{a}(B)=-1-a$ yields $c_{3}({\mathcal{T}}_{\sigma})=6-2a\geq 0$, so $a\in\{-1,0,1,2,3\}$, but from
 From the short exact sequence
@@
 $\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=8-2a,$
-which cannot coincide for the values $a\in\{1,2,3\}$, hence we obtain a contradiction. ∎
+which can only coincide for $a=0$. ∎
+We note that we have no explicit construction for the unstable case above to occur as an example.
 4 Pencils of cubics and degree 6 curves inside quadric surfaces
 In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces.
-The results are derived from the previous sections, Section 3.1 and Section 4, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [14], [21], [15], [3] and [16].
+The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [undefn], [undefw], [undefo], [undefb] and [undefp].
+Two major results which are used are the following ones, both assuming that $\kappa$ is an algebraically closed field of characteristic zero.
+Theorem 28 ([undefp], Theorem 1.1).
+Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$. Assume $c_{1}\geq-3$ and $h^{0}(\mathcal{E})=0$. Define the integers:
+$A$ $\doteq\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil,\text{resp. ditto +1 if }c_{1}=1,3;$
+$B$ $\doteq\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil,\text{resp. ditto +1 if }c_{1}=2,4;\text{ditto -1 if }c_{1}=-3.$
+Then $c_{2}\geq A$. Furthermore:
+  • (a)
+    If $A\leq c_{2}\leq B$, then
+    $c_{3}\leq(c_{1}+4)c_{2}-2\binom{c_{1}+3}{3}-2.$
+  • (b)
+    If $c_{2}>B$, then
+    $c_{3}\leq c_{2}^{2}-c_{2}(2B-c_{1}-5)+B^{2}-B-2\binom{c_{1}+3}{3}-2.$
+With this theorem, we obtain an upper bound for $m(\sigma)$ lower than $m_{0}$ when $\sigma$ is incompressible, so that $h^{0}({\mathcal{T}}_{\sigma})=0$, in the case $d_{f}=1,d_{g}=2$.
+In the case of pencils of cubics, however, we apply this to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, so we must assume $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ and treat the case $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ separately.
+Another important result is the following, which enables us to obtain a lower bound for the existence of a syzygy in some cases:
+Theorem 29 ([undefo], Theorem 0.1).
+Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$ with $c_{1}=0$ of $c_{1}=-1$ and with $c_{2}\geq 0$. Let $t\in\mathbb{Z}$ such that either
+  • (a)
+    $c_{1}=0$ and $t>\sqrt{3c_{2}+1}-2$, or
+  • (b)
+    $c_{1}=-1$ and $t>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}$.
+Then $H^{0}(\mathcal{E}(t))\neq 0$.
 4.1 Pencils of cubics
-In this section, we focus on the case of pencils of cubics, i.e., $d_{f}=d_{g}=2$.
-The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
+Now, assume $d_{f}=d_{g}=2$. The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
 $\operatorname{Bour}(\sigma)=e(e-4)+12-m(\sigma).$
 When $e=1$, the formula above implies $m(\sigma)\leq 9$, from the inequality $\operatorname{Bour}(\sigma)\geq 0$.
-The assumption $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ enables us to apply the result [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}(1)$ and obtain the following upper bounds for $m(\sigma)$ and the third Chern class $c_{3}({\mathcal{T}}_{\sigma})$:
+Applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get
+$c_{1}(\mathcal{E})=c_{1}({\mathcal{T}}_{\sigma}(1))=-d+2=-2$
+and the quantities of the Theorem become
+$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(4-4+3)\right\rceil=1$
+$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil=\left\lceil\frac{1}{3}(4-12+8)\right\rceil=0.$
+Moreover, another important invariant is:
+$c_{2}(\mathcal{E})$ $=c_{2}({\mathcal{T}}_{\sigma})-d+1$
+$=m_{0}-m(\sigma)-d+1$
+$=9-m(\sigma).$
+Direct application of the result shows the next pair of propositions:
 Proposition 30.
@@
   • (a)
-    if $7\leq m(\sigma)\leq 8$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
+    if $8\leq m(\sigma)\leq 9$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
   • (b)
@@
     $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72.$
-The following result is obtained using [15, Theorem 0.1] and the bound obtained in Section 3.1, $(a)$:
+Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=9-m(\sigma)\geq 0$ from the previous computations, assuming $e>1$.
+Then, the Theorem says that, for
+$t$ $>\sqrt{3c_{2}+1}-2=\sqrt{27-3m(\sigma)+1}-2$
+we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$.
+Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives minimum $t=2,1,1$ satisfying the inequality above, so we obtain
 Proposition 32.
@@
 $m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
-6 $e\leq 3$
 7 $e\leq 3$
@@
 Then $m(\sigma)=0$, the number of singular members is $32$, and all singular members have one singular point, in particular $m(\sigma)=0$.
-Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$.
-(see Example 11).
+Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$ (see Example 10), and the minimal free resolution for the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ is of the form
+$0\to\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0,$
+given by the Buchsbaum–Rim complex.
 Proof.
-This follows from intersection theory for the bundle of principal parts (see [10], Proposition $7.1$ and Proposition $7.4$). ∎
-Rewriting Section 3.1, $(d)$ for pencils of cubics, we obtain:
+This follows from intersection theory for the bundle of principal parts (see [undefi], Proposition $7.1$ and Proposition $7.4$). ∎
+Rewriting Theorem A Theorem A, $(d)$ for pencils of cubics, we obtain:
 Proposition 34.
-Let $\sigma$ be a normal pencil of cubics. Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 34).
+Let $\sigma$ be a normal pencil of cubics.
+Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 35), and in this case ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)$.
 Example 35 (Free, compressible pencil of cubics).
@@
 $\nu=\begin{pmatrix}0&-x_{0}x_{1}^{3}\\ 0&x_{0}^{3}x_{1}\\ 1&0\\ 0&-x_{0}^{3}x_{3}+x_{1}^{3}x_{3}\end{pmatrix}$
-Here, $m(\sigma)=12$, the annihilator ideal and the 0-th Fitting ideal coincide, with the support being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
-Using Section 4, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
+Here, $m(\sigma)=12$, with the irreducible components of $(\Xi_{\sigma})_{1}$ being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
+Using Theorem B Theorem B, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
 Proposition 36.
@@
 Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$.
-Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 33, 36 and 38.
+Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 34, 37 and 39.
 Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
-Since $\sigma$ is not free, we can apply the result Section 4, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$.
+Since $\sigma$ is not free, we can apply the result Theorem B Theorem B, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$.
 On the other hand, we have
@@
 $=8-m(\sigma)>5$
-if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Section 4, $(b)$. ∎
+if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Theorem B Theorem B, $(b)$. ∎
 Proposition 37.
@@
 Proof.
-The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 29 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$.
+The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 30 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$.
 From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
@@
 ∎
-Example 37 (Free, unstable incompressible pencil of cubics ($m(\sigma)=9$)).
+Example 38 (Free, unstable and incompressible pencil of cubics).
 Consider the sequence $\sigma=(x_{1}(x_{2}^{2}-x_{1}^{2}),x_{3}x_{2}(x_{0}-x_{1}))$. Then the matrix $\nabla\sigma$ is given by:
@@
 Proof.
-If $m(\sigma)=8$, by 31, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 2$.
-If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [3, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$.
-If we assume $e=1$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 7 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain $c_{3}({\mathcal{T}}_{\sigma})=4$, on the other hand, 30 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction and this case cannot happen.
-On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$.
+If $m(\sigma)=8$, by 32, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$.
+If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [undefb, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$.
+If we assume $e\in{1,3}$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 6 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain:
+$e=1$ $\Rightarrow c_{3}=4$
+$e=3$ $\Rightarrow c_{3}=0$
+On the other hand, 31 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction with the case $e=1$.
+For the case $e=3$, we would have a nearly-free locally free pencil, which cannot happen by 19.
+Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$.
 ∎
@@
 with corresponding Jacobian matrix given by:
-$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$
-Using Macaulay2 software, we compute $e=2$ and $m(\sigma)=7$, so that $\operatorname{Bour}(\sigma)=1$.
+$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$.
+There are four irreducible components in $(\Xi_{\sigma})_{1}$, all lines, corresponding to the four prime ideals below:
+$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
+$\mathfrak{p}_{2}$ $=(x_{1}-x_{2},x_{3})$
+$\mathfrak{p}_{3}$ $=(x_{1},x_{2})$
+$\mathfrak{p}_{4}$ $=(x_{0},x_{2})$
+For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and the elementary operation $R_{1}^{\prime}=uR_{1}$, to obtain the matrix
+$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}1&u(x_{0}^{2}+x_{2}^{2})&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$,
+which sends the first basis vector to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$ and therefore we may compute the cokernel by considering the second line of the remaining matrix, namely
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3}(2x_{2}-x_{1}),x_{2}(x_{2}-x_{1}))}$
+$\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3},x_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}}\simeq\kappa,$
+so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
+For the prime $\mathfrak{p}_{2}=(x_{1}-x_{2},x_{3})$, the same element $u\in R_{\mathfrak{p}_{2}}$ is invertible, so that we compute the cokernel analogously to obtain:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(-x_{2}x_{3},x_{3}(2x_{2}-x_{1}),x_{2}(x_{2}-x_{1}))}$
+$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{1}-x_{2})}\simeq\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
+so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}+u[x_{2}x_{3}]R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to obtain
+$(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}u(2x_{0}(x_{1}-x_{2})-x_{2}^{2})&1&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&0&x_{3}(2x_{2}-x_{1})-ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&x_{2}(x_{2}-x_{1})-ux_{2}^{3}x_{3}\end{pmatrix}$,
+which sends $e_{2}\in R_{\mathfrak{p}_{3}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{3}}^{2}$. Thus, we may compute the cokernel by considering the remaining columns of the second row:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{(f_{1},f_{2})}$
+where
+$f_{1}$ $=x_{3}(2x_{2}-x_{1})+ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))$
+$f_{2}$ $=x_{2}(x_{2}-x_{1})+ux_{2}^{3}x_{3}.$
+Considering $f_{1}\in I$, we obtain
+$2x_{2}(1+ux_{2}(x_{1}-x_{0}+x_{3})-ux_{0}^{2})\equiv x_{1}\mod I.$
+Substituting this into $f_{2}\in I$, we get:
+$0\equiv x_{2}^{2}\left(1-(2(1+ux_{2}(x_{1}-x_{0}+x_{3}))-ux_{0}^{2})+ux_{2}x_{3}\right)\mod I,$
+and the element multiplying $x_{2}^{2}$ is an unit in $R_{\mathfrak{p}_{3}}$, we conclude that $x_{2}^{2}\in I$.
+Hence $x_{1}^{2}\in I$ by $f_{1}\in I$, and we may rewrite $f_{1}$ and $f_{2}$ mod $I$ as:
+$f_{1}$ $=x_{2}(2-ux_{0}^{2})-x_{1}\mod I$
+$f_{2}$ $=-x_{1}x_{2}\mod I,$
+and thus the ideal $I=(x_{1}^{2},x_{2}^{2},x_{1}x_{2},x_{2}v-x_{1})$ where $v\in R_{\mathfrak{p}_{3}}$ is a unit. Hence, one may write
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa\langle 1,x_{1},x_{2}\rangle,$
+so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
+For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and the elementary operations $R_{1}^{\prime}=R_{1}-u[-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})]R_{2}\doteq R_{1}-uQR_{2}$ and $R_{2}^{\prime}=uR_{2}$, to obtain:
+$(\nabla\sigma)_{\mathfrak{p}_{4}}\sim\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}&0&x_{2}^{2}-uQx_{2}(x_{2}-x_{1})\\ 0&-ux_{2}x_{3}&1&ux_{2}(x_{2}-x_{1})\end{pmatrix}$,
+sending $e_{3}\in R_{\mathfrak{p}_{4}}^{4}$ to $(0,1)\in R_{\mathfrak{p}_{4}}^{2}$. Hence, we compute the cokernel as:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{4}}}{I}$
+with
+$f_{1}$ $=2x_{0}x_{1}-x_{2}(2x_{0}+x_{2})=2x_{0}(x_{1}-x_{2})-x_{2}^{2}$
+$f_{2}$ $=(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}$
+$f_{3}$ $=x_{2}^{2}-uQx_{2}(x_{2}-x_{1})$
+$u$ $=(x_{3}(2x_{2}-x_{1}))^{-1}$
+$Q$ $=-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}).$
+Thus,
+$f_{2}$ $=x_{0}^{2}+x_{2}^{2}-ux_{3}x_{0}^{2}x_{2}+2u(x_{1}-x_{0}+x_{3})x_{3}x_{2}^{2}$
+$=x_{0}^{2}(1-ux_{3}x_{2})+x_{2}^{2}(1+2u(x_{1}-x_{0}+x_{3})x_{3})$
+$f_{3}$ $=x_{2}^{2}+u(x_{2}-x_{1})x_{0}^{2}x_{2}-2u(x_{2}-x_{1})(x_{1}-x_{0}+x_{3})x_{2}^{2},$
+where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have
+$x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I,$
+and thus
+$f_{2}\equiv x_{0}(x_{0}(1-ux_{3}x_{2})+2(x_{1}-x_{2})+2u(x_{1}-x_{2})(x_{1}-x_{0}+x_{3})x_{3})\mod I,$
+but since the element inside the parenthesis is invertible in $R_{\mathfrak{p}_{4}}$, we conclude that $x_{0}\in I$.
+From $f_{1}\in I$, we obtain that $x_{2}^{2}\in I$.
+Taking this into account, we conclude that
+$f_{1}$ $\equiv-x_{2}^{2}\mod I$
+$f_{2}$ $\equiv x_{2}^{2}\mod I$
+$f_{3}$ $\equiv x_{2}^{2}(1-2ux_{2}(x_{1}+x_{3})+2ux_{1}(x_{1}+x_{3}))\equiv x_{2}^{2}\mod I,$
+and therefore
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},x_{2}^{2})}\simeq\kappa\langle 1,x_{2}\rangle,$
+with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
+$m(\sigma)=1\cdot 1+1\cdot 1+3\cdot 1+2\cdot 1=7.$
+On the other hand, $e(\sigma)=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
+$0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0.$
 Proposition 43.
-Let $\sigma=(f,g)$ be a normal pencil of cubics such that $m(\sigma)=7$. Then, the only possible $\mu$-semistable cases are:
-  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 40);
+Let $\sigma=(f,g)$ be a normal, $\mu$-semistable pencil of cubics such that $m(\sigma)=7$. Then, the only possible cases are:
+  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 41);
   • $e=3$ and $\sigma$ is locally free, $\operatorname{Bour}(\sigma)=2$ and $B$ is a pair of skew lines (or their degeneration) (see Example 23).
 Proof.
-For $e\geq 2$, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [3, Lemma 2.1], we conclude that the only possible cases are $c_{3}=0$ (stable case) or $c_{3}=2$ (strictly semistable case).
-Thus, for $e=2$, we must have $c_{3}({\mathcal{T}}_{\sigma})=2$, and thus we obtain the second case.
-For the third case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [22] fitting in a sequence of the form
+From 32, we know that $e\leq 3$.
+For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [undefb, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case).
+For the strictly semistable case $e=2$ we obtain $\operatorname{Bour}(\sigma)=1$, so we get the first case.
+For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [undefx] fitting in a sequence of the form
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}(2)\rightarrow\mathcal{I}_{B}(1)\rightarrow 0$
 where $B$ is a pair of skew lines or their degeneration, thus $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=3$. ∎
-As we mentioned in Remark 27, the only possible case with $m(\sigma)=7$ for which we do not have examples is when $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=8$.
-We summarize the results of this subsection in the following theorem:
-Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
+Theorem C.
+Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
   • (a)
@@
 Proof.
-Item $(a)$ is 33 and item $(b)$ with 36 and Example 39. Item $(c)$ follows from 42 and item $(d)$ is 35. ∎
+Item $(a)$ is 34 and item $(b)$ with 37 and Example 40. Item $(c)$ follows from 43 and item $(d)$ is 36. ∎
 To finish this study, we consider a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
@@
 $\nabla\sigma=\begin{pmatrix}3x_{0}^{2}&3x_{1}^{2}&3x_{2}^{2}&3x_{3}^{2}\\ 3x_{0}^{2}&3x_{1}^{2}&x_{3}^{2}&2x_{2}x_{3}\end{pmatrix}$.
-The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$ with multiplicity $4$.
-The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-where the first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2,\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=4$.
-This is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
+The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$.
+Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$, first $R_{1}^{\prime}=R_{1}-R_{2}$, and then $R_{2}^{\prime}=u/3\cdot R_{2}$, where $u=(x_{1}^{2})^{-1}\in R_{\mathfrak{p}}$, to obtain the form:
+$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}0&0&3x_{2}^{2}-x_{3}^{2}&3x_{3}^{2}-2x_{2}x_{3}\\ ux_{0}^{2}&1&(u/3)x_{3}^{2}&(2/3)ux_{2}x_{3}\end{pmatrix}$,
+so it sends the second basis vector to $(0,1)\in R_{\mathfrak{p}}^{2}$, and then we may compute the cokernel using the first line and remaining columns:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{P}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
+$\simeq\kappa\langle 1,x_{2},x_{3},x_{2}^{2}\rangle,$
+and thus we conclude $\operatorname{{length}}(\mathcal{Q})_{\mathfrak{p}}=4$.
+Since $\deg(R/\mathfrak{p})=1$, using the associativity formula we obtain $m(\sigma)=4$.
+A first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2$, and thus $\operatorname{Bour}(\sigma)=4$.
+The sheaf ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
+obtained using Macaulay2, so this is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
+To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not only depend on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
+Example 45.
+Consider the cubic polynomials:
+$f$ $=x_{3}(x_{1}^{2}-x_{0}x_{2})+x_{1}^{2}(x_{0}-x_{1})$
+$g$ $=x_{3}(x_{1}^{2}-x_{0}x_{2})+x_{1}^{3}$
+$g^{\prime}$ $=x_{0}(x_{3}^{2}-x_{1}x_{2})+x_{3}^{3}.$
+Here, $V(f)$ is a cubic surface with singularity type $4A_{1}$, with primary ideals $(x_{0},x_{1},x_{2})$, $(x_{0},x_{1}^{2},x_{3})$ and $(x_{1},x_{2},x_{3})$.
+On the other hand, the surfaces $V(g)$ and $V(g^{\prime})$ are both of type $A_{1}2A_{2}$ and are isomorphic, via the change of coordinates
+$x_{0}$ $\mapsto x_{1}$
+$x_{1}$ $\mapsto x_{3}$
+$x_{3}$ $\mapsto x_{0}$
+and fixing $x_{3}$.
+The sequence $\sigma=(f,g)$ is non-normal with $V(x_{1})$ in the divisorial component of the Jacobian scheme.
+On the other hand, $\sigma^{\prime}=(f,g^{\prime})$ is a normal, generic pencil with $m(\sigma^{\prime})=0$.
 4.2 Degree 6 curves inside quadric surfaces
@@
 $\operatorname{Bour}(\sigma)=e(e-3)+7-m(\sigma).$
-Assuming $\sigma$ is incompressible, as $c_{1}({\mathcal{T}}_{\sigma})=-3$, we can apply [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}$ without further restrictions, and obtain the following upper bounds for $m(\sigma)$ and $c_{3}({\mathcal{T}}_{\sigma})$:
+We start by applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible. The quantities of the theorem become
+$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(9-6+3)\right\rceil=0$
+$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil-1=\left\lceil\frac{1}{3}(9-9+8)\right\rceil-1=1.$
+Moreover, another important invariant is: $c_{2}(\mathcal{E})=7-m(\sigma)$. Direct application of the result shows the next pair of propositions:
 Proposition 46.
 If $\sigma=(f,g)$ is an incompressible normal sequence with $d_{f}=1,d_{g}=2$, then $m(\sigma)\leq 5$.
+Proof.
+Let us assume that $e\in[1,d-2]$ and $d\geq 3$. Then, from the Bourbaki formula and the fact that $\operatorname{Bour}(\sigma)\geq 0$, we obtain
+$m(\sigma)\leq m_{0}+e(e-d)\leq m_{0}+e-d\leq m_{0}-2,$
+which gives the claim. This covers the case $e=1$. If we now assume $e>1$ and apply Theorem 28 to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a reflexive sheaf with $c_{1}=-1$, we obtain the constant is
+$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(1-2+3)\right\rceil=0.$
+Hence, $5-m(\sigma)=c_{2}({\mathcal{T}}_{\sigma}(1))=c_{2}(\mathcal{E})\geq A=0$ and thus $m(\sigma)\leq 5$. ∎
 Proposition 47.
@@
   • (a)
-    If $4\leq m(\sigma)\leq 5$, then $c_{3}\leq 5-m(\sigma))$. In particular, when $m(\sigma)=5$, $\sigma$ is locally free;
+    If $6\leq m(\sigma)\leq 7$, then $c_{3}\leq 5-m(\sigma)$.
   • (b)
-    If $0\leq m(\sigma)<4$, then $c_{3}\leq m(\sigma)^{2}-10m(\sigma)+25.$
-From Section 3.1, $(d)$, we obtain:
+    If $0\leq m(\sigma)<6$, then $c_{3}\leq m(\sigma)^{2}-14m(\sigma)+47$.
+From Theorem A Theorem A, $(d)$, we obtain:
 Proposition 48.
@@
 gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
-The following result is obtained using ([15], $0.1$):
+Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, the theorem says that, for
+$t$ $>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}=\sqrt{15-3m(\sigma)+\frac{1}{4}}-\frac{3}{2}$
+we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$.
+Direct application of this bound for high values of $m(\sigma)=4,5$ we get $t=0,t=1$, respectively. This shows the following proposition.
 Proposition 50.
@@
 Proof.
-To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 49.
-From 4, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
+To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 51.
+From 3, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
 Assuming $e=2k$ is even, we obtain
@@
 Proof.
-Assuming $\sigma$ is non-compressible, by 48, we obtain that $e=1$.
+Assuming $\sigma$ is non-compressible, by 50, we obtain that $e=1$.
 But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free.
@@
 $\begin{pmatrix}0&x_{0}(x_{0}-x_{1})\\ 0&-x_{1}(x_{0}-x_{1})\\ x_{2}&0\\ -x_{3}&-x_{3}(x_{0}+x_{1})\end{pmatrix}$
-gives linearly-independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
+gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
 Proposition 55.
@@
 Proof.
-Using 48, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$.
+Using 50, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$.
 Using Remark 6 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples.
@@
 $\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&-x_{3}&-x_{2}\\ x_{1}x_{3}&x_{3}(x_{0}-x_{2})&-x_{1}x_{3}&x_{1}(x_{0}-x_{2})\end{pmatrix}$.
-In this case, we have equality of schemes $\Xi_{\sigma}=\operatorname{supp}(\mathcal{Q}_{\sigma})$.
-The Jacobian scheme has a structure of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$, which is outside the three lines, and therefore $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$.
-The free resolution of ${\mathcal{T}}_{\sigma}$ is (obtained computationally):
+The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines, so $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$.
+We may compute $m(\sigma)=4$, where the first line above occurs with length two and the other two occur with length one.
+For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform two elementary operations, namely $R_{2}^{\prime}=R_{2}-ux_{3}(x_{0}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite the matrix:
+$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{1}&1&-ux_{3}&-ux_{2}\\ x_{1}x_{3}(1-u(x_{0}-x_{2}))&0&-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u)&(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u)\end{pmatrix}$.
+Thus, it sends the second basis vector $e_{2}\in R_{\mathfrak{p}_{1}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$, and we may compute the cokernel from the second row of the remaining matrix:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{1}x_{3}(1-u(x_{0}-x_{2})),-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u),(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u))}$
+$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{1}x_{3},-x_{3}^{2},x_{1}-x_{2}x_{3}u)}$
+$\simeq\kappa\langle 1,x_{3}\rangle,$
+simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
+For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform two elementary operations $R_{2}^{\prime}=R_{2}+ux_{1}x_{3}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite $\nabla\sigma$ as
+$(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}ux_{1}&ux_{0}&1&-ux_{2}\\ x_{1}x_{3}(1+ux_{1})&x_{3}((x_{0}-x_{2})+ux_{0}x_{1})&0&x_{1}(x_{0}-x_{2})-ux_{1}x_{2}x_{3}\end{pmatrix}$.
+Since it sends $e_{3}$ to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, we may compute the cokernel using the second row and remaining columns:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I},$
+where
+$f_{1}$ $=x_{1}x_{3}(1+ux_{1})$
+$f_{2}$ $=x_{3}((x_{0}-x_{2})+ux_{0}x_{1})$
+$f_{3}$ $=x_{1}((x_{0}-x_{2})-ux_{2}x_{3}).$
+Since $f_{1}\in I$ and $x_{3}(1+ux_{1})\in R_{\mathfrak{p}_{2}}^{\times}$, we obtain $x_{1}\in I$. From this and $f_{2}\in I$, we obtain:
+$x_{3}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
+and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
+For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and the elementary operations $R_{2}^{\prime}=R_{2}-ux_{1}x_{3}R_{1}$, $R_{1}^{\prime}=uR_{1}$ to obtain
+$(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}1&ux_{0}&-ux_{3}&-ux_{2}\\ 0&x_{3}((x_{0}-x_{2})-ux_{0}x_{1})&x_{1}x_{3}(-1+ux_{3})&x_{1}(x_{0}-x_{2})+ux_{1}x_{2}x_{3}\end{pmatrix}$.
+Since it sends $e_{1}$ to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, we may compute the cokernel using the second row and remaining columns:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I},$
+where
+$f_{1}$ $=x_{3}((x_{0}-x_{2})-ux_{1}x_{0})$
+$f_{2}$ $=x_{1}x_{3}(ux_{3}-1)$
+$f_{3}$ $=x_{1}((x_{0}-x_{2})+ux_{2}x_{3}).$
+Since $f_{1}\in I$ and $(x_{0}-x_{2})-ux_{1}x_{0}\in R_{\mathfrak{p}_{2}}^{\times}$, we obtain $x_{3}\in I$. From this and $f_{3}\in I$, we obtain:
+$x_{1}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
+and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
+From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain
+$m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4.$
+From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly free sequence.
+The complete minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\xrightarrow{M}{\mathcal{T}}_{\sigma}\rightarrow 0$
@@
 $M=\begin{pmatrix}x_{0}x_{1}+x_{3}(x_{2}-x_{0})&x_{0}^{2}&x_{0}x_{3}\\ -x_{1}^{2}+x_{1}x_{3}&-x_{0}x_{1}&0\\ x_{1}x_{2}&x_{2}^{2}&x_{0}x_{1}+x_{3}(x_{2}-x_{1})\\ 0&-x_{2}x_{3}&x_{1}x_{3}-x_{3}^{2}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}\\ x_{3}-x_{1}\\ -x_{2}\end{pmatrix}$.
-So we obtain $m(\sigma)=4,e=2$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=1$ corresponding to the point $p$, which is an irreducible component of $\Xi_{\sigma}$ of codimension three.
+which corresponds to the nearly free resolution given in 16.
+We have $c_{3}({\mathcal{T}}_{\sigma})=1$, corresponding to the point $p$, the unique irreducible component of $\Xi_{\sigma}$ of codimension three.
 Proposition 57.
-Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$.
-If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$.
+Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$, such that $m(\sigma)=3$ and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
+Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$.
 We have examples for the cases $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
 Proof.
-First, the bound $c_{3}\leq 4$ is obtained from 45 for $m=m(\sigma)=3$. As we shown in 28, $e\geq 2$. We claim that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ in this case.
 We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$.
 Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$.
-This follows from [21, Proposition 1.1] for $c_{3}=0$, [3, Lemma 2.4] for $c_{3}=2$ and [14, Lemma 9.6] for $c_{3}=4$.
+This follows from [undefw, Proposition 1.1] for $c_{3}=0$, [undefb, Lemma 2.4] for $c_{3}=2$ and [undefn, Lemma 9.6] for $c_{3}=4$.
 ∎
+The unstable case of the above Proposition is numerically described in 27.
 Example 58 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=2$).
@@
 $\nabla\sigma=\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$.
-Here, the saturation of the annihilator ideal coincides with the saturation of the 0-th Fitting ideal, and the first Fitting ideal has codimension three.
-The support of $\mathcal{Q}_{\sigma}$ in codimension two consists of two lines $V(x_{3},x_{0})\cup V(x_{3},x_{0}-x_{1})$, and the first one has multiplicity two.
-Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
+Here, $(\Xi_{\sigma})_{1}$ has two irreducible components, the lines $V(x_{3},x_{0})$ and $V(x_{3},x_{0}-x_{1})$. Let the corresponding primes be denoted by $\mathfrak{p}_{1},\mathfrak{p}_{2}$.
+Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$.
+Consider the elementary operations $R_{2}=R_{2}-u(x_{0}x_{1}+3x_{3}^{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite
+$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{3}&-ux_{3}&0&1\\ 2x_{0}x_{2}+x_{1}x_{3}-ux_{3}&x_{0}x_{3}+uQx_{3}&x_{0}^{2}&0\end{pmatrix}$,
+so the matrix sends $e_{4}\in R_{\mathfrak{p}_{1}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$, so we may compute the cokernel from the second row and remaining columns:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{f_{1},f_{2},f_{3}}\doteq\frac{R_{\mathfrak{p}_{1}}}{I},$
+where
+$f_{1}$ $=2x_{0}x_{2}+x_{1}x_{3}-u(x_{0}x_{1}+3x_{3}^{2})x_{3}$
+$f_{2}$ $=x_{3}(1+u(x_{0}x_{1}+3x_{3}^{2}))$
+$f_{3}$ $=x_{0}^{2}.$
+Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$. Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain
+$uf_{1}=2x_{0}^{3}x_{1}x_{2},$
+and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})}\simeq\kappa\langle 1,x_{0}\rangle$
+so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
+Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider the elementary operation and $R_{2}^{\prime}=uR_{2}$, to obtain the matrix
+$(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$,
+which sends $e_{3}\in R_{\mathfrak{p}_{2}}^{4}$ to $(0,1)\in R_{\mathfrak{p}_{2}}^{2}$, so we may compute the cokernel using the remaining columns of the first row, namely:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}=\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
+and therefore $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+Thus, from the additivity formula, we obtain $m(\sigma)=3$. Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$.
+The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-so we get $\operatorname{Bour}(\sigma)=2,m(\sigma)=3$ and $c_{3}({\mathcal{T}}_{\sigma})=2$.
+so we get $c_{3}({\mathcal{T}}_{\sigma})=2$ and the Bourbaki scheme is a union of two skew lines.
 Example 59 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=4$).
@@
 with corresponding Jacobian matrix given by
-$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{2}x_{3}-x_{3}^{2}&x_{3}^{2}-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})+2x_{3}(x_{1}-x_{0})\end{pmatrix}$.
-Here, the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $R$.
-The codimension two locus of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of three simple lines.
-Thus, $m(\sigma)=3$.
-Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
+$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{3}(x_{2}-x_{3})&x_{3}(x_{3}-x_{2})&x_{3}(x_{0}-x_{1})&(x_{0}-x_{1})(x_{2}-2x_{3})\end{pmatrix}$.
+Here, the irreducible components of $(\Xi_{\sigma})_{1}$ are three lines, given by the associated primes:
+$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
+$\mathfrak{p}_{2}$ $=(x_{0}-x_{1},x_{3})$
+$\mathfrak{p}_{3}$ $=(x_{2}-x_{3},x_{0}-x_{1}).$
+Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}-ux_{3}(x_{3}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite:
+$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}2ux_{0}&1&2ux_{2}&2ux_{3}\\ x_{3}(x_{2}-x_{3})(1+ux_{0})&0&x_{3}((x_{0}-x_{1})-2ux_{2}(x_{3}-x_{2}))&(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2})\end{pmatrix}$
+Thus, it sends $e_{2}$ to $(1,0)$ and compute the cokernel by considering the remaining columns of the second row, namely
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{I}$
+with $I=(f_{1},f_{2},f_{3})$ and
+$f_{1}$ $=x_{3}(x_{2}-x_{3})$
+$f_{2}$ $=x_{3}((x_{0}-x_{1})-2ux_{3}(x_{3}-x_{2}))$
+$f_{3}$ $=(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2}).$
+Since $f_{1}\in I$, $x_{3}(x_{2}-x_{3})\in I$, and we may rewrite $f_{2}$ as
+$f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\in I\Rightarrow x_{3}(x_{0}-x_{1})\in I,$
+and therefore $x_{3}\in I$. Now, turning to $f_{3}\in I$ and removing terms with $x_{3}$, we obtain:
+$f_{3}\equiv x_{2}(x_{0}-x_{1})\Rightarrow x_{2}\in I,$
+since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
+Over $\mathfrak{p}_{2}=((x_{0}-x_{1}),x_{3})$, we may choose the same elementary operations and thus
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{I}$
+with $I=(f_{1},f_{2},f_{3})$, the polynomials as before. Since $f_{1}\in I$, we conclude:
+$f_{3}=(x_{0}-x_{1})(x_{2}-2x_{3})+2ux_{3}^{2}(x_{2}-x_{3})\Rightarrow(x_{0}-x_{1})(x_{2}-2x_{3})\in I.$
+Since $(x_{2}-2x_{3})$ is invertible in $R_{\mathfrak{p}_{2}}$, we obtain $(x_{0}-x_{1})\in I$.
+From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+Over the prime $\mathfrak{p}_{3}$, we may also do the same elementary operations and consider the same generators for the ideal $I$.
+From $f_{1}\in I$ we get $(x_{2}-x_{3})\in I$, since $x_{3}$ is now invertible.
+Moreover, from $f_{2}\in I$ we obtain:
+$f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\Rightarrow x_{3}(x_{0}-x_{1})\in I\Rightarrow(x_{0}-x_{1})\in I,$
+so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$.
+Thus, from the associativity formula $m(\sigma)=3$.
+Moreover, $e=2$ and thus $\operatorname{Bour}(\sigma)=2$.
+Here, the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
-so that $\operatorname{Bour}(\sigma)=2$ and $c_{3}({\mathcal{T}}_{\sigma})=4$.
+so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane smooth quadric.
 Corollary 60.
-Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
+Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
 Proof.
-Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is stable if and only if $e>1$.
-From the result above, the only possibilities of $\operatorname{Bour}(\sigma)\leq 2$ are when $m(\sigma)\geq 3$, and moreover if $m(\sigma)=3$ we cannot have $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
-For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Section 4, $(a)$, $e\neq 1$.
+Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable if and only if $e>1$.
+For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Theorem B Theorem B, $(a)$, $e\neq 1$.
 ∎
@@
 Proof.
-If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [14, Section 9].
-By the proof of [14, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
+If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [undefn, Section 9].
+By the proof of [undefn, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
 From Remark 6, we obtain $p_{a}(B)=1$, and thus $B$ must be a plane cubic curve, with a resolution of the form
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [14, Lemma $9.3$].
+From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [undefn, Lemma $9.3$].
 If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7).
-From [3, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$.
+From [undefb, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$.
 ∎
@@
 $\nabla\sigma=\begin{pmatrix}-x_{1}&x_{2}-x_{0}&x_{1}-x_{3}&-x_{2}\\ x_{1}^{2}&2x_{0}x_{1}&3x_{2}^{2}+2x_{2}x_{3}&x_{2}^{2}\end{pmatrix}$.
-Here the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the codimension two part of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of a line $V(x_{1},x_{2})$ with multiplicity two.
-The saturation of the first Fitting ideal is $(0)$ in this case.
-The sheaf ${\mathcal{T}}_{\sigma}$ admits the following free resolution:
+The irreducible components of $(\Xi_{\sigma})_{1}$ consist only of a line $V(x_{1},x_{2})$.
+Consider $\mathfrak{p}=(x_{1},x_{2})$ the corresponding prime ideal in $R$, and let $u=(x_{2}-x_{0})^{-1}\in R_{\mathfrak{p}}$.
+The elementary operations $R_{2}^{\prime}=R_{2}-2ux_{0}x_{1}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ turn the matrix into
+$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&1&u(x_{1}-x_{3})&-ux_{2}\\ x_{1}^{2}+2ux_{0}x_{1}^{2}&0&x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})&x_{2}^{2}+2ux_{0}x_{1}x_{2}\end{pmatrix}$,
+which sends $e_{2}\in R_{\mathfrak{p}}^{4}$ to $(1,0)\in R_{\mathfrak{p}}^{2}$, so we may compute the cokernel $(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ by considering the second row and the remaining entries:
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I},$
+where
+$f_{1}$ $=x_{1}^{2}(1+2ux_{0})$
+$f_{2}$ $=x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})$
+$f_{3}$ $=x_{2}(x_{2}+2ux_{0}x_{1}).$
+Since $1+2ux_{0}\in R_{\mathfrak{p}}^{\times}$ and $f_{1}\in I$, we conclude $x_{1}^{2}\in I$. This simplifies $f_{2}\in I$ to:
+$x_{2}(3x_{2}+2x_{3})\equiv-(2ux_{0}x_{3})x_{1}\mod I,$
+so there is a unit $v\in R_{\mathfrak{p}}^{\times}$ such that $x_{1}\equiv vx_{2}\mod I$, which together with $x_{1}^{2}\in I$ implies $x_{2}^{2}\in I$. From this, $f_{3}\in I$ simplifies to
+$f\equiv 2ux_{0}x_{1}x_{2}\in I\Rightarrow x_{1}x_{2}\in I,$
+and therefore we have
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{1}\rangle,$
+with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=2$.
+Thus, $m(\sigma)=2$ (from the associativity formula), and $e=2$, yielding $\operatorname{Bour}(\sigma)=3$.
+The following resolution may also be obtained computationally:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
-so that $e=2$, $\operatorname{Bour}(\sigma)=3$ and $m(\sigma)=2$, with $c_{3}({\mathcal{T}}_{\sigma})=7$.
+describing the Bourbaki scheme as a twisted cubic and giving $c_{3}({\mathcal{T}}_{\sigma})=7$.
 Example 63 ($m(\sigma)=2$, $e=3$, $\operatorname{Bour}(\sigma)=5$, $c_{3}=3$).
@@
 $\nabla\sigma=\begin{pmatrix}-x_{1}&-x_{0}&x_{3}&x_{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}+x_{3}^{2}&x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2}\end{pmatrix}$.
-Here, the saturation of the 0-th Fitting ideal coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $(1)$.
-Moreover, $m(\sigma)=2$ with $\operatorname{supp}(\mathcal{Q}_{\sigma})$ being a double line structure at $V(x_{0},x_{3})$.
-Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
+Here, there is a unique irreducible component of $(\Xi_{\sigma})_{1}$, the line $V(x_{0},x_{3})$.
+Considering $\mathfrak{p}=(x_{0},x_{3})$ the associated prime ideal, we compute $m(\sigma)=2$ using the associativity formula.
+Take $u=(x_{2})^{-1}\in R_{\mathfrak{p}}$ and consider the elementary operations $R_{2}^{\prime}=R_{2}-u(x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2})R_{1}\doteq R_{2}-uQR_{1}$, $R_{1}^{\prime}=uR_{1}$ so that
+$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&-ux_{0}&ux_{3}&1\\ 2x_{0}x_{2}+x_{1}x_{3}+uQx_{1}&x_{0}x_{3}+uQx_{0}&x_{0}^{2}+x_{3}^{2}-uQx_{3}&0\end{pmatrix}$.
+Arguing as usual, we may compute the cokernel by
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I}$
+with
+$f_{1}$ $=x_{0}(2x_{2}+ux_{1}^{2})+x_{1}x_{3}(2ux_{2}+3ux_{3}+1)$
+$f_{2}$ $=x_{0}(x_{3}(2ux_{2}+3ux_{3}+1)+ux_{0}x_{1})$
+$f_{3}$ $=x_{0}^{2}+x_{3}(x_{3}-ux_{0}x_{1}-2ux_{2}x_{3}-3ux_{3}^{2})$
+Writing
+$f_{2}-\frac{x_{0}}{x_{1}}f_{1}$ $=x_{0}\left(-\frac{x_{0}}{x_{1}}(2x_{2}+ux_{1}^{2})+ux_{0}x_{1}\right)$
+$=x_{0}^{2}\left(-\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)\in I,$
+and since the term inside parenthesis is invertible in $R_{\mathfrak{p}}$, we conclude $x_{0}^{2}\in I$.
+On the other hand, we consider
+$B\doteq f_{3}-x_{0}^{2}+\frac{x_{3}}{x_{1}}f_{1}$ $=-x_{3}\left(x_{0}\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{0}x_{1}-2x_{3}\right)$
+$=-x_{3}x_{0}\left(\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)+2x_{3}^{2}\in I.$
+and moreover, let
+$A\doteq x_{0}x_{3}(2ux_{2}+3ux_{3}+1)=f_{2}-ux_{0}^{2}x_{1}x_{3}\in I.$
+Then
+$A-\frac{3}{2}ux_{0}B$ $=x_{0}x_{3}(2ux_{2}+1)+3x_{3}^{2}-3x_{3}^{2}+\frac{3}{2}ux_{0}^{2}x_{3}\left(\frac{-(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right),$
+and after removing the term with $x_{0}^{2}$ factor, we obtain
+$x_{0}x_{3}(2ux_{2}+1)\in I\Rightarrow x_{0}x_{3}\in I,$
+since $2ux_{2}+1$ is invertible in $R_{\mathfrak{p}}$.
+Now, removing the $x_{0}x_{3}$ term in $B$ we conclude $x_{3}^{2}\in I$.
+So, up until now, we have $x_{0}^{2},x_{0}x_{3}$ and $x_{3}^{2}\in I$.
+Finally, considering
+$f_{1}-3ux_{1}x_{3}^{2}=x_{0}(2x_{2}+ux_{1}^{2})+x_{3}x_{1}(2ux_{2}+1)\in I,$
+this means there are $a,b\in R_{\mathfrak{p}}^{\times}$ such that $ax_{0}+bx_{3}\in I$. Therefore, we conclude that
+$I=(x_{0}^{2},x_{3}^{2},x_{0}x_{3},ax_{0}+bx_{3}),$
+and these mean that we can write
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{0}\rangle\simeq\langle 1,x_{3}\rangle$
+and obtain that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=1$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 3}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
@@
 We summarize the results of this subsection in the following theorem:
-Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
+Theorem D.
+Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
   • (a)
@@
   • (d)
-    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
+    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$.
+Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
+  • (e)
+    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
 Proof.
-Item $(a)$ is 46 and item $(b)$ follows with 51. Item $(c)$ is shown in 53, and the stability result in $(d)$ is in 58. ∎
+Item $(a)$ is 48 and item $(b)$ follows with 53. Item $(c)$ is shown in $5$, and item $(d)$ is described in 57. The stability result in $(e)$ is in 60. ∎

### v4

@@
 Abstract
-The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials — defined as the kernel of the associated Jacobian matrix — naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation.
-As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
-The present work focuses on sequences of two homogeneous polynomials in four variables.
+The present work focuses on studying the logarithmic tangent sheaf associated with sequences of two homogeneous polynomials in four variables.
 We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis.
@@
 We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants.
-Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
-We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly free cases.
-In particular, one of the nearly free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
+Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension-one foliation in $\mathbb{P}^{3}$.
+We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic.
+In particular, one of the nearly-free examples induces an unstable, non-split tangent sheaf for a codimension-one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
 1 Introduction
-Let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$ variables with coefficients in an algebraically closed field $\kappa$.
-For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{1}+1,\ldots,d_{k}+1$, respectively, with $d_{1}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
+Let $\kappa$ be an algebraically closed field of characteristic zero and let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$.
+For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{f}+1,\ldots,d_{k}+1$, respectively, with $d_{f}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
 $\nabla\sigma=\begin{pmatrix}\nabla f_{1}\\ \nabla f_{2}\\ \vdots\\ \nabla f_{k}\end{pmatrix}:\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i}).$
-The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$.
-This definition is in analogy (see [undefj]) with the case of the tangent logarithmic sheaf associated to divisors in $\mathbb{P}^{n}$, namely for $k=1$.
-A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
-In the case of divisors on $\mathbb{P}^{n}$, where $k=1$, one has a short exact sequence of the form:
+The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the logarithmic tangent sheaf associated with the sequence $\sigma$.
+This definition is in analogy (see [11]) with the case of divisors in $\mathbb{P}^{n}$, namely for $k=1$.
+A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a direct sum of line bundles on $\mathbb{P}^{n}$.
+In the case of divisors, one has a short exact sequence of the form:
 $0\rightarrow{\mathcal{T}}_{f}\rightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\mathcal{I}_{J_{f}}(d)\rightarrow 0,$
-where $J_{f}=\langle\partial_{0}f,\ldots,\partial_{n}f\rangle\trianglelefteq R$ is the Jacobian ideal of the homogeneous polynomial $f$.
-The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated to the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$.
-For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated to the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
-For the generalization of $k>1$, one should consider sequences rather than the associated ideals $I=(f_{1},\ldots,f_{k})$, as pointed out by [undefj, Example 2.7].
-When $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and furthermore ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [undefj, Lemma 2.14]).
-We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$, called the initial degree.
-Following the terminology of [undefj], we call a sequence $\sigma=(f_{1},\ldots,f_{k})$ compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$.
-This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$.
-Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [undefj, Lemma 2.8]).
-One may also consider the relationship with Bourbaki ideals, as in [undefr] and [undeff].
+where $J_{f}=(\partial_{0}f,\ldots,\partial_{n}f)\subseteq R$ is the Jacobian ideal of the homogeneous polynomial $f$.
+The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated with the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$.
+For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated with the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
+We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a nonzero syzygy of the matrix $\nabla\sigma$, called the initial degree.
 For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
@@
 The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice.
-In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [undefr].
+In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [17].
 By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence.
-The focus of this work is to understand the concept above for $n=3$ and $k=2$.
-For $k=1$, the singular scheme of $V(f)$, defined by the Jacobian ideal $\Sigma_{f}\doteq V(J_{f})\subset\mathbb{P}^{n}$, plays an important role in this study.
-If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
-$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma)}{s!}t^{s}+O(t^{s-1}),$
-and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above.
+The goal of this work is to study the concept above for $n=3$ and $k=2$.
+For $k=1$, the singular scheme $\Sigma_{f}$ of $V(f)$, defined by the Jacobian ideal, plays an important role. If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
+$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma_{f})}{s!}t^{s}+O(t^{s-1}),$
+and, by definition, the leading coefficient $\deg(\Sigma_{f})$ is the degree of the singular scheme.
 Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$.
-For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [undefr]):
-$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}),$
-where $e=\operatorname{indeg}({\mathcal{T}}_{f})$.
-Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for the logarithmic sheaf ${\mathcal{T}}_{f}$ to get an ideal sheaf as the cokernel.
-This problem is addressed in a recent work [undeff] where the authors consider another kind of Bourbaki sequences, recently developed in [undefq], by choosing generators of maximal degree.
-We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$.
-In [undefj], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$.
-In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$.
-The generic case is when $\dim(\mathcal{Q}_{\sigma})\leq n-2$, and in particular $c_{1}(\mathcal{Q}_{\sigma})=0$.
-In this case, we call $\sigma$ a normal sequence.
-For now on, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$.
-We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$.
-Since the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
-If $\sigma$ is a normal sequence, $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
-$H(\mathcal{Q}_{\sigma},t)=\deg(\mathcal{Q}_{\sigma})t+O(1),$
-We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional.
-From the associativity formula applied to the graded homogeneous module $\mathcal{Q}$ associated to $\mathcal{Q}_{\sigma}$, we obtain the expression (see, for example, [undefs, Theorem $14.7$])
-$m(\sigma)=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(R/P)=2}\operatorname{{length}}_{R_{P}}\mathcal{Q}_{P}\cdot\deg(R/P).$
-Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [undefr]. We show a formula for the Bourbaki degree of $\sigma$ given by
-$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma),$
-so that $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. A first result of the paper is some bounds for the discrete quantities defined above.
-Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
+For $n=2$, this means singularities are isolated, and $\deg(\Sigma_{f})$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$.
+The formula of the Bourbaki degree in this case is ([17, Theorem 2.1]):
+$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}).$
+We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix and call the scheme $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$ the Jacobian scheme of $\sigma$, given by the zero locus of the $(k\times k)$-minors of the matrix.
+The set-theoretical support of $\mathcal{Q}_{\sigma}$ is the reduced locus $(\Xi_{\sigma})_{\text{red}}$, although the scheme structure of both may differ.
+When $\dim(\mathcal{Q}_{\sigma})\leq n-2$, we call $\sigma$ a normal sequence.
+Hereafter, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$.
+We also assume sequences $\sigma$ are normal, as defined above, so $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$.
+Then $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
+$H(\mathcal{Q}_{\sigma},t)=m(\sigma)t+O(1),$
+where $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$ and $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional.
+From the associativity formula ([18, Theorem $14.7$]) applied to the graded homogeneous module $Q$ associated with $\mathcal{Q}_{\sigma}$, we obtain the expression
+$m(\sigma)=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ann}}(\mathcal{Q}_{\sigma}),\dim(R/\mathfrak{p})=2}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}Q_{\mathfrak{p}}\cdot\deg(R/\mathfrak{p}).$
+Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [17]. We show the following formula for the Bourbaki degree of $\sigma$:
+Proposition: Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$.
+If $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, then the Bourbaki degree of $\sigma$ is given by
+$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma).$
+A first result of the paper is some bounds for the discrete quantities defined above.
+Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
   • (a)
@@
     • (4)
-      $\sigma$ is compressible;
-We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 8): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable.
-In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely by the matrices $\nabla\sigma$ for which the Buchsbaum–Rim complex is a minimal resolution, or equivalently $m(\sigma)=0$.
-Geometrically, all regular pencils ($d_{f}=d_{g}$) with isolated singularities satisfy this property (see Remark 9).
-In Section 2.2, we relate our notion to the notion of the Bourbaki degree for a projective plane curve, by considering a reduced polynomial $g\in\kappa[x_{0},x_{1},x_{2}]$ together with its associated surface $V(g)\subset\mathbb{P}^{3}$ and projective curve $X=V(g)\subset\mathbb{P}^{2}$, to obtain $\operatorname{Bour}(X)=\operatorname{Bour}(g,x_{3})$.
-Afterwards, in Section 2.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal syzygy.
-Related to the free resolutions, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
-$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1.$
-The converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$. We point out nearly free sequences are never locally free, as we show in 19.
-In Section 3, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [undefj, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections.
-For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [undefr, Corollary 2.11]).
+      $\sigma$ is compressible, that is, up to a linear change of coordinates of $\mathbb{P}^{3}$, both polynomials in $\sigma$ are independent of some variable $x_{0},\ldots,x_{3}$;
+We also show inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 9).
+In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely whenever $m(\sigma)=0$.
+In this case, the cokernel $\mathcal{Q}_{\sigma}$ is minimally resolved by the Buchsbaum–Rim complex.
+In particular, ${\mathcal{T}}_{\sigma}(d)\simeq\operatorname{coker}(\nabla\sigma)^{T}$.
+In Section 2.2, we consider sequences $\sigma=(x_{3},g)$, where $g\in\kappa[x_{0},x_{1},x_{2}]$ is a reduced homogeneous polynomial.
+We show that $\operatorname{Bour}(\sigma)$ coincides with the Bourbaki degree of the plane curve defined by $V(g)\subset\mathbb{P}^{2}=V(x_{3})$.
+In Section 2.3, we relate free resolutions of ${\mathcal{T}}_{\sigma}$ and of the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal degree syzygy.
+As a consequence, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
+$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
+where $\operatorname{gpdim}({\mathcal{T}}_{\sigma})$ is the minimal length of a free resolution of the $R$-module $H^{0}_{*}({\mathcal{T}}_{\sigma})=\bigoplus_{l\in\mathbb{Z}}H^{0}({\mathcal{T}}_{\sigma}(l))$.
+The converses do not hold, see Example 16 and Example 23.
+In Section 3, we explore the structure of codimension-one foliation associated with $\sigma$, presented in [11, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$.
+For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [17, Corollary 2.11]).
 The main theorem of the section is:
@@
     If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
-In Section 4, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$.
+We note that the bounds above are independent of the degrees $d_{f},d_{g}$.
+In Section 4, we study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$.
 To do so, we use classical results on bounds for Chern classes of reflexive sheaves in $\mathbb{P}^{3}$.
-We classify all free and nearly free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized below:
-Theorem C.(Theorem C)Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
-  • (a)
-    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
-  • (b)
-    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
-  • (c)
-    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42), which is strictly $\mu$-semistable;
-  • (d)
-    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
-In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 15), but the bound given above for stability may not be sharp.
-Theorem D.(Theorem D) Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
-  • (a)
-    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
-  • (b)
-    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
-  • (c)
-    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 56 and Example 14);
-  • (d)
-    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$.
-Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
-  • (e)
-    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
-Example 14 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable.
-This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [undefa]:
-Conjecture. If the tangent sheaf of a codimension one foliation on $\mathbb{P}^{3}$ is not split, then it is $\mu$-semistable.
-Along the text, we describe examples developed computationally with aid of Macaulay2 software ([undefl]) for the two families of sequences.
+We classify all free and nearly-free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized at the tables below (for precise statements, see Theorem Theorem C and Theorem Theorem D).
+Finally, we provide a counter-example for a conjecture posed by Calvo-Andrade, Correa and Jardim, in [3].
+They proposed that a tangent sheaf of a codimension-one foliation on $\mathbb{P}^{3}$ is either split or $\mu$-semistable.
+Example 15 gives a sequence $\sigma$ which induces a codimension-one foliation whose tangent sheaf is neither split nor $\mu$-semistable.
+Along the text, we describe examples developed computationally with aid of Macaulay2 software ([13]) for the two families of sequences.
+Classification results for $d_{f}=d_{g}=2$.
+$m(\sigma)$ $\operatorname{Bour}(\sigma)$ Behavior minimal resolution for logarithmic tangent sheaf
+12 0 free, compressible $\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)$
+9 0 free, unstable $\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$
+8 0 free, $\mu$-semistable $\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}$
+7 1 unique nearly-free $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0$
+$\leq 6$ $\geq 2$ $\mu$-semistable
+$\leq 2$ $\geq 6$ $\mu$-stable
+0 12 $\mu$-stable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\xrightarrow{(\nabla\sigma)^{T}}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0$
+Classification results for $d_{f}=1$, $d_{g}=2$.
+$m(\sigma)$ $\operatorname{Bour}(\sigma)$ Behavior minimal resolution for logarithmic tangent sheaf
+7 0 free, compressible $\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$
+5 0 free, unstable $\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$
+4 1 nearly-free, unstable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-3)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\to{\mathcal{T}}_{\sigma}\to 0$
+4 1 nearly-free, $\mu$-stable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0$
+<3<3 $\geq 3$ $\mu$-stable
+0 7 $\mu$-stable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{(\nabla\sigma)^{T}}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0$
 Acknowledgments
@@
 In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials.
-We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3), in particular introducing the class of nearly free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
+We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3).
 2.1 Framework and first results
@@
 and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$.
-We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [undefj].
+A sequence $\sigma$ will be called free if ${\mathcal{T}}_{\sigma}$ splits as a direct sum of line bundles.
+We assume throughout that the sequences $\sigma$ satisfy the condition $\operatorname{{codim}}(\mathcal{Q}_{\sigma})\geq 2$, and in this case we call $\sigma$ a normal sequence.
+With this assumption, we obtain $\operatorname{{rk}}(\mathcal{Q}_{\sigma})=0$ and $c_{1}(\mathcal{Q}_{\sigma})=0$.
 The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
@@
 Proof.
-Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
+This is a direct application of the Hirzebruch-Riemann-Roch Theorem. ∎
 We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$.
-From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
-As in the introduction, the Jacobian scheme of $\nabla\sigma$ is defined as
+Since $m(\sigma)$ is the leading coefficient of a Hilbert polynomial, $m(\sigma)\geq 0$, and it is zero if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
+The Jacobian scheme of $\nabla\sigma$ is defined as
 $\Xi_{\sigma}\doteq V\left(\bigwedge^{2}\nabla\sigma\right)=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
 the zero-locus of the $2\times 2$ minors of $\nabla\sigma$ or, in other words, the $0$-th Fitting scheme of $\nabla\sigma$.
-From the general theory of Fitting ideals (see, for example, [undefh, Chapter 20]), the annihilator ideal of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal, and they have the same support, namely
-$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
-as closed subsets, and therefore the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$.
-In general, however, their schematic structure may be different, as we explore in the following example.
+From the general theory of Fitting ideals, the annihilator ideal $\operatorname{\operatorname{Ann}}\mathcal{Q}_{\sigma}$ of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal $\operatorname{Fitt}_{0}(\nabla\sigma)$, and they have the same reduced support, namely
+$V(\sqrt{\operatorname{\operatorname{Ann}}(\mathcal{Q}_{\sigma}})=\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))_{\text{red}}=V(\sqrt{\operatorname{Fitt}_{0}(\nabla\sigma)})$
+as closed subsets of $\mathbb{P}^{3}$.
+In particular, the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$.
+The respective schematic structures may be different, as we explore in the next example.
 Denote by $(\Xi_{\sigma})_{1}$ the one-dimensional part of the scheme $\Xi_{\sigma}$.
@@
 Let $\sigma=(2x_{1}x_{3}-x_{1}^{2},3x_{2}x_{3}^{2}-3x_{0}x_{1}x_{3}+x_{1}^{3})$. The Jacobian matrix is of the form
-$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$
-and the matrix below
+$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$.
+The annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, namely given by:
+$\begin{cases}\operatorname{\operatorname{Ann}}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
+so both schemes $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ are non-reduced, with degrees 44 and $6$, respectively. The matrix below
 $\begin{pmatrix}x_{3}&x_{0}x_{1}-x_{1}^{2}-2x_{2}x_{3}\\ 0&-x_{1}x_{3}\\ x_{1}&-2x_{2}x_{3}\\ 0&-x_{1}x_{3}+x_{3}^{2}\end{pmatrix}$
-gives two linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
-In this case, the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, given by:
-$\begin{cases}\operatorname{Ann}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
-so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively.
-From the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$, since ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ we conclude $m(\sigma)=5$.
-The right approach to computing $m(\sigma)$ in general is using the associativity formula, as we will in the next examples.
-Any saturated syzygy of the Jacobian matrix $\nu$, of degree $e\in\mathbb{Z}$, induces a short exact sequence:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
-where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described as the Serre correspondence (see [undefn, Theorem $4.1$]) between the sheaf ${\mathcal{T}}_{\sigma}$ and the scheme $B$.
-Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$.
-Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$.
-The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
+gives two linearly independent syzygies for the matrix $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ and $\sigma$ is free.
+From this and the additivity of the Hilbert polynomial on short exact sequences, one may compute $m(\sigma)=5$.
+The following proposition gives a sufficient condition to forbid the anomaly above to occur.
 Proposition 3.
-Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
-For any saturated syzygy $\nu\in H^{0}({\mathcal{T}}(e))$ of degree $e\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated closed subscheme of pure dimension one.
+Let $\sigma=(f,g)$ be a normal sequence. If $\operatorname{{codim}}(\operatorname{Fitt}_{1}(\nabla\sigma))\geq 3$, then
+$m(\sigma)=\deg(\Xi_{\sigma})=\deg(\operatorname{supp}\mathcal{Q}_{\sigma}).$
+Proof.
+Let $Q_{\sigma}$ denote the graded $R$-module associated to $\mathcal{Q}_{\sigma}$. All three invariants above may be described using the associativity formula, as follows:
+$\deg(\Xi)$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(\operatorname{Fitt}_{0}(\nabla\sigma)),\;\dim(R/\mathfrak{p})=\dim(\operatorname{Fitt}_{0}(\nabla\sigma)_{\mathfrak{p}})}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}\bigl((R/\operatorname{Fitt}_{0}(\nabla\sigma))_{\mathfrak{p}}\bigr)\cdot\deg(R/\mathfrak{p})$
+$\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(\operatorname{Fitt}_{0}(\nabla\sigma)),\;\dim(R/\mathfrak{p})=\dim((Q_{\sigma})_{\mathfrak{p}})}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}\bigl((R/\operatorname{\operatorname{Ann}}Q_{\sigma})_{\mathfrak{p}}\bigr)\cdot\deg(R/\mathfrak{p})$
+$m(\sigma)$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(\operatorname{Fitt}_{0}(\nabla\sigma)),\;\dim(R/\mathfrak{p})=\dim((Q_{\sigma})_{\mathfrak{p}})}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}\bigl((Q_{\sigma})_{\mathfrak{p}}\bigr)\cdot\deg(R/\mathfrak{p}).$
+The sets of primes involved in each sum are the same, so we only need to check that the lengths coincide over minimal prime ideals $\mathfrak{p}$ of height two.
+Since $\operatorname{{codim}}(\operatorname{Fitt}_{1}(\nabla\sigma))\geq 3$, the rank of the localization $(\nabla\sigma)_{\mathfrak{p}}$ is one, and we may change coordinates locally to write
+$A_{\mathfrak{p}}\doteq(\nabla\sigma)_{\mathfrak{p}}=\begin{pmatrix}1&0&0&0\\ 0&a&b&c\end{pmatrix}$.
+In these coordinates, we obtain $\operatorname{Fitt}_{0}(A_{\mathfrak{p}})=(a,b,c)$ from the $2\times 2$-minors, and moreover $(Q_{\sigma})_{\mathfrak{p}}\simeq R_{\mathfrak{p}}/(a,b,c)$.
+We claim that $\operatorname{\operatorname{Ann}}(Q_{\sigma})_{\mathfrak{p}}=(a,b,c)$.
+Since $(a,b,c)=\operatorname{Fitt}_{0}(\nabla\sigma)\subset\operatorname{\operatorname{Ann}}(Q_{\sigma})$, we show the other inclusion.
+Let $r\in R_{\mathfrak{p}}$ such that $r(Q_{\sigma})_{\mathfrak{p}}=0$.
+Writing $(Q_{\sigma})_{\mathfrak{p}}=\operatorname{coker}(A_{\mathfrak{p}})$, we have
+$r\cdot\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}=\begin{pmatrix}rv_{1}\\ rv_{2}\end{pmatrix}\in\text{im }A_{\mathfrak{p}}$
+for $(v_{1},v_{2})\in R_{\mathfrak{p}}^{2}$.
+From the second equation, one obtains $r\cdot v_{2}=au_{1}+bu_{2}+cu_{3}$, for some $u_{i}\in A_{p}$.
+Taking $v_{2}=1$ yields $r\in(a,b,c)$.
+This shows that
+$\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}}/\operatorname{Fitt}_{0}(\nabla\sigma)_{\mathfrak{p}})=\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}(R_{\mathfrak{p}}/\operatorname{\operatorname{Ann}}(Q_{\sigma})_{\mathfrak{p}})=\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}((Q_{\sigma})_{\mathfrak{p}}),$
+and since $(Q_{\sigma})_{\mathfrak{p}}\simeq R_{\mathfrak{p}}/(a,b,c)$, the claim follows. ∎
+Given a sequence $\sigma=(f,g)$ and a saturated syzygy of the Jacobian matrix $\nu$ of degree $l\in\mathbb{Z}$, we obtain a short exact sequence:
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
+where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated with $\nu$. Using the Hilbert polynomial, one can write formulas (see [14, Proposition $4.1$]):
+$\deg(B_{\nu})$ $=c_{2}({\mathcal{T}}_{\sigma}(l))$
+$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2l).$
+Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $l\geq 0$.
+The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $l$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
+Proposition 4.
+Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
+For a saturated syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(l))$ of degree $l\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated curve.
 Then, we have the following equation:
-$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+m_{0}-m(\sigma),$
+$\deg(B_{\nu})=l^{2}-l(d_{f}+d_{g})+m_{0}-m(\sigma),$
 where $m_{0}\doteq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$.
@@
 We can relate the Chern character and the Chern classes by the formula
-$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
+$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
+$=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
 $=\frac{(d_{f}+d_{g})^{2}}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}-c_{2}({\mathcal{T}}_{\sigma})$
@@
 $c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma).$ (1)
-On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(e))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
-$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(e))=c_{2}({\mathcal{T}}_{\sigma})+e\cdot c_{1}({\mathcal{T}}_{\sigma})+e^{2}.$
+On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(l))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
+$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(l))=c_{2}({\mathcal{T}}_{\sigma})+l\cdot c_{1}({\mathcal{T}}_{\sigma})+l^{2}.$
 From this, together with 1, we obtain
-$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
-We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [undefr, Definition 2.4].
+$\deg(B_{\nu})=l^{2}-l(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
+We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [17, Definition 2.4].
 Definition 5.
-Let $\sigma$ be a normal sequence and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. The Bourbaki degree of a normal sequence $\sigma$ is defined by:
+Let $\sigma=(f,g)$ be a normal sequence, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $d\doteq d_{f}+d_{g}$. The Bourbaki degree of $\sigma$ is defined by:
 $\operatorname{Bour}(\sigma)\doteq\deg(B_{\nu})=e(e-d)+m_{0}-m(\sigma),$
-for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. We note that every non-trivial syzygy of minimal degree is saturated.
+for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$.
+We note that every non-trivial syzygy of minimal degree is saturated, and $\operatorname{Bour}(\sigma)$ is independent of the choice of $\nu$.
 Remark 6.
@@
 $0\to\mathcal{O}_{\mathbb{P}^{3}}(-e)\to{\mathcal{T}}_{\sigma}\to\mathcal{O}_{\mathbb{P}^{3}}(e-d)\to 0$
-splits.
-Remark 6.
-From [undefm, Proposition 4.1], we have the formula
-$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
-relating the third Chern class of ${\mathcal{T}}_{\sigma}$ and the discrete invariants of $B$, when $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $B$ is the zero locus of a non-zero section in $H^{0}({\mathcal{T}}_{\sigma}(e))$.
-Moreover, dualizing the sequence
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B}(e-d)\rightarrow 0$
-we conclude that the singular set of the sheaf ${\mathcal{T}}_{\sigma}$ is contained in $B$.
+splits, as the extension group $\operatorname{{Ext}}^{1}(\mathcal{O}_{\mathbb{P}^{3}}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(d-2e))$ vanishes.
+Whenever, up to a change of coordinates, the elements of a sequence $\sigma=(f,g)$ do not depend on some variable in $R$, $\sigma$ is called compressible, see [11].
+The following lemma is a consequence of their results in our context:
+Lemma 7.
+[[11, Lemmas 2.7 and 2.8]] Let $\sigma=(f,g)$ be a normal sequence in $\mathbb{P}^{3}$. Then:
+  • (a)
+    A sequence $\sigma$ is compressible if and only if $h^{0}({\mathcal{T}}_{\sigma})\neq 0$;
+  • (b)
+    If $\sigma$ is compressible, then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
 Now, we obtain some bounds for the quantities $\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ in terms of the degrees $d_{f},d_{g}$.
@@
   • (a)
-    We have $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
+    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
   • (b)
@@
 are a syzygies of degree $d_{f}+d_{g}$.
-Since there is at least one nonzero $2\times 2$ minor, from the hypothesis of algebraically independent, at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
+Since $\sigma$ is algebraically independent, there is at least one nonzero $2\times 2$ minor, and at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
 For $(b)$, using that $e\leq d$, we can use that
@@
 so that $m(\sigma)\leq m_{0}+e(e-d)$, but $e(e-d)\leq 0$ and the claim follows.
-For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [undefj, Lemma 2.7].
-Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [undefj, Lemma 2.8].
+For $(c)$, the equivalence $(2)\iff(3)$ is the content of 7, $(a)$.
+Moreover, $(3)\Rightarrow(4)$ is 7, $(b)$.
 The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
@@
 From the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=e(e-d)$.
-Since $\operatorname{Bour}(\sigma)\geq 0$ and from part (a) we have $e\leq d$, it follows that $e=0$ or $e=d$.
-Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free.
-Since $c_{1}({\mathcal{T}}_{\sigma})=-d$ is additive, it follows that $e=0$, otherwise we would have $c_{1}({\mathcal{T}}_{\sigma})<-d$.
+Since $\operatorname{Bour}(\sigma)\geq 0$ and, from part (a), $e(e-d)\leq 0$, it follows that either $e=0$ or $e=d$.
+Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free, say ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-a)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-b)$, with $0\leq a\leq b$.
+If $e=d$, then $a=d>0$, but with $c_{1}({\mathcal{T}}_{\sigma})=-(a+b)=-d$ one gets a contradiction with $a\leq b$, and thus $e=0$.
 ∎
@@
 The bound obtained in Theorem A, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory.
-We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [undefj, Section 9]).
+We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension-one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [11, Section 9]).
 Let $C$ be the one-dimensional part of the singular scheme of this foliation.
-From the formulas of discrete invariants in [undefa, Theorem 3.1] given for generic foliations,
+From the formulas of discrete invariants in [3, Theorem 3.1]:
 $c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C).$
@@
 $\deg(C)-d-d_{f}d_{g}-1=m(\sigma)$ $\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g},$
-and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [undefv, Corollary 4.8].
-From simple observations about the formula $\operatorname{Bour}(\sigma)$, we are able to obtain the following inequalities related to $\mu$-stability and freeness of the logarithmic sheaves.
+and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [21, Corollary 4.8].
+The following inequalities relate the discrete invariants with $\mu$-stability and freeness of the logarithmic sheaves.
 Proposition 9.
-Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
+Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
   • (a)
@@
     • (2)
-      A minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
-      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{1})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{2})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
-      a Buchsbaum–Rim resolution, where
-      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\phi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
-      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$. In particular, $e=d$.
+      ${\mathcal{T}}_{\sigma}$ admits a minimal resolution by the Buchsbaum–Rim complex:
+      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{g})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
+      where
+      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\varphi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
+      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$.
+In particular, $e=d$ and the logarithmic tangent sheaf is identified with the cokernel of the transpose matrix ${\mathcal{T}}_{\sigma}(d)\simeq\operatorname{coker}(\nabla\sigma)^{T}$.
   • (b)
@@
 Proof.
-The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$.
-This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility.
-The converse also follows simply from the formula, assuming that $m(\sigma)=0$ and $e=d$.
-To show the equivalences above, we start by noting that $m(\sigma)=0$ implies that the grade of the ideal of $(2\times 2)$-minors is three, and therefore the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$.
-From this fact, we obtain that $e=d$.
-For the converse, since $e=d$ and the Buchsbaum–Rim complex is a minimal resolution, we obtain that the grade of the ideal of minors of $\nabla\sigma$ is at least three, hence $\dim(\mathcal{Q}_{\sigma})=0$ and thus $m(\sigma)=0$.
-By construction, $m(\sigma)\geq 0$, and therefore for a given value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, the Bourbaki degree of $\sigma$ can be at most
+The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)=e(e-d)+m_{0}-m(\sigma)$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$.
+This also shows equality occurs if and only if both e(ed)e(e-d) and $m(\sigma)$ are zero, therefore $e=d$ or $e=0$.
+But $e=0$ gives $\operatorname{Bour}(\sigma)=0$, from compressibility, and $m_{0}\neq 0$ implies $e=d$.
+To show the equivalences above, note that $m(\sigma)=0$ if and only if the grade of the ideal of $(2\times 2)$-minors is three, which occurs if and only if the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$, when in particular $e=d$.
+To show $(b)$, note that
 $\operatorname{Bour}(\sigma)\leq e(e-d)+m_{0},$
 and this is a function $H=H(e)$ which attains its minimum at $e=d/2$.
-The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, and if $e>d/2$, ${\mathcal{T}}_{\sigma}$ is $\mu$-stable, since $\mu({\mathcal{T}}_{\sigma})=-d/2$.
-Thus, the maximum value $\operatorname{Bour}(\sigma)$ in the range $e\in\{1,\ldots,d/2\}$ for any possible value of $m(\sigma)$ is $H(1)$, which is the expression above on the right-hand side.
-If $\operatorname{Bour}(\sigma)$ is higher than this, then $e>d/2$, and thus we obtain $(b)$.
-With the same strategy as $(b)$, since the function $H(e)$ attains its minimum at $e=d/2$, for $m(\sigma)$ satisfying the inequality of the claim we obtain that
-$\operatorname{Bour}(\sigma)=H(e)\geq H(d/2)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
-and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. ∎
+The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, so when $\operatorname{Bour}(\sigma)>H(1)=(d_{f}-1)(d_{f}+d_{g})+d_{g}^{2}+1$, then $e>d/2$, and thus ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
+To prove $(c)$, note that $H$ attains its minimum at $e=d/2$, so we obtain that
+$\operatorname{Bour}(\sigma)=H(e)-m(\sigma)\geq H(d/2)-m(\sigma)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
+and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and $\sigma$ is not free. ∎
 Remark 10.
-Let $\sigma=(f,g)$ be a sequence of surfaces of the same degree, say $d_{f}=d_{g}=p$, called a pencil of surfaces. It defines a 11-parameter family of projective surfaces as:
+Let $\sigma=(f,g)$ with $d_{f}=d_{g}=p$. It defines a pencil of projective surfaces:
 $V_{\sigma}\doteq\{V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}:z=[z_{0}:z_{1}]\in\mathbb{P}^{1}\},$
-and each hypersurface $V(z_{0}f+z_{1}g)$ is called a member of the pencil $V_{\sigma}$.
-The set where each surface is singular is the vanishing locus of a polynomial $p=p(z)$ (the discriminant) inside $\mathbb{P}^{1}$, so it must be either a finite subset or the whole family.
-When this set is finite, the pencil is called a regular pencil, otherwise is called irregular.
-In [undefj, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}$ for pencils of surfaces as the union of singular loci:
+where each $V(z_{0}f+z_{1}g)$ is a member of $V_{\sigma}$. The pencil is called regular when the generic member is non-singular, otherwise it is called irregular.
+In [11, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ for pencils of surfaces as the union of singular loci:
 $(\Xi_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g)).$
-Hence, whenever a normal sequence $\sigma=(f,g)$ is a regular pencil where every singular member has isolated singularities, we conclude $\dim(\Xi_{\sigma})=0$ from the description above, and thus $m(\sigma)=0$.
+Hence, whenever a normal sequence $\sigma=(f,g)$ defines a regular pencil whose members have at most isolated singularities, then $\dim(\Xi_{\sigma})=0$ and $m(\sigma)=0$.
 Example 11.
 Consider the sequence
-$\sigma=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
+$\sigma=(f,g)=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
 where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$.
-Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-given by the Buchsbaum–Rim complex.
+Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$.
+Using 9, $(a)$, it follows ${\mathcal{T}}_{\sigma}$ has a minimal free resolution given by the Buchsbaum–Rim complex:
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
 2.2 Plane curves as pairs of surfaces
-Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$.
-We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$.
-Denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme, given as the zero locus of the Jacobian ideal of $g$.
+Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free homogeneous polynomial of degree $d$, with the associated curve $X=V(g)\subset\mathbb{P}^{2}=V(x_{3})$ which has at most isolated singularities, and let $\sigma=(x_{3},g)$.
+We consider $S=V(g)\subset\mathbb{P}^{3}$, which is the cone over the curve $X$ and denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme.
 The matrix $\nabla\sigma$ will be given by
@@
 ${\mathcal{T}}_{\overline{g}}\doteq\ker(\nabla\overline{g})\hookrightarrow\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}\xrightarrow{\nabla\overline{g}}\mathcal{O}_{\mathbb{P}^{3}}(d_{g})$
-the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
+the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, we form the following diagram with exact columns:
 ${\mathcal{T}}_{\overline{g}}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}$$\mathcal{O}_{\mathbb{P}^{3}}(d)$$\mathcal{O}_{Z_{g}}(d)$${\mathcal{T}}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}$$\mathcal{O}_{\mathbb{P}^{3}}(d)\oplus\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{Q}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{O}_{\mathbb{P}^{3}}$$\scriptstyle{\nabla\overline{g}}$$\scriptstyle{\nabla\sigma}$$\scriptstyle{\cdot 1}$
-and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$.
-Denoting $i:H\hookrightarrow\mathbb{P}^{2}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely.
-From the discrete invariants considered $(d_{f}=0)$, we have the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
+and by the snake lemma, there are isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$.
+On $\mathbb{P}^{2}$, we have the sequence defining the logarithmic tangent sheaf of the curve $X$:
+$0\to{\mathcal{T}}_{g}\to\mathcal{O}_{\mathbb{P}^{2}}\xrightarrow{(\partial_{0}g,\partial_{1}g,\partial_{2}g)}\mathcal{O}_{\mathbb{P}^{3}}(d)\to 0,$
+and a defined Bourbaki degree $\operatorname{Bour}(X)$ ([17]), which coincides with the Chern class $c_{2}({\mathcal{T}}_{g}(e))$ when $e=\operatorname{indeg}({\mathcal{T}}_{g})$.
+By the block-form of $\nabla\sigma$,
+$e=\operatorname{indeg}({\mathcal{T}}_{g})=\operatorname{indeg}({\mathcal{T}}_{\sigma}).$
+Denoting by $i:\mathbb{P}^{2}\simeq V(x_{3})\hookrightarrow\mathbb{P}^{3}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely.
+Now, compare the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
 $\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}({\mathcal{T}}_{g}(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
-where $\tau(X)$ is the Tjurina number of the curve $X$, since $d_{f}=0$.
-From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:H\hookrightarrow\mathbb{P}^{3}$ and thus
+where $\tau(X)$ is the Tjurina number of the curve $X$. From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$, so that
 $1-d[H]+(m_{0}-m(\sigma))[H]^{2}$ $=i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$
-$=1-d[H]+(d^{2}-\tau(X))[H]^{2},$
-and since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$.
-Moreover, doing the analogous comparison for the total classes of the twist ${\mathcal{T}}_{\sigma}(e)$, we obtain
-$i^{*}c({\mathcal{T}}_{\sigma}(e))=1+(e-d)[H]+\operatorname{Bour}(\sigma)[H]^{2}=1+(e-d)[H]+\operatorname{Bour}(X)[H]^{2}=c(i^{*}{\mathcal{T}}_{\sigma}(e)),$
-and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$.
-Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does.
-A more algebraic way to see this is through the associativity formula.
-Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$, with the effect of adding a variable to the graded ring of the plane $\mathbb{P}^{2}\hookrightarrow\mathbb{P}^{3}$, and the associated map of modules is given by the quotient map
-$0\to(x_{3})\to\kappa[x_{0},x_{1},x_{2},x_{3}]\to\kappa[x_{0},x_{1},x_{2}]\to 0.$
-Then, for any associated prime $P$ of the Jacobian ideal of $X$ at $\kappa[x_{0},x_{1},x_{2}]$, the extension will satisfy
-$\dim R/(R\cdot P)$ $=\dim(\kappa[x_{0},x_{1},x_{2}]/P)+1$
-$\deg R/(R\cdot P)$ $=\deg(\kappa[x_{0},x_{1},x_{2}]/P).$
-since the generators do not include $x_{3}$.
-Moreover, the length of the modules over $P$ will be preserved, as the element $x_{3}\not\in P$ will be invertible.
-Together with the associativity formula, we obtain:
-$\tau(X)$ $=\sum_{P\in\operatorname{Ann}(\mathcal{O}_{\operatorname{Sing}X}(d)),\dim(\kappa[x_{0},x_{1},x_{2}]/P)=1}\operatorname{{length}}(\mathcal{O}_{\operatorname{Sing}X}(d))_{P}\cdot\deg(\kappa[x_{0},x_{1},x_{2}]/P)$
-$=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(\kappa[x_{0},x_{1},x_{2},x_{3}]/P)=2}\operatorname{{length}}(\mathcal{Q}_{\sigma})_{P}\cdot\deg(R/P)=m(\sigma).$
-The following example of the nature above is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ but $m(\sigma)\neq 0$, showing that a converse for 8, $(a)$ is not possible without assuming the Buchsbaum–Rim resolution.
+$=1-d[H]+(d^{2}-\tau(X))[H]^{2}.$
+Since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$.
+Geometrically, this means that $m(\sigma)$ counts the singular lines of the cone $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does.
+Another perspective for this is through the associativity formula.
+Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$ or, in terms of rings, if $S\doteq\kappa[x_{0},x_{1},x_{2}]$, we have
+$\frac{S[x_{3}]}{(\partial_{0}g,\partial_{1}g,\partial_{2}g)}\otimes_{S}\frac{S[x_{3}]}{(x_{3})}\simeq\frac{S}{(\partial_{0}g,\partial_{1}g,\partial_{2}g)},$
+and then coming back to $R$ via a flat change of basis $\otimes_{\kappa}\kappa[x_{3}]$, since $\partial_{i}g\in\kappa[x_{0},x_{1},x_{2}]$.
+Therefore, for any associated prime $\mathfrak{p}$ of the Jacobian ideal $J_{g}\subset S$, the extension satisfies
+$\dim R/(\mathfrak{p}R)$ $=\dim(S/\mathfrak{p})+1$
+$\deg R/(\mathfrak{p}R)$ $=\deg(S/\mathfrak{p}),$
+and each associated prime to the cone ideal $J_{g}R$ comes by an extension to an associated prime of $J_{g}$ in $S$. Now, using the associativity formula, we get:
+$\tau(X)$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(J_{g}),\dim(S/\mathfrak{p})=1}\operatorname{\operatorname{length}}_{S_{\mathfrak{p}}}(S/J_{g})_{\mathfrak{p}}\cdot\deg(S/\mathfrak{p})$
+$=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(J_{g}R),\dim(R/\mathfrak{p})=2}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}(R/J_{g}R)_{\mathfrak{p}}\cdot\deg(R/\mathfrak{p})=m(\sigma).$
+The following example is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ is maximal but $m(\sigma)\neq 0$.
 Example 12.
@@
 Hence, $m(\sigma)=\tau(X)=1\neq 0$, accounting for the nodal singularity, but $e=d=3$.
-2.3 Free resolutions and the Bourbaki scheme
-In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [undefr, Theorem 2.1, (c)].
-The next result follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$, as it is the case for $\mathbb{P}^{2}$.
+2.3 Free resolutions and Bourbaki schemes
+In this section, after a choice of syzygy $\nu$ of minimal degree, we relate the resolutions for the ideal sheaf of its zero-locus $\mathcal{I}_{B_{\nu}}$ and for ${\mathcal{T}}_{\sigma}$.
+We use this lemma to characterize different classes of sequences.
 Lemma 13.
-Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a non-zero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated to $\nu$ in a short exact sequence:
+Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a nonzero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated with $\nu$ in a short exact sequence:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0.$
@@
     $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\xrightarrow{\omega}\mathcal{I}_{B}\rightarrow 0$
-    lifts for a free resolution of the form
-    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0$
-    for ${\mathcal{T}}_{\sigma}$.
+    lifts to a free resolution
+    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0.$
   • (b)
-    For a minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
+    Every minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
     $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\lambda,\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0,$
-    it induces a free resolution for $\mathcal{I}_{B}$ of the form:
+    induces a free resolution of $\mathcal{I}_{B}$:
     $0\rightarrow F_{2}(d-e)\rightarrow F_{1}(d-e)\rightarrow{}F_{0}(d-e)\xrightarrow{\lambda(d-e)}\mathcal{I}_{B}\rightarrow 0.$
 Proof.
+To summarize, item $(a)$ follows from the Horseshoe lemma since line bundles have vanishing $\operatorname{{Ext}}^{1}(-,\mathcal{O}_{\mathbb{P}^{3}})$, and $(b)$ comes from the mapping cone, after factoring the trivial factors $\mathcal{O}_{\mathbb{P}^{3}}(-e)$.
 To show $(a)$, we apply the functor $\operatorname{Hom}(F_{0}(e-d),-)$ to the short exact sequence
@@
 since $\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(F_{0}^{\vee}(-2e-d))=0$, as $F_{0}^{\vee}$ is a direct sum of line bundles and these have vanishing first cohomology in $\mathbb{P}^{3}$.
-Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\omega(e-d)=\tilde{\omega}$.
+Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\tilde{\omega}=\omega(e-d)$.
 We now consider the map $\omega(e-d)\oplus\nu$ in the following commutative diagram with short exact sequences as the central two columns:
@@
 Definition 14.
-Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say $\sigma$ is:
+Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say that $\sigma$ is:
   • nearly free if $\operatorname{Bour}(\sigma)=1$.
-  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of Lemma 12, $(b)$.
-Example 14 (Nearly free sequence with $d_{f}=1,d_{g}=2$).
+  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of 13, $(b)$.
+Example 15 (Nearly-free sequence with $d_{f}=1,d_{g}=2$).
 We consider the following sequence with $d_{f}=1$, $d_{g}=2$:
@@
 The scheme $(\Xi_{\sigma})_{1}$ has three primary components in dimension one, given by the prime ideals $\mathfrak{p}_{1}=(x_{0},x_{3})$, $\mathfrak{p}_{2}=(x_{0},x_{2})$ and $\mathfrak{p}_{3}=(x_{0},x_{1})$.
-We also note that there is an easy syzygy of degree one, namely $\nu=(-x_{3},0,0,x_{0})^{T}$, hence $e=1$.
-To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated to the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
-Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{P}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}_{P}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$.
+We also note that $\nu=(-x_{3},0,0,x_{0})^{T}$ is a syzygy for $\nabla\sigma$ of degree one, so $e\leq 1$.
+We claim that $m(\sigma)=4$, which shows that $\sigma$ cannot be compressible (by Theorem A Theorem A) and hence $e=1$.
+Moreover, with these data, the Bourbaki degree formula yields $\operatorname{Bour}(\sigma)=1$, so this is a nearly-free sequence.
+To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated with the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
+Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{\mathfrak{p}}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$.
 From this, we may rewrite the matrix as
@@
 first by adding the first row scaled by $(-\frac{3}{2}x_{3})$ to the second row, and then multiplying the first row by $u/2$.
-Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{P}^{2}$, and thus
+Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{\mathfrak{p}}^{2}$, and thus
 $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0}x_{1},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
@@
 since $x_{1}\in R_{\mathfrak{p}_{2}}^{\times}$.
-Thus, it follows $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+Thus, it follows $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
 For the ideal $\mathfrak{p}_{3}$, we note that the approach is analogous as the previous one, using the same elementary operations, and obtaining the same isomorphism at the end $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{3}}\simeq\kappa$, so that from the associativity formula:
-$m(\sigma)=\sum_{i=1}^{3}\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4.$
-Using the Bourbaki formula with $m_{0}=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}=7$, $d=3$ and the obtained discrete invariants $m(\sigma)=4$ and $e=1$, we conclude that $\operatorname{Bour}(\sigma)=1$.
-From the minimal free resolution of the ideal of a line in $\mathbb{P}^{3}$, using Lemma 12, the minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-From the resolution, one obtains $c_{3}({\mathcal{T}}_{\sigma})=3$.
+$m(\sigma)=\sum_{i=1}^{3}\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4,$
+as we wanted to show.
 Example 16 ($3$-syzygy pencil of cubics which is not nearly-free).
-Considering the following pencil of cubics:
-$\sigma=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
+Consider the following pencil of cubics:
+$\sigma=(f,g)=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
 where $f$ is a hyperplane arrangement and $g$ is the union of a plane and a smooth quadric, with the Jacobian matrix:
 $\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
-The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals of $R$:
+We claim that here, $m(\sigma)=5$ and that $\operatorname{Bour}(\sigma)=4\neq 1$, so $\sigma$ is not nearly free. Macaulay2 gives the following minimal free resolution for ${\mathcal{T}}_{\sigma}$:
+$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
+which is of $3$-syzygy type. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals in $R$:
 $\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
@@
 $\mathfrak{p}_{4}$ $=(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
-We claim that $m(\sigma)=5$, a quantity which will be obtained using the associativity formula. For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we start by noting that the quantities:
+For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we note that the elements
 $h$ $=3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$
@@
 $t$ $=(x_{0}-x_{1})$
-are all invertible inside the ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$.
+are all invertible in the local ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$.
 Then, we do two elementary operations:
@@
 $\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3},utx_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2},x_{3})}\simeq\kappa,$
-and hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
+and hence $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
 For $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, note that $v=x_{0}x_{2}\in R_{\mathfrak{p}_{2}}$ is a unit and denote $u=v^{-1}$. We consider some elementary operations:
@@
 is of the form a unit $x_{2}$ plus something inside the maximal ideal $(x_{3},x_{0}-x_{1})_{\mathfrak{p}_{2}}$, it is also a unit.
-Thus, we obtain $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
+Thus, we obtain $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
 An analogous consideration also holds for the case of the prime $\mathfrak{p}_{3}=(x_{2},x_{0}-x_{1})$, with similar elementary operations, now to trivialize the fourth column as $(0,1)$, and then to show that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa$.
@@
 Thus, we obtain that the length at each prime is one, but one of them has multiplicity two, therefore $m(\sigma)=5$ from the associativity formula.
-The main result of the next section is:
-$e=1$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 2$
-$e=2$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 5,$
-so for $m_{0}=12$ we conclude that these two correspond to $m(\sigma)\geq 7$ and $m(\sigma)\geq 6$, respectively. Since $m(\sigma)=5$, we conclude automatically that $e\geq 3$.
-Using Macaulay2, we obtain a minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the 33-syzygy kind, given by:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-so $e=3$ and, with $m(\sigma)=5$, we conclude $\operatorname{Bour}(\sigma)=4$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for the Bourbaki ideal $\mathcal{I}_{B}$ will be of the form:
+Using 13 with the minimal free resolution for ${\mathcal{T}}_{\sigma}$ above, after a choice of minimal degree syzygy, one gets the following resolution for $B=B_{\nu}$:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}\rightarrow\mathcal{I}_{B}\rightarrow 0,$
-presenting $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
+which presents $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
 For the rest of the section, we study some aspects of these special classes of sequences.
 Proposition 17.
-Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then:
-  • (a)
-    $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
+Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-1)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-    where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
-  • (b)
-    if $\sigma$ is nearly free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
+where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. Moreover, when $\sigma$ is nearly free, the zero-dimensional part of $\Xi_{\sigma}$ is aligned.
 Proof.
-This follows from Lemma 12, using the minimal free resolution for a line in $l\subset\mathbb{P}^{3}$ as the intersection of two planes:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\rightarrow\mathcal{I}_{l}\rightarrow 0$
-since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for a choice of syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$.
-The isolated zeros of the minors will be in the support of the sheaf $\operatorname{supp}({\mathcal{T}}_{\sigma})$, and these will be inside the line $l$ from the resolution obtained.
-∎
+Note that $\operatorname{Bour}(\sigma)=1$ if and only if $B=B_{\nu}$ is a line, for a given choice of non-zero syzygy of minimal degree $\nu$.
+Then, the equivalence follows from 13.
+In general, the singular scheme of ${\mathcal{T}}_{\sigma}$ is supported at the zero-dimensional locus of $\Xi$, and this will be contained inside $B$, a fact which may be seen from the short exact sequence
+$0\to\mathcal{O}(-e)\to{\mathcal{T}}_{\sigma}\to\mathcal{I}_{B}(e-d)\to 0.$
 Remark 18.
-The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [undefd], related to rational cuspidal curves.
-In [undefr, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [undefd] (see [undefr, Definition 2.17]).
-Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
-The notion of 33-syzygy divisors is also present in a number of previous works, for example [undef], [undefe] and [undefg].
+The notion of nearly-free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [7], related to rational cuspidal curves.
+In [17, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [7] (see [17, Definition 2.17]).
+The notion of $3$-syzygy divisors is also present in a number of previous works, for example [1], [8] and [9].
 Example 19.
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-1)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)^{\oplus 2}\xrightarrow{M}\mathcal{T}_{\sigma}\rightarrow 0,$
-and in particular $\sigma$ is a nearly free sequence.
+and in particular $\sigma$ is a nearly-free sequence.
 Proposition 20.
-If $\sigma=(f,g)$ is a nearly free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
+If $\sigma=(f,g)$ is a nearly-free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
 Proof.
@@
 $e=\frac{d+2}{2},$
-which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
+which implies that $d$ must be even and that $e=d/2+1$, giving $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
 If we denote by $E\doteq{\mathcal{T}}_{\sigma}(d/2)$ the normalized vector bundle, we must have $c_{1}(E)=0$ and
@@
 $=\frac{d^{2}}{4}-\frac{d^{2}}{2}-d+\frac{4d}{4}+1+m_{0}-m(\sigma),$
-which yields
-$m_{0}-m(\sigma)-\frac{d^{2}}{4}=0,$
-so that $c_{2}(E)=0$.
-But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ ([undefn, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis.
+which yields $m_{0}-m(\sigma)-\frac{d^{2}}{4}=0$, so that $c_{2}(E)=0$.
+But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ (by [14, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis.
 ∎
@@
 If a normal sequence $\sigma=(f,g)$ is a $3$-syzygy, then $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$.
-Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
+Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))\setminus\{0\}$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
 Proof.
 First, if we assume $\sigma$ is $3$-syzygy, then there is a free resolution of the form:
-$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0,$
-so we split the resolution into two short exact sequences:
-$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma},$
+$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0$
+with $\operatorname{{rk}}(F_{0})=2$. Split the resolution into two short exact sequences:
+$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\twoheadrightarrow{\mathcal{T}}_{\sigma},$
 and focus on the second one.
-The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [undefn, Proposition 1.1].
-Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
+The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [14, Proposition 1.1].
+Furthermore, since $\operatorname{{rk}}(F_{0})=2$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
 $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma}$
 is a free resolution for ${\mathcal{T}}_{\sigma}$, concluding $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$.
-For the equivalence stated above, if we start with a sequence $\sigma$ which is $3$-syzygy and then apply Lemma 12, $(b)$, we obtain a resolution for $\mathcal{I}_{B}$ which is of the form
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
-since $\operatorname{{rk}}(F_{0})=2$, thus concluding $B$ must be a complete intersection scheme.
-On the other hand, if $B$ is a complete intersection, then there is a resolution for $\mathcal{I}_{B}$ of the form
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\xrightarrow{r,s}\mathcal{I}_{B}\rightarrow 0,$
-induced by the two equations $r,s$ defining $B=V(r,s)\subset\mathbb{P}^{3}$.
-Thus, applying Lemma 12, $(a)$, we can lift the resolution above for a free resolution for ${\mathcal{T}}_{\sigma}$ of the form
-$0\rightarrow F_{1}\rightarrow F_{0}^{\prime}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
-so that $\operatorname{{rk}}(F_{0}^{\prime})=3$, thus $\operatorname{{rk}}(F_{0})=2$ and $\sigma$ is $3$-syzygy. ∎
+With this and 13, we conclude that $\sigma$ is $3$-syzygy if and only if $B=B_{\nu}$ admits a resolution of the form:
+$0\to\mathcal{O}_{\mathbb{P}^{3}}(-k)\to\mathcal{O}_{\mathbb{P}^{3}}(-a)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-b)\to\mathcal{I}_{B}\to 0,$
+which is equivalent to the fact that $B\subset\mathbb{P}^{3}$ is a complete intersection. ∎
 Remark 22.
@@
 $\sigma\text{ is nearly free }\Rightarrow\sigma\text{ is $3$-syzygy }\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
-where the first follows from 20 since $\sigma$ is nearly free iff $B_{\nu}$ is a line for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and every line is a complete intersection of two planes.
-The converses do not hold, as we explore in the next examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 15), pencils of cubics which satisfy $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but are not $3$-syzygy (Example 22) and pencils of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (Example 23).
-We also construct two pencils of cubics with the same discrete invariants $(m(\sigma),\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma))$ and the same Chern classes, which are distinguished by their homological behavior: one is $3$-syzygy and the other satisfies $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (see Example 22 and Example 24).
-Example 22 (pencil of cubics which is not $3$-syzygy and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$).
-Considering the following pencil of cubics:
+from 17 and 21.
+The converses do not hold, as we explore in the examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 16), and the next example is a pencil of cubics satisfying $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but not $3$-syzygy.
+Example 23 ($\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$, not $3$-syzygy).
+Consider the following pencil of cubics:
 $\sigma=(x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{2}^{3}+x_{1}x_{2}x_{3}+x_{3}^{3})$
@@
 $\nabla\sigma=\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\\ 0&x_{2}x_{3}&3x_{2}^{2}+x_{1}x_{3}&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$.
-The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines and a plane quadric curve:
-$L_{1}$ $=V(x_{0},x_{3})\doteq V(\mathfrak{p}_{1})$
-$L_{2}$ $=V(x_{2},x_{3})=V(\mathfrak{p}_{2})$
-$Q$ $=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})=V(\mathfrak{p}_{3}),$
-Now, we proceed to computing the length of the cokernel module $Q$ over each prime ideal $\mathfrak{p}_{i}$, in order to show that $m(\sigma)=5$ using the associativity formula.
-Starting with the prime ideal $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that the element $3x_{2}^{2}\in R_{\mathfrak{p}_{1}}$ is invertible, and hence the sum $v=3x_{2}^{2}+x_{1}x_{3}$ is also invertible, as $x_{1}x_{3}\in\mathfrak{p}_{1}$.
-Let $u=v^{-1}$, so we perform two elementary operations on the matrix $\nabla\sigma$ over $\mathfrak{p}_{1}$:
-$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&x_{2}x_{3}&v&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$
-$\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&ux_{2}x_{3}&1&u(x_{1}x_{2}+3x_{3}^{2})\end{pmatrix}$.
-The first transform is to add the scaled second row (by $ux_{0}^{2}$) to the first, and then we scale the second row by $u$.
-Then, the third basis vector goes via $\nabla\sigma$ to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and therefore we may compute the cokernel from the quotient
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(2x_{0}x_{2}+x_{1}x_{3},x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3},x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2}))}\simeq\frac{R_{\mathfrak{p}_{1}}}{I},$
-where $I=(f_{1},f_{2},f_{3})$ are the generators above.
-We may factor $f_{2}=x_{0}x_{3}(1-ux_{0}x_{2})$, so that since $(1-ux_{0}x_{2})$ is a unit in the local ring, we conclude $x_{0}x_{3}\in I$.
-From this, multiplying $x_{3}f_{1}$ we conclude that $x_{1}x_{3}^{2}\in I$, and since $x_{1}$ is also a unit, we obtain $x_{3}^{2}\in I$, and analogously taking $x_{0}f_{1}$ we obtain $x_{0}^{2}x_{2}\in I$, and since $x_{2}$ is a unit we get $x_{0}^{2}\in I$.
-Thus
-$(x_{0}^{2},x_{0}x_{3},x_{3}^{2})\subset I.$
-Now, from this, looking at $f_{3}$:
-$f_{3}=x_{0}x_{1}+3x_{3}^{2}-x_{0}^{2}u(x_{1}x_{2}+3x_{3}^{2})\in I\Rightarrow x_{0}x_{1}\in I\Rightarrow x_{0}\in I,$
-and looking at $f_{1}$, we obtain
-$f_{1}=2x_{0}x_{2}+x_{1}x_{3}\in I\Rightarrow x_{1}x_{3}\in I\Rightarrow x_{3}\in I,$
-so that $I=(x_{0},x_{3})$ and thus $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\kappa$, with $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
-For the case $\mathfrak{p}_{2}=(x_{2},x_{3})$, let $u=(x_{0}^{2})^{-1}$, $q=(3x_{2}^{2}+x_{1}x_{3})$ and consider the elementary operations given by $R_{2}^{\prime}=R_{2}-uqR_{1}$ and then $R_{1}^{\prime}=uR_{1}$:
-$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&u(x_{0}x_{1}+3x_{3}^{2})\\ -uq(2x_{0}x_{2}+x_{1}x_{3})&x_{2}x_{3}-uqx_{0}x_{3}&0&x_{1}x_{2}+3x_{3}^{2}-uq(3x_{3}^{2}+x_{0}x_{1})\end{pmatrix}$.
-The Jacobian matrix sends the third basis vector to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, and thus we may compute:
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I}$
-where
-$f_{1}$ $=(3x_{2}^{2}+x_{1}x_{3})(2x_{0}x_{2}+x_{1}x_{3})$
-$f_{2}$ $=x_{3}(x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3})$
-$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3ux_{3}^{2}(3x_{2}^{2}+x_{1}x_{3})-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})$
-First, we consider
-$f_{2}-\frac{x_{3}}{x_{1}}f_{3}=\frac{3x_{3}^{3}}{x_{1}}(1-u(3x_{2}^{2}+x_{1}x_{3}))\in I,$
-and thus $x_{3}^{3}\in I$. Then
-$x_{3}f_{2}=x_{3}^{2}x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3}^{3}\in I\Rightarrow x_{3}^{2}x_{2}\in I,$
-and moreover
-$x_{3}f_{1}$ $=x_{2}x_{3}(6x_{0}x_{2}^{2}+3x_{1}x_{2}x_{3}+2x_{0}x_{1}x_{3})+x_{1}^{2}x_{3}^{3}\in I$
-$\Rightarrow 6x_{0}x_{2}^{3}x_{3}+x_{3}^{2}x_{2}(3x_{1}x_{2}+2x_{0}x_{1})\in I$
-$\Rightarrow 6x_{0}x_{2}^{3}x_{3}\in I$
-$\Rightarrow x_{2}^{3}x_{3}\in I.$
-On the other hand,
-$x_{2}x_{3}f_{3}$ $=-3ux_{3}^{3}q+3x_{2}x_{3}^{3}+x_{2}x_{3}(x_{1}x_{2}-ux_{0}x_{1}q)\in I$
-$\Rightarrow x_{1}x_{2}^{2}x_{3}-3ux_{0}x_{1}x_{2}^{3}x_{3}-ux_{0}x_{1}^{2}x_{2}x_{3}^{2}\in I$
-$\Rightarrow x_{1}x_{2}^{2}x_{3}\in I\Rightarrow x_{2}^{2}x_{3}\in I.$
-Now, consider
-$x_{3}f_{3}$ $=x_{1}x_{2}x_{3}+3x_{3}^{3}-uqx_{3}^{3}-uqx_{0}x_{1}x_{3}\in I$
-$\Rightarrow x_{3}x_{2}(x_{2}-3ux_{0}x_{2}^{2}-ux_{0}x_{1}x_{3})\in I$
-$\Rightarrow x_{1}x_{3}(x_{2}-ux_{0}x_{1}x_{3})\doteq A\in I,$
-and
-$x_{3}A\in I\Rightarrow x_{1}x_{3}^{2}x_{2}-ux_{0}x_{1}^{2}x_{3}^{2}\in I\Rightarrow x_{3}^{2}\in I.$
-This simplifies
-$f_{2}=x_{3}(x_{2}-uqx_{0})\in I\Rightarrow x_{2}x_{3}(1-3ux_{2})\in I\Rightarrow x_{2}x_{3}\in I.$
-Considering
-$f_{1}$ $=6x_{0}x_{2}^{3}+x_{2}x_{3}(3x_{1}x_{2}+2x_{0}x_{1})+x_{1}^{2}x_{3}^{2}\in I$
-$\Rightarrow x_{2}^{3}\in I,$
-and then
-$x_{2}f_{3}$ $=x_{1}x_{2}^{2}+x_{2}(3x_{3}^{2}-3ux_{3}^{2}q)-uqx_{0}x_{1}x_{2}\in I$
-$\Rightarrow x_{1}x_{2}^{2}-3ux_{0}x_{1}x_{2}^{3}-ux_{0}x_{1}^{2}x_{2}x_{3}\in I$
-$\Rightarrow x_{1}x_{2}^{2}\in I\Rightarrow x_{2}^{2}\in I,$
-simplifying
-$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3uqx_{3}^{2}-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})\in I$
-$\Rightarrow x_{1}(x_{2}-ux_{0}x_{1}^{2}x_{3})\in I\Rightarrow x_{2}-ux_{0}x_{1}x_{3}\in I,$
-so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa\langle 1,x_{2}\rangle$, meaning $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=2$, since the equations $f_{1},f_{2}$ are trivialized after $x_{3}^{2},x_{2}^{2},x_{2}x_{3}\in I$.
-Let us consider the case $\mathfrak{p}_{3}=(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})\doteq(g_{1},g_{2})$.
-One may be tempted to choose $u=(2x_{0}x_{2}+x_{1}x_{3})^{-1}$ and do the elementary operations $R_{1}^{\prime}=uR_{1}$, but
-$2x_{0}x_{2}+x_{1}x_{3}=2x_{0}g_{1}+g_{2}.$
-Take $u\doteq(x_{0}^{2})^{-1}\in R_{\mathfrak{p}_{3}}$. If $x_{0}^{2}\in\mathfrak{p}_{3}$, then:
-$x_{0}\in\mathfrak{p}_{3}\Rightarrow x_{2}\in\mathfrak{p}_{3}\Rightarrow x_{1}x_{3}\in\mathfrak{p}_{3}$
-and $\text{height}(\mathfrak{p}_{3})\geq 3$ would give a contradiction. Thus, we may perform the same elementary operations as we did for $\mathfrak{p}_{2}$, and thus
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{I}$
-with generators $I=(f_{1},f_{2},f_{3})$ given as before. By the same reasoning, $(3x_{2}^{2}+x_{1}x_{3})\not\in\mathfrak{p}_{3}$, and thus
-$f_{1}\in I\Rightarrow 2x_{0}x_{2}+x_{1}x_{3}\in I.$
-From the previous description of the element above, it suffices to show $g_{i}\in I$ for some $i\in\{1,2\}$ to conclude $I=\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$.
-Let $A\doteq(g_{2}+x_{2}^{2})$, so that $u=A^{-1}$.
-We can rewrite $f_{3}$ as:
-$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-u(g_{2}+x_{2}^{2})(3x_{3}^{2}+x_{0}x_{1})$
-$=3x_{3}^{2}(1-u(g_{2}+x_{2}^{2}))+x_{1}(x_{2}-ux_{0}(g_{2}+x_{2}^{2}))$
-$=3x_{3}^{2}(1-1)+x_{1}(x_{2}-ux_{0})$
-$=-x_{1}g_{1}\in I,$
-and since $x_{1}\notin\mathfrak{p}_{4}$, $g_{1}\in I$ and $I=\mathfrak{p}_{4}R_{\mathfrak{p}_{4}}$, concluding $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=1$.
-From the associativity formula, we obtain
-$m(\sigma)=1\cdot 1+2\cdot 1+1\cdot 2=5.$
-Since $m(\sigma)\leq 6$, we conclude using the bounds from the next section that $e\geq 3$ The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
+The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines $V(x_{0},x_{3}),V(x_{2},x_{3})$ and a plane quadric curve $V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})$.
+Using the associativity formula, one may compute $m(\sigma)=5$.
+The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{M}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
@@
 $\sigma=(x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3},x_{2}x_{3}(x_{2}-x_{1})),$
-considered in [undefj, Theorem 8.1].
+considered in [11, Theorem 8.1].
 From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle.
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-So $e=3$ and $m(\sigma)=7$.
-The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$.
-Localizing the cokernel at the first two lines gives length one, so over last line should have length $5$.
-Example 24 (pencils of cubics with the same discrete invariants and different homological behavior).
-Considering the following pencil of cubics:
+So $e=3$ and $m(\sigma)=7$. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$.
+For normal sequences $\sigma=(f,g)$, the triples $(\operatorname{indeg}({\mathcal{T}}_{\sigma},\operatorname{Bour}(\sigma),m(\sigma))$ are not enough ensure the homological behavior of the sheaf ${\mathcal{T}}_{\sigma}$.
+The resolutions for the associated Bourbaki schemes also fix the arithmetic genus $p_{a}(B)$, which describes the third Chern class $c_{3}({\mathcal{T}}_{\sigma})$ of the sheaf.
+Even if these are fixed, there are curves with fixed degree and genus with different minimal free resolutions.
+The next example has the same triple $(\operatorname{indeg}({\mathcal{T}}_{\sigma},\operatorname{Bour}(\sigma),m(\sigma))$ as in the $3$-syzygy Example 16 seen before, but this one has $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$.
+Example 25.
 $\sigma=(x_{0}^{3}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{3}^{3}+x_{1}x_{3}^{2}+x_{0}x_{1}x_{3}+x_{0}^{2}x_{2}),$
@@
 $\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}+x_{3}^{2}&x_{0}^{2}&x_{0}x_{1}+2x_{1}x_{3}+3x_{3}^{2}\end{pmatrix}$.
-Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$.
-Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
+Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$. Using Macaulay2, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 6}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 6}\rightarrow\mathcal{I}_{B}\rightarrow 0.$
-This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as Example 15, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
-In Section 4 we characterize all nearly free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly free sequences with $d_{f}=1,d_{g}=2$.
+In Section 4 we characterize all nearly-free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly-free sequences with $d_{f}=1,d_{g}=2$.
 3 Extreme cases of low initial degree
-In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$.
-We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two.
-From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 27.
-We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example, [undefc]).
-As explored in [undefj, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
-$0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
-where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the associated foliation.
-Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
+As explored in [11, Section 9], a sequence $\sigma=(f,g)$ defines a codimension-one foliation with a corresponding short exact sequence:
+$\mathcal{F}_{\sigma}:0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
+where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is called the singular scheme of $\mathcal{F}_{\sigma}$.
+Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a nonzero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
 […]
@@
 defining a subscheme $W\subset\mathbb{P}^{3}$, called the singular scheme of the associated foliation by curves.
-It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}$.
-The classification of such foliations by curves of low degree provides the following bounds on their singular schemes:
+It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}(e-1)$.
+We derive numerical restrictions for $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two.
+A similar strategy was employed to study distributions on $\mathbb{P}^{3}$ with unstable tangent sheaf in [2], and general considerations about the minimal degree of sub-foliations have been considered in [19].
+One may localize the diagram above at a minimal prime of the scheme $B$, arguing as in [6, Proposition 3.4].
+Over this prime, both sheaves ${\mathcal{T}}_{\sigma}$ and $\mathbb{T}\mathbb{P}^{3}$ split as sums of line bundles, and we obtain a diagram
+$\mathcal{O}_{\mathbb{P}^{3},p}(1-e)$$\mathcal{O}_{\mathbb{P}^{3},p}(1-e)$$\mathcal{O}_{\mathbb{P}^{3},p}(-a)\oplus\mathcal{O}_{\mathbb{P}^{3},p}(-b)$$\bigoplus_{i=1}^{3}\mathcal{O}_{\mathbb{P}^{3},p}(l_{i})$$\mathcal{I}_{B}(e-d+1)_{p}$$G_{p}$$\scriptstyle{\nu}$$\scriptstyle{\nu^{\prime}}$$\scriptstyle{L}$
+describing $W$ and $B$ via matrix presentations, as the vanishing set of $\nu^{\prime}$ and $\nu$, respectively.
+From the composition of matrices $L\cdot\nu=\nu^{\prime}$ we conclude minors of $\nu^{\prime}$ are factors of the minors of $\nu$, and thus $B\subset W$ at every associated prime, hence $B\subset W$ schematically.
+Now, we turn to the classification of schemes $W$ for foliations by curves on $\mathbb{P}^{3}$ of degrees $e=1,2$:
+Lemma 26.
+Let
+$\mathcal{F}:0\to\mathcal{O}_{\mathbb{P}^{3}}(1-e)\to\mathbb{T}\mathbb{P}^{3}\to G\to 0$
+be a foliation by curves of degree $e\geq 0$ in $\mathbb{P}^{3}$, with singular scheme $W\subset\mathbb{P}^{3}$. Then:
   • (a)
-    ([undefk, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1.
-In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
+    ([12, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or a pair of skew lines, in which case $G^{\vee}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}$.
+In particular, $\deg(W)\leq 2$.
   • (b)
-    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$ or is zero-dimensional.
+    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$.
+Proof.
 For completeness, we include an argument for $(b)$, from V. Cordeiro: using the short exact sequence
 $0\to G^{\vee}\to\Omega^{1}_{\mathbb{P}^{3}}\to\mathcal{I}_{W}(1)\to 0,$
-one obtains $4\leq c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [undefc, 4.1, 4.2]).
-Now, if $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$.
-Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a section to form a sequence
+one obtains $c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [5, 4.1, 4.2]).If $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$, hence $\deg(W)\leq 4$.
+Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a nonzero section to form a sequence
 $0\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\to G^{\vee}\to\mathcal{I}_{Y}(-3)\to 0$
-where $Y$ is the vanishing locus of this section.
-Hence, from additivity one gets $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$.
-From the first inequality, we conclude $c_{2}(G^{\vee})\geq 6$, and thus $\deg(W)\leq 5$.
+where $Y$ is the vanishing locus of this section. From the sequence, we get $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$, so $c_{2}(G^{\vee})\geq 6$, and $\deg(W)\leq 5$. ∎
 Theorem B.
@@
 Proof.
-To show $(a)$, we dualize the following short exact sequence, obtained above for $e=1$:
-$0\rightarrow\mathcal{I}_{B}(2-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
-to get a long exact sequence, after simplifying, of the form:
-$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-2)\rightarrow$
-$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+2)\rightarrow$
-$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
-where $\mathcal{U}$ is defined by the short exact sequence $0\to\mathcal{U}\to\mathcal{O}_{\Gamma_{\sigma}}\to\mathcal{O}_{Y}\to 0$ and $Y$ is the one-dimensional component of $\Gamma\sigma$.
-Moreover, since $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$, we may consider the final piece of the long exact sequence
-$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+2)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
-where $W^{\prime}\subset W$ is a pure one-dimensional subscheme.
-Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$, and from the classification $\deg(W^{\prime})\leq\deg(W)\leq 2$, hence the result follows.
-For $(b)$, we proceed analogously, dualizing the sequence
-$0\rightarrow\mathcal{I}_{B}(3-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0$
-to obtain a long exact sequence, after simplifying, of the form:
-$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-3)\rightarrow$
-$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+1)\rightarrow$
-$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
-where $\mathcal{U}\hookrightarrow\mathcal{O}_{\Gamma_{\sigma}}\rightarrow\mathcal{O}_{Y}$ and $Y$ is the one-dimensional component of $\Gamma_{\sigma}$.
-Moreover, $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$ and from the final piece of the sequence
-$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+1)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
-where $W^{\prime}\subset W$ is a pure one-dimensional subscheme.
-Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{O}_{\Gamma_{\sigma}},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$.
-Since $W^{\prime}$ is a subscheme of $W$, we have $\deg(W^{\prime})\leq\deg(W)$.
-The classification of foliations by curves of degree $2$ implies $\deg(W)\leq 5$, which completes the proof.
-∎
+The result follows from the schematic inclusion $B\subset W$ and from the classification results for the singular scheme $W$, since $\deg(B)\leq\deg(W)$. ∎
 To be able to present the next applications, we will need the following result on the structure of degree $2$ space curves in $\mathbb{P}^{3}$:
 Theorem 27.
-[undeft, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree 22 and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
+[20, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree $2$ and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
   • (a)
@@
 If $a=0$ and $B$ is a union of two skew lines, then $\omega_{B}\simeq\mathcal{O}_{B}(-2)$.
+For the case $e=1$ and $\operatorname{Bour}(\sigma)=2$, one gets $B=W$ and we can predict the Chern classes for the logarithmic sheaf, by the following proposition.
+We note that we have not been able to construct examples of sequences of this kind.
+Proposition 28.
+Let $\sigma=(f,g)$ be a normal sequence with $e=1$ and $\operatorname{Bour}(\sigma)=2$. Then, the Chern classes of ${\mathcal{T}}_{\sigma}$ are (d,d+1,2d)(-d,d+1,2d).
 Proof.
-Items $(a)$ and $(b)$ are in the original paper, and to obtain $(c)$, for $a\geq 1$ we can apply the functor $\operatorname{\mathcal{H}\textit{om}}(-,\omega_{\mathbb{P}^{3}})$ to the short exact sequence in $(ii)$, use that $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ and twist by $a+2$ to obtain a short exact sequence
-$0\rightarrow\mathcal{O}_{L}(a)\rightarrow\omega_{B}(a+2)\rightarrow\mathcal{O}_{L}\rightarrow 0$
-which coincides with the short sequence in $(b)$, hence we obtain the isomorphism in the claim.
-For the case $a=0$, $B$ must be either a multiplicity two structure as in $(b)$ with $a=0$ or a union of skew lines.
-In the first case, the proof follows as before.
-If $B=L_{1}\cup L_{2}$, then $\omega_{B}\simeq\omega_{L_{1}}\oplus\omega_{L_{2}}$ and the claim follows from the fact $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ in each component.
-∎
-Remark 26.
-The results above can be used to restrict the possible values of numerical invariants.
-For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results to show the only possible third Chern class is $c_{3}({\mathcal{T}}_{\sigma})=8$.
-Indeed, from [undefu, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 6 we obtain $c_{3}({\mathcal{T}}_{\sigma})=2(p_{a}(B)+5)=8-2a$, where $p_{a}(B)=-1-a$.
-From the short exact sequence
-$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(4-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
-we obtain the Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(4-a))=10-a$ and
-$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=10-2a,$
-and additivity of Euler characteristic gives a contradiction whenever $a\neq 0$, so $c_{3}({\mathcal{T}}_{\sigma})=8$ if this occurs.
-In the next section, we show the above cannot happen either (see 43).
-As another application, we are able to show that sequences $\sigma$ with mixed degrees $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$ cannot be unstable.
+From the formulas $c_{2}({\mathcal{T}}_{\sigma})=m_{0}-m(\sigma)$ and $\operatorname{Bour}(\sigma)=2$, we get $c_{2}({\mathcal{T}}_{\sigma})=d+1$.
+Since $B=W$ is a pair of skew lines in this case, $p_{a}(B)=-1$ and from
+$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
+it follows that $c_{3}({\mathcal{T}}_{\sigma})=2d$. ∎
+The last proposition of the section also uses the foliation structure, and it follows from [6, Proposition B].
 Proposition 29.
-Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$ and $m(\sigma)=3$.
-If ${\mathcal{T}}_{\sigma}$ is unstable, then the possible discrete invariants are $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=6$.
-Proof.
-Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$.
-From Theorem A Theorem A, $\sigma$ is incompressible.
-Using Remark 6 we have for $\deg(B)=2$, $p_{a}(B)=-1-a$ yields $c_{3}({\mathcal{T}}_{\sigma})=6-2a\geq 0$, so $a\in\{-1,0,1,2,3\}$, but from
-From the short exact sequence
-$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(3-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
-we get Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(3-a))=8-a$ and
-$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=8-2a,$
-which can only coincide for $a=0$. ∎
-We note that we have no explicit construction for the unstable case above to occur as an example.
+Let $\sigma=(f,g)$ be a normal sequence with $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. If $\sigma$ is not free, then ${\mathcal{T}}_{\sigma}$ is not locally free.
 4 Pencils of cubics and degree 6 curves inside quadric surfaces
 In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces.
-The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [undefn], [undefw], [undefo], [undefb] and [undefp].
-Two major results which are used are the following ones, both assuming that $\kappa$ is an algebraically closed field of characteristic zero.
-Theorem 28 ([undefp], Theorem 1.1).
+The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [14], [22], [15], [4] and [16].
+We recall two major results:
+Theorem 30 ([16], Theorem 1.1).
 Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$. Assume $c_{1}\geq-3$ and $h^{0}(\mathcal{E})=0$. Define the integers:
@@
 Another important result is the following, which enables us to obtain a lower bound for the existence of a syzygy in some cases:
-Theorem 29 ([undefo], Theorem 0.1).
+Theorem 31 ([15], Theorem 0.1).
 Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$ with $c_{1}=0$ of $c_{1}=-1$ and with $c_{2}\geq 0$. Let $t\in\mathbb{Z}$ such that either
@@
 $\operatorname{Bour}(\sigma)=e(e-4)+12-m(\sigma).$
-When $e=1$, the formula above implies $m(\sigma)\leq 9$, from the inequality $\operatorname{Bour}(\sigma)\geq 0$.
-Applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get
-$c_{1}(\mathcal{E})=c_{1}({\mathcal{T}}_{\sigma}(1))=-d+2=-2$
-and the quantities of the Theorem become
-$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(4-4+3)\right\rceil=1$
-$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil=\left\lceil\frac{1}{3}(4-12+8)\right\rceil=0.$
-Moreover, another important invariant is:
-$c_{2}(\mathcal{E})$ $=c_{2}({\mathcal{T}}_{\sigma})-d+1$
-$=m_{0}-m(\sigma)-d+1$
-$=9-m(\sigma).$
-Direct application of the result shows the next pair of propositions:
+When $e=1$, $\operatorname{Bour}(\sigma)\geq 0$ gives $m(\sigma)\leq 9$.
+Applying Theorem 30 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get $c_{1}(\mathcal{E})=-2$ and the quantities of the Theorem become $A=1$ and $B=2$.
+Since $c_{2}(\mathcal{E})=9-m(\sigma)$ direct application of the result shows the next proposition.
 Proposition 32.
-If $\sigma=(f,g)$ is a normal pencil of cubics with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 9$.
+If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 8$. Furthermore:
+  • (a)
+    if $7\leq m(\sigma)\leq 8$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
+  • (b)
+    if $0\leq m(\sigma)<7$, then $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72$.
+Next, we apply Theorem 31 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=8-m(\sigma)\geq 0$, assuming $e>1$.
+Then, for $t>\sqrt{3c_{2}+1}-2=\sqrt{24-3m(\sigma)+1}-2$ we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$.
+Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives the following result.
 Proposition 33.
-If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then
-  • (a)
-    if $8\leq m(\sigma)\leq 9$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
-  • (b)
-    if $0\leq m(\sigma)<7$, then
-    $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72.$
-Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=9-m(\sigma)\geq 0$ from the previous computations, assuming $e>1$.
-Then, the Theorem says that, for
-$t$ $>\sqrt{3c_{2}+1}-2=\sqrt{27-3m(\sigma)+1}-2$
-we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$.
-Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives minimum $t=2,1,1$ satisfying the inequality above, so we obtain
-Proposition 32.
-Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$, and we have the following table of sharper bounds for each possibility of $m=m(\sigma)$:
-$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
-7 $e\leq 3$
-8 $e\leq 3$
+Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$. If $m(\sigma)=6,7$, $e\leq 3$, and if $m(\sigma)=8$, $e\leq 2$.
 The following result gives a picture of the generic case of a pencil of cubics.
@@
 Proof.
-This follows from intersection theory for the bundle of principal parts (see [undefi], Proposition $7.1$ and Proposition $7.4$). ∎
+This follows from intersection theory for the bundle of principal parts (see [10], Proposition $7.1$ and Proposition $7.4$). ∎
 Rewriting Theorem A Theorem A, $(d)$ for pencils of cubics, we obtain:
@@
 Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$.
-Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 34, 37 and 39.
+Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 35, 38 and 40.
 Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
@@
 $=8-m(\sigma)>5$
-if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Theorem B Theorem B, $(b)$. ∎
+if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$. ∎
 Proposition 38.
@@
 Proof.
-The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 30 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$.
+The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 32 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$.
 From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
@@
 Proof.
-If $m(\sigma)=8$, by 32, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$.
-If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [undefb, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$.
-If we assume $e\in{1,3}$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 6 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain:
-$e=1$ $\Rightarrow c_{3}=4$
-$e=3$ $\Rightarrow c_{3}=0$
-On the other hand, 31 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction with the case $e=1$.
-For the case $e=3$, we would have a nearly-free locally free pencil, which cannot happen by 19.
-Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$.
+If $m(\sigma)=8$, by $3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 2$, and by 32 it follows that ${\mathcal{T}}_{\sigma}$ is locally free.
+If $e=1$, then $\operatorname{Bour}(\sigma)=1$, but by 20 nearly-free sequences are never locally free, so this gives a contradiction.
+Since $\sigma$ is incompressible, it follows that $e=2$, and for $\operatorname{Bour}(\sigma)=0$ we obtain ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}$.
+Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ gives $m(\sigma)=8$.
 ∎
@@
 Proposition 42.
-Let $\sigma=(f,g)$ be a nearly free pencil of cubics.
+Let $\sigma=(f,g)$ be a nearly-free pencil of cubics.
 Then, the only possible discrete invariants are $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $m(\sigma)=7$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 43).
 Proof.
-Using Remark 6 for $p_{a}(B)=0$ and $\deg(B)=1$, we obtain
-$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2e)$
-$=-2+8-2e=2(3-e).$
-On the other hand, if $m(\sigma)=7$, then $e=2$ and the formula for $c_{3}$ in Remark 6 yields $c_{3}({\mathcal{T}}_{\sigma})=2$.
-As we have observed in Example 40 and before, we must have $m(\sigma)\leq 7$ to be able to obtain $\operatorname{Bour}(\sigma)=1$.
-On the other hand, assuming $m(\sigma)\leq 6$, we obtain
-$\operatorname{Bour}(\sigma)=12-m(\sigma)+e(e-4)\geq 6+e(e-4)\geq 2$
-hence $\operatorname{Bour}(\sigma)\neq 1$. ∎
-Example 42 (Nearly free pencil of Cubics).
+By the previous results, $m(\sigma)\leq 7$.
+But, if $m(\sigma)\leq 6$, we get $\operatorname{Bour}(\sigma)\geq 2$, so $\sigma$ is not nearly free.
+For $m(\sigma)=7$, any value of $e\neq 2$ gives $\operatorname{Bour}(\sigma)\neq 1$.
+Thus, it follows that $e=2$.
+Moreover, $p_{a}(B)=0$ and $\deg(B)=1$ imply $c_{2}({\mathcal{T}}_{\sigma})=3$.
+∎
+Example 43 (Nearly-free pencil of Cubics).
 We consider the following sequence of cubics:
@@
 $\mathfrak{p}_{4}$ $=(x_{0},x_{2})$
-For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and the elementary operation $R_{1}^{\prime}=uR_{1}$, to obtain the matrix
+For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and an elementary operation to obtain the matrix
 $(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}1&u(x_{0}^{2}+x_{2}^{2})&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$,
@@
 $\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3},x_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}}\simeq\kappa,$
-so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
+so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
 For the prime $\mathfrak{p}_{2}=(x_{1}-x_{2},x_{3})$, the same element $u\in R_{\mathfrak{p}_{2}}$ is invertible, so that we compute the cokernel analogously to obtain:
@@
 $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{1}-x_{2})}\simeq\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
-so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
-For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}+u[x_{2}x_{3}]R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to obtain
+so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform elementary operations and obtain
 $(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}u(2x_{0}(x_{1}-x_{2})-x_{2}^{2})&1&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&0&x_{3}(2x_{2}-x_{1})-ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&x_{2}(x_{2}-x_{1})-ux_{2}^{3}x_{3}\end{pmatrix}$,
@@
 $f_{2}$ $=x_{2}(x_{2}-x_{1})+ux_{2}^{3}x_{3}.$
-Considering $f_{1}\in I$, we obtain
+Since $f_{1}\in I$, we obtain
 $2x_{2}(1+ux_{2}(x_{1}-x_{0}+x_{3})-ux_{0}^{2})\equiv x_{1}\mod I.$
@@
 $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa\langle 1,x_{1},x_{2}\rangle,$
-so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
-For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and the elementary operations $R_{1}^{\prime}=R_{1}-u[-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})]R_{2}\doteq R_{1}-uQR_{2}$ and $R_{2}^{\prime}=uR_{2}$, to obtain:
+so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
+For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and perform elementary matrix operations to obtain:
 $(\nabla\sigma)_{\mathfrak{p}_{4}}\sim\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}&0&x_{2}^{2}-uQx_{2}(x_{2}-x_{1})\\ 0&-ux_{2}x_{3}&1&ux_{2}(x_{2}-x_{1})\end{pmatrix}$,
@@
 $f_{3}$ $=x_{2}^{2}+u(x_{2}-x_{1})x_{0}^{2}x_{2}-2u(x_{2}-x_{1})(x_{1}-x_{0}+x_{3})x_{2}^{2},$
-where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have
-$x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I,$
-and thus
+where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have $x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I$, and thus
 $f_{2}\equiv x_{0}(x_{0}(1-ux_{3}x_{2})+2(x_{1}-x_{2})+2u(x_{1}-x_{2})(x_{1}-x_{0}+x_{3})x_{3})\mod I,$
@@
 $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},x_{2}^{2})}\simeq\kappa\langle 1,x_{2}\rangle,$
-with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
+with $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
 $m(\sigma)=1\cdot 1+1\cdot 1+3\cdot 1+2\cdot 1=7.$
-On the other hand, $e(\sigma)=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
+On the other hand, $e=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
 $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0.$
@@
 Let $\sigma=(f,g)$ be a normal, $\mu$-semistable pencil of cubics such that $m(\sigma)=7$. Then, the only possible cases are:
-  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 41);
+  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 42);
   • $e=3$ and $\sigma$ is locally free, $\operatorname{Bour}(\sigma)=2$ and $B$ is a pair of skew lines (or their degeneration) (see Example 24).
 Proof.
-From 32, we know that $e\leq 3$.
-For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [undefb, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case).
+From 33, we know that $e\leq 3$.
+For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [4, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case).
 For the strictly semistable case $e=2$ we obtain $\operatorname{Bour}(\sigma)=1$, so we get the first case.
-For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [undefx] fitting in a sequence of the form
+For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [23] fitting in a sequence of the form
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}(2)\rightarrow\mathcal{I}_{B}(1)\rightarrow 0$
 where $B$ is a pair of skew lines or their degeneration, thus $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=3$. ∎
+Summarizing the main results for normal pencils of cubics, we get the following Theorem:
 Theorem C.
@@
   • (c)
-    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42), which is strictly $\mu$-semistable;
+    There is only one case of nearly-free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 43), which is strictly $\mu$-semistable;
   • (d)
@@
 Proof.
-Item $(a)$ is 34 and item $(b)$ with 37 and Example 40. Item $(c)$ follows from 43 and item $(d)$ is 36. ∎
-To finish this study, we consider a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
+Item $(a)$ is 35 and item $(b)$ with 38 and Example 41. Item $(c)$ follows from 44 and item $(d)$ is 37. ∎
+Next, we consider some examples. First, a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
 Example 45 ($m(\sigma)=4$, $e=2$, $\operatorname{Bour}(\sigma)=4$, $3$-syzygy).
-Considering the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
+Consider the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
 $\sigma=(x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3},x_{0}^{3}+x_{1}^{3}+x_{2}x_{3}^{2}),$
@@
 The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$.
-Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$, first $R_{1}^{\prime}=R_{1}-R_{2}$, and then $R_{2}^{\prime}=u/3\cdot R_{2}$, where $u=(x_{1}^{2})^{-1}\in R_{\mathfrak{p}}$, to obtain the form:
+Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$ to obtain the form:
 $(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}0&0&3x_{2}^{2}-x_{3}^{2}&3x_{3}^{2}-2x_{2}x_{3}\\ ux_{0}^{2}&1&(u/3)x_{3}^{2}&(2/3)ux_{2}x_{3}\end{pmatrix}$,
 so it sends the second basis vector to $(0,1)\in R_{\mathfrak{p}}^{2}$, and then we may compute the cokernel using the first line and remaining columns:
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{P}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
+$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{\mathfrak{p}}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
 $\simeq\kappa\langle 1,x_{2},x_{3},x_{2}^{2}\rangle,$
-and thus we conclude $\operatorname{{length}}(\mathcal{Q})_{\mathfrak{p}}=4$.
+and thus we conclude $\operatorname{\operatorname{length}}(\mathcal{Q})_{\mathfrak{p}}=4$.
 Since $\deg(R/\mathfrak{p})=1$, using the associativity formula we obtain $m(\sigma)=4$.
@@
 obtained using Macaulay2, so this is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
-To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not only depend on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
+To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not depend only on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
 Example 46.
@@
 $\operatorname{Bour}(\sigma)=e(e-3)+7-m(\sigma).$
-We start by applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible. The quantities of the theorem become
-$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(9-6+3)\right\rceil=0$
-$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil-1=\left\lceil\frac{1}{3}(9-9+8)\right\rceil-1=1.$
-Moreover, another important invariant is: $c_{2}(\mathcal{E})=7-m(\sigma)$. Direct application of the result shows the next pair of propositions:
+We start by applying Theorem 30 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible.
+The quantities become $A=B=2$.
+Moreover, $c_{2}(\mathcal{E})=7-m(\sigma)$, so direct application of the result shows the following:
 Proposition 47.
-If $\sigma=(f,g)$ is an incompressible normal sequence with $d_{f}=1,d_{g}=2$, then $m(\sigma)\leq 5$.
-Proof.
-Let us assume that $e\in[1,d-2]$ and $d\geq 3$. Then, from the Bourbaki formula and the fact that $\operatorname{Bour}(\sigma)\geq 0$, we obtain
-$m(\sigma)\leq m_{0}+e(e-d)\leq m_{0}+e-d\leq m_{0}-2,$
-which gives the claim. This covers the case $e=1$. If we now assume $e>1$ and apply Theorem 28 to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a reflexive sheaf with $c_{1}=-1$, we obtain the constant is
-$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(1-2+3)\right\rceil=0.$
-Hence, $5-m(\sigma)=c_{2}({\mathcal{T}}_{\sigma}(1))=c_{2}(\mathcal{E})\geq A=0$ and thus $m(\sigma)\leq 5$. ∎
-Proposition 47.
-Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then, the following hold:
+Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then $m(\sigma)\leq 5$ and the following hold:
   • (a)
-    If $6\leq m(\sigma)\leq 7$, then $c_{3}\leq 5-m(\sigma)$.
+    If $m(\sigma)=5$, then $c_{3}({\mathcal{T}}_{\sigma})=0$.
   • (b)
-    If $0\leq m(\sigma)<6$, then $c_{3}\leq m(\sigma)^{2}-14m(\sigma)+47$.
+    If $0\leq m(\sigma)<4$, then $c_{3}\leq m(\sigma)^{2}-12m(\sigma)+35$.
 From Theorem A Theorem A, $(d)$, we obtain:
@@
 Example 49 (Free and compressible sequence with $d_{f}=1,d_{g}=2$).
-Considering the sequence
+Consider the sequence
 $\sigma=(x_{0}(x_{1}-x_{2}),x_{0}^{3}+x_{1}^{3}+x_{2}^{3}),$
@@
 gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
-Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, the theorem says that, for
+Next, we apply Theorem 31 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, for
 $t$ $>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}=\sqrt{15-3m(\sigma)+\frac{1}{4}}-\frac{3}{2}$
-we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$.
-Direct application of this bound for high values of $m(\sigma)=4,5$ we get $t=0,t=1$, respectively. This shows the following proposition.
+we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$. Direct application of this bound shows the following proposition.
 Proposition 50.
 Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1$, $d_{g}=2$.
-Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$, and we have the following table of bounds for the following possibilities of $m(\sigma)$:
-$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
-4 $e\leq 2$
-5 $e\leq 1$
-As an easy consequence of Remark 6, we obtain:
+Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$.
+Whenever $m(\sigma)=4$, $e\leq 2$ and, whenever $m(\sigma)=5$, $e\leq 1$.
+We explore some parity results in the next two propositions about local freeness.
 Proposition 51.
-Let $\sigma=(f,g)$ be a normal sequence which is not a pencil, (i.e., with $d_{f}\neq d_{g}$) such that $d$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
+Let $\sigma=(f,g)$ be a normal sequence such that $d=d_{f}+d_{g}$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
 Proof.
-Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore
-$2p_{a}(B)$ $=2-\operatorname{Bour}(\sigma)(4+d-2e)$
-$p_{a}(B)$ $=1-\operatorname{Bour}(\sigma)(2+\frac{d}{2}-e)=1-(2-e)\operatorname{Bour}(\sigma)+\frac{d\operatorname{Bour}(\sigma)}{2},$
-which is not an integer if $\operatorname{Bour}(\sigma)$ is odd, so we get a contradiction. ∎
+Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore $2p_{a}(B)-2=-\operatorname{Bour}(\sigma)(4+d-2e)$.
+But the left-hand-side is even, and the right-hand side is a product of two odd numbers, hence odd, so we get a contradiction.
+∎
 Proposition 52.
@@
 Proof.
-To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 51.
-From 3, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
-Assuming $e=2k$ is even, we obtain
+We show that, in any of these cases above, the Bourbaki degree $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$ is odd, thus the result follows from 51.
+Whenever $m(\sigma)$ is even, $7-m(\sigma)$ is odd.
+We claim $e^{2}-3e$ is always an even number for $e\geq 0$, and therefore $\operatorname{Bour}(\sigma)$ is always odd for $m(\sigma)$ even.
+First, assuming $e=2k$ is even, we obtain
 $e^{2}-3e=4k^{2}-6k=2(2k^{2}-3k),$
@@
 Proof.
-Assuming $\sigma$ is non-compressible, by 50, we obtain that $e=1$.
+Assuming $\sigma$ is incompressible, by 50, we obtain that $e=1$.
 But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free.
@@
 Example 54 (Free and incompressible sequence, $m(\sigma)=5$).
-Considering the sequence
+Consider the sequence
 $\sigma=(x_{0}x_{1},x_{3}x_{2}(x_{0}-x_{1})),$
@@
     ${\mathcal{T}}_{\sigma}$ is unstable with $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 15);
-Furthermore, these are the only two possibilities of numerical invariants for nearly free sequences with $d_{f}=1,d_{g}=2$.
+Furthermore, these are the only two possibilities of numerical invariants for nearly-free sequences with $d_{f}=1,d_{g}=2$.
 Proof.
-Using 50, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$.
-Using Remark 6 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples.
-To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
+Using 50, $e\in\{1,2\}$, and both cases imply $\operatorname{Bour}(\sigma)=1$ when $m(\sigma)=4$.
+Using $p_{a}(B)=0$ and $\deg(B)=1$, we obtain the $c_{3}$’s above, and both can be realized.
+To conclude the last claim, for $m(\sigma)\leq 3$,
 $\operatorname{Bour}(\sigma)=7-m(\sigma)+e(e-3)\geq 4+e(e-3)\geq 2$
-for $e\in\{1,2,3\}$. ∎
-Example 56 (Nearly free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
+for possible values $1\leq e\leq 3$. ∎
+Example 56 (Nearly-free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
 We consider the following normal sequence with $d_{f}=1$, $d_{g}=2$:
@@
 $\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&-x_{3}&-x_{2}\\ x_{1}x_{3}&x_{3}(x_{0}-x_{2})&-x_{1}x_{3}&x_{1}(x_{0}-x_{2})\end{pmatrix}$.
-The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines, so $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$.
-We may compute $m(\sigma)=4$, where the first line above occurs with length two and the other two occur with length one.
-For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform two elementary operations, namely $R_{2}^{\prime}=R_{2}-ux_{3}(x_{0}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite the matrix:
+The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines.
+We compute $m(\sigma)=4$, where the first line above occurs with multiplicity two and the other two are simple.
+For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform elementary operations to rewrite the matrix as
 $(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{1}&1&-ux_{3}&-ux_{2}\\ x_{1}x_{3}(1-u(x_{0}-x_{2}))&0&-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u)&(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u)\end{pmatrix}$.
@@
 $\simeq\kappa\langle 1,x_{3}\rangle,$
-simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
-For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform two elementary operations $R_{2}^{\prime}=R_{2}+ux_{1}x_{3}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite $\nabla\sigma$ as
+simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
+For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform elementary operations to rewrite $\nabla\sigma$ as
 $(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}ux_{1}&ux_{0}&1&-ux_{2}\\ x_{1}x_{3}(1+ux_{1})&x_{3}((x_{0}-x_{2})+ux_{0}x_{1})&0&x_{1}(x_{0}-x_{2})-ux_{1}x_{2}x_{3}\end{pmatrix}$.
@@
 $x_{3}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
-and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
-For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and the elementary operations $R_{2}^{\prime}=R_{2}-ux_{1}x_{3}R_{1}$, $R_{1}^{\prime}=uR_{1}$ to obtain
+and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
+For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and an elementary operations to obtain
 $(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}1&ux_{0}&-ux_{3}&-ux_{2}\\ 0&x_{3}((x_{0}-x_{2})-ux_{0}x_{1})&x_{1}x_{3}(-1+ux_{3})&x_{1}(x_{0}-x_{2})+ux_{1}x_{2}x_{3}\end{pmatrix}$.
@@
 $x_{1}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
-and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
-From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain
-$m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4.$
-From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly free sequence.
-The complete minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
+and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
+From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain $m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4$.
+From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly-free sequence.
+The minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\xrightarrow{M}{\mathcal{T}}_{\sigma}\rightarrow 0$
@@
 $M=\begin{pmatrix}x_{0}x_{1}+x_{3}(x_{2}-x_{0})&x_{0}^{2}&x_{0}x_{3}\\ -x_{1}^{2}+x_{1}x_{3}&-x_{0}x_{1}&0\\ x_{1}x_{2}&x_{2}^{2}&x_{0}x_{1}+x_{3}(x_{2}-x_{1})\\ 0&-x_{2}x_{3}&x_{1}x_{3}-x_{3}^{2}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}\\ x_{3}-x_{1}\\ -x_{2}\end{pmatrix}$.
-which corresponds to the nearly free resolution given in 16.
+which corresponds to the nearly-free resolution given in 17.
 We have $c_{3}({\mathcal{T}}_{\sigma})=1$, corresponding to the point $p$, the unique irreducible component of $\Xi_{\sigma}$ of codimension three.
@@
 Proof.
-We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$.
-Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$.
-This follows from [undefw, Proposition 1.1] for $c_{3}=0$, [undefb, Lemma 2.4] for $c_{3}=2$ and [undefn, Lemma 9.6] for $c_{3}=4$.
+The sheaf ${\mathcal{T}}_{\sigma}(1)$ is a stable rank two reflexive sheaf with Chern classes $(-1,2,c_{3})$, and $c_{3}\leq c_{2}^{2}=4$, using [14, Theorem 8.2].
+All three possibilities $c_{3}\in\{0,2,4\}$ imply $e=2$.
+Indeed, this follows from [22, Proposition 1.1] for $c_{3}=0$, [4, Lemma 2.4] for $c_{3}=2$ and [14, Lemma 9.6] for $c_{3}=4$.
 ∎
-The unstable case of the above Proposition is numerically described in 27.
+Next examples illustrate some cases of the proposition above.
 Example 58 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=2$).
@@
 Here, $(\Xi_{\sigma})_{1}$ has two irreducible components, the lines $V(x_{3},x_{0})$ and $V(x_{3},x_{0}-x_{1})$. Let the corresponding primes be denoted by $\mathfrak{p}_{1},\mathfrak{p}_{2}$.
-Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$.
-Consider the elementary operations $R_{2}=R_{2}-u(x_{0}x_{1}+3x_{3}^{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite
+Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$. Consider elementary operations to rewrite
 $(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{3}&-ux_{3}&0&1\\ 2x_{0}x_{2}+x_{1}x_{3}-ux_{3}&x_{0}x_{3}+uQx_{3}&x_{0}^{2}&0\end{pmatrix}$,
@@
 $f_{3}$ $=x_{0}^{2}.$
-Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$. Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain
-$uf_{1}=2x_{0}^{3}x_{1}x_{2},$
-and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
+Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$.
+Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain $uf_{1}=2x_{0}^{3}x_{1}x_{2}$, and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
 $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})}\simeq\kappa\langle 1,x_{0}\rangle$
-so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
-Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider the elementary operation and $R_{2}^{\prime}=uR_{2}$, to obtain the matrix
+so that $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
+Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider an elementary operation to obtain a matrix
 $(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$,
@@
 $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}=\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
-and therefore $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
-Thus, from the additivity formula, we obtain $m(\sigma)=3$. Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$.
+and therefore $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+Thus, from the additivity formula, we obtain $m(\sigma)=3$.
+Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$.
 The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
 $\mathfrak{p}_{3}$ $=(x_{2}-x_{3},x_{0}-x_{1}).$
-Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}-ux_{3}(x_{3}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite:
+Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations and rewrite:
 $(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}2ux_{0}&1&2ux_{2}&2ux_{3}\\ x_{3}(x_{2}-x_{3})(1+ux_{0})&0&x_{3}((x_{0}-x_{1})-2ux_{2}(x_{3}-x_{2}))&(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2})\end{pmatrix}$
@@
 $f_{3}\equiv x_{2}(x_{0}-x_{1})\Rightarrow x_{2}\in I,$
-since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
+since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
 Over $\mathfrak{p}_{2}=((x_{0}-x_{1}),x_{3})$, we may choose the same elementary operations and thus
@@
 Since $(x_{2}-2x_{3})$ is invertible in $R_{\mathfrak{p}_{2}}$, we obtain $(x_{0}-x_{1})\in I$.
-From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
+From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
 Over the prime $\mathfrak{p}_{3}$, we may also do the same elementary operations and consider the same generators for the ideal $I$.
@@
 $f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\Rightarrow x_{3}(x_{0}-x_{1})\in I\Rightarrow(x_{0}-x_{1})\in I,$
-so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$.
+so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$.
 Thus, from the associativity formula $m(\sigma)=3$.
@@
 $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
-so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane smooth quadric.
+so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane conic.
 Corollary 60.
@@
 ∎
-Proposition 61.
-Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$.
-If $m(\sigma)=2$, then $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, $c_{3}\in\{1,3,5,7\}$ and $c_{3}=7$ if and only if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, with $\operatorname{Bour}(\sigma)=3$ and $p_{a}(B)=0$ (see Example 62).
-Proof.
-If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [undefn, Section 9].
-By the proof of [undefn, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
-From Remark 6, we obtain $p_{a}(B)=1$, and thus $B$ must be a plane cubic curve, with a resolution of the form
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
-which by Lemma 12 yields a resolution
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
-From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [undefn, Lemma $9.3$].
-If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7).
-From [undefb, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$.
-∎
-Example 62 ($m(\sigma)=2,e=2$, $\operatorname{Bour}(\sigma)=3$ , $c_{3}=7$).
-Consider the following sequence with $d_{f}=1$, $d_{g}=2$:
-$\sigma=(-x_{0}x_{1}+x_{1}x_{2}-x_{2}x_{3},x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3}),$
-with corresponding Jacobian matrix given by:
-$\nabla\sigma=\begin{pmatrix}-x_{1}&x_{2}-x_{0}&x_{1}-x_{3}&-x_{2}\\ x_{1}^{2}&2x_{0}x_{1}&3x_{2}^{2}+2x_{2}x_{3}&x_{2}^{2}\end{pmatrix}$.
-The irreducible components of $(\Xi_{\sigma})_{1}$ consist only of a line $V(x_{1},x_{2})$.
-Consider $\mathfrak{p}=(x_{1},x_{2})$ the corresponding prime ideal in $R$, and let $u=(x_{2}-x_{0})^{-1}\in R_{\mathfrak{p}}$.
-The elementary operations $R_{2}^{\prime}=R_{2}-2ux_{0}x_{1}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ turn the matrix into
-$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&1&u(x_{1}-x_{3})&-ux_{2}\\ x_{1}^{2}+2ux_{0}x_{1}^{2}&0&x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})&x_{2}^{2}+2ux_{0}x_{1}x_{2}\end{pmatrix}$,
-which sends $e_{2}\in R_{\mathfrak{p}}^{4}$ to $(1,0)\in R_{\mathfrak{p}}^{2}$, so we may compute the cokernel $(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ by considering the second row and the remaining entries:
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I},$
-where
-$f_{1}$ $=x_{1}^{2}(1+2ux_{0})$
-$f_{2}$ $=x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})$
-$f_{3}$ $=x_{2}(x_{2}+2ux_{0}x_{1}).$
-Since $1+2ux_{0}\in R_{\mathfrak{p}}^{\times}$ and $f_{1}\in I$, we conclude $x_{1}^{2}\in I$. This simplifies $f_{2}\in I$ to:
-$x_{2}(3x_{2}+2x_{3})\equiv-(2ux_{0}x_{3})x_{1}\mod I,$
-so there is a unit $v\in R_{\mathfrak{p}}^{\times}$ such that $x_{1}\equiv vx_{2}\mod I$, which together with $x_{1}^{2}\in I$ implies $x_{2}^{2}\in I$. From this, $f_{3}\in I$ simplifies to
-$f\equiv 2ux_{0}x_{1}x_{2}\in I\Rightarrow x_{1}x_{2}\in I,$
-and therefore we have
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{1}\rangle,$
-with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=2$.
-Thus, $m(\sigma)=2$ (from the associativity formula), and $e=2$, yielding $\operatorname{Bour}(\sigma)=3$.
-The following resolution may also be obtained computationally:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
-describing the Bourbaki scheme as a twisted cubic and giving $c_{3}({\mathcal{T}}_{\sigma})=7$.
-Example 63 ($m(\sigma)=2$, $e=3$, $\operatorname{Bour}(\sigma)=5$, $c_{3}=3$).
-We consider the following sequence:
-$\sigma=(x_{2}x_{3}-x_{0}x_{1},x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{2}x_{3}^{2}+x_{3}^{3})$
-with corresponding Jacobian matrix given by
-$\nabla\sigma=\begin{pmatrix}-x_{1}&-x_{0}&x_{3}&x_{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}+x_{3}^{2}&x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2}\end{pmatrix}$.
-Here, there is a unique irreducible component of $(\Xi_{\sigma})_{1}$, the line $V(x_{0},x_{3})$.
-Considering $\mathfrak{p}=(x_{0},x_{3})$ the associated prime ideal, we compute $m(\sigma)=2$ using the associativity formula.
-Take $u=(x_{2})^{-1}\in R_{\mathfrak{p}}$ and consider the elementary operations $R_{2}^{\prime}=R_{2}-u(x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2})R_{1}\doteq R_{2}-uQR_{1}$, $R_{1}^{\prime}=uR_{1}$ so that
-$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&-ux_{0}&ux_{3}&1\\ 2x_{0}x_{2}+x_{1}x_{3}+uQx_{1}&x_{0}x_{3}+uQx_{0}&x_{0}^{2}+x_{3}^{2}-uQx_{3}&0\end{pmatrix}$.
-Arguing as usual, we may compute the cokernel by
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I}$
-with
-$f_{1}$ $=x_{0}(2x_{2}+ux_{1}^{2})+x_{1}x_{3}(2ux_{2}+3ux_{3}+1)$
-$f_{2}$ $=x_{0}(x_{3}(2ux_{2}+3ux_{3}+1)+ux_{0}x_{1})$
-$f_{3}$ $=x_{0}^{2}+x_{3}(x_{3}-ux_{0}x_{1}-2ux_{2}x_{3}-3ux_{3}^{2})$
-Writing
-$f_{2}-\frac{x_{0}}{x_{1}}f_{1}$ $=x_{0}\left(-\frac{x_{0}}{x_{1}}(2x_{2}+ux_{1}^{2})+ux_{0}x_{1}\right)$
-$=x_{0}^{2}\left(-\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)\in I,$
-and since the term inside parenthesis is invertible in $R_{\mathfrak{p}}$, we conclude $x_{0}^{2}\in I$.
-On the other hand, we consider
-$B\doteq f_{3}-x_{0}^{2}+\frac{x_{3}}{x_{1}}f_{1}$ $=-x_{3}\left(x_{0}\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{0}x_{1}-2x_{3}\right)$
-$=-x_{3}x_{0}\left(\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)+2x_{3}^{2}\in I.$
-and moreover, let
-$A\doteq x_{0}x_{3}(2ux_{2}+3ux_{3}+1)=f_{2}-ux_{0}^{2}x_{1}x_{3}\in I.$
-Then
-$A-\frac{3}{2}ux_{0}B$ $=x_{0}x_{3}(2ux_{2}+1)+3x_{3}^{2}-3x_{3}^{2}+\frac{3}{2}ux_{0}^{2}x_{3}\left(\frac{-(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right),$
-and after removing the term with $x_{0}^{2}$ factor, we obtain
-$x_{0}x_{3}(2ux_{2}+1)\in I\Rightarrow x_{0}x_{3}\in I,$
-since $2ux_{2}+1$ is invertible in $R_{\mathfrak{p}}$.
-Now, removing the $x_{0}x_{3}$ term in $B$ we conclude $x_{3}^{2}\in I$.
-So, up until now, we have $x_{0}^{2},x_{0}x_{3}$ and $x_{3}^{2}\in I$.
-Finally, considering
-$f_{1}-3ux_{1}x_{3}^{2}=x_{0}(2x_{2}+ux_{1}^{2})+x_{3}x_{1}(2ux_{2}+1)\in I,$
-this means there are $a,b\in R_{\mathfrak{p}}^{\times}$ such that $ax_{0}+bx_{3}\in I$. Therefore, we conclude that
-$I=(x_{0}^{2},x_{3}^{2},x_{0}x_{3},ax_{0}+bx_{3}),$
-and these mean that we can write
-$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{0}\rangle\simeq\langle 1,x_{3}\rangle$
-and obtain that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=1$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
-$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 3}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
-so we obtain $\operatorname{Bour}(\sigma)=5$ and $c_{3}({\mathcal{T}}_{\sigma})=3$.
-We summarize the results of this subsection in the following theorem:
+We summarize the main results of this subsection in the following theorem:
 Theorem D.
@@
   • (c)
-    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 56 and Example 14);
+    There are two cases of nearly-free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 56 and Example 15);
   • (d)